### v1

# Holomorphic Yang-Mills fields on $B$-branes

Abstract.
Considering the $B$-branes over a complex manifold $Y$ as objects of the bounded derived category $D^{b}(Y)$, we define holomorphic gauge fields on $B$-branes and the Yang-Mills functional for these fields. These definitions are a generalization to $B$-branes of concepts that are well known in the context of vector bundles. Given ${\mathscr{F}}^{\bullet}\in D^{b}(Y),$ we show that the Atiyah class $a({\mathscr{F}}^{\bullet})\in{\rm Ext}^{1}({\mathscr{F}}^{\bullet},\,\Omega^{1}({\mathscr{F}}^{\bullet}))$ is the obstruction to the existence of gauge fields on ${\mathscr{F}}^{\bullet}.$ When $Y$ be either the projective space ${\mathbb{P}}^{n}$ or the variety of complete flags in ${\mathbb{C}}^{3},$ we prove that the cardinal of the set of holomorphic gauge fields on any $B$-brane over $Y$ is $\leq 1.$ We prove that the set of Yang-Mills fields on the $B$-brane ${\mathscr{F}}^{\bullet},$ if it is nonempty, is in bijective correspondence with the points of an algebraic subset of ${\mathbb{C}}^{m}$ defined by $m$ polynomial equations of degree $\leq 3$, where $m={\rm dim}\,{\rm Ext}^{0}({\mathscr{F}}^{\bullet},\,\Omega^{1}({\mathscr{F}}^{\bullet})).$
Key words and phrases:
Yang-Mills fields, holomorphic connections, coherent reflexive sheaves
MSC 2020: 53C05, 58E15, 18G10
1. Introduction
In this article, we extend the well-known concepts of gauge field and Yang-Mills field on vector bundles to $B$-branes. From the mathematical point of view, a $B$-brane over a complex manifold $Y$ is an object of $D^{b}(Y)$, the bounded derived category of coherent analytic sheaves over $Y$ [2, Sect. 5.4] [3, Sect. 5.3].
The simplest $B$-banes are the holomorphic vector bundles. A gauge field on a holomorphic vector bundle $V\to M,$ in mathematical terms a connection on $V,$ defines a derivation of sections along vectors of $V$ giving rise to sections not necessarily holomorphic. A holomorphic gauge field on $V,$ i., e. holomorphic connection on $V$ [4], allows us to define a derivative of the holomorphic sections of $V$ along any “direction”, giving rise to holomorphic sections.
But no every holomorphic vector bundle admits a holomorphic connection. The obstruction for this existence is the Atiyah class of the bundle [4]. Hence, the vanishing of this class is a necessary condition for the existence of such gauge fields on the vector bundle. Assuming that there is no obstruction to the existence of holomorphic gauge fields, the set of these fields on a vector bundle over a compact manifold is a finite dimensional affine space, in contrast to the smooth case. This property is a consequence of the finite dimensionality of the coherent cohomology groups.
The two features discussed above concerning the holomorphic connections on vector bundles: the possible obstruction to its existence, and the finite dimension of the space of those fields, when it is not empty, are also present in the extension of the holomorphic gauge fields to general branes.
Gauge fields on coherent sheaves. A gauge field on $V$ determines an isomorphism between the fibers over “close” points of $M$. In fact, the connection is determined by the identification of the fibers at points “infinitesimally close” in $M;$ i. e. by the identifications in the directions defined by the tangent vectors of the base. The idea of being “infinitesimally close” can be formulated by means of the first infinitesimal neighborhood of the diagonal of $M$. In this setting, a gauge field on $V$ can be regarded as a right inverse of the projection $\pi:J^{1}(V)\to V,$ where $J^{1}(V)$ is the $1$-jet bundle of $V$ [30].
That approach admits a natural translation to the context of the coherent sheaves over complex manifolds. In this way, if ${\mathscr{F}}$ is a coherent sheaf over the compact analytic manifold $Y,$ we define a holomorphic gauge field on ${\mathscr{F}}$ as a right inverse of the natural morphism ${\mathscr{J}}^{1}({\mathscr{F}})\to{\mathscr{F}}$, where ${\mathscr{J}}^{1}({\mathscr{F}})$ is the corresponding $1$-jet sheaf. This inverse determines a morphism of abelian sheaves
$\nabla:{\mathscr{F}}\to\Omega^{1}({\mathscr{F}})=\Omega^{1}_{Y}\otimes_{{\mathscr{O}}_{Y}}{\mathscr{F}},$
which satisfies the Leibniz’s rule. In this context, the obstruction to the existence of a holomorphic connection on the sheaf ${\mathscr{F}}$ is an element $a({\mathscr{F}})$ of the group ${\rm Ext}^{1}({\mathscr{F}},\,\Omega^{1}({\mathscr{F}}))$. Furthermore, when the set of holomorphic gauge fields on ${\mathscr{F}}$ is nonempty, it is an affine space associated to the finite dimensional vector space ${\rm Hom}({\mathscr{F}},\,\Omega^{1}({\mathscr{F}}))$.
Yang-Mills fields on shaves. Setting ${\mathcal{S}}$ for the singularity set of ${\mathscr{F}}$, let us assume that on the locally free sheaf ${\mathscr{F}}|_{Y\setminus{\mathcal{S}}}$ there is defined a Hermitian metric, we say that ${\mathscr{F}}$ is a Hermitian sheaf. If $Y$ is a Kähler manifold, on the set of holomorphic connections over ${\mathscr{F}}$, one defines the Yang-Mills functional $\mathcal{YM}$. Essentially, the value of $\mathcal{YM}$ at a connection $\nabla$ is the squared norm $\lVert K_{\nabla}\rVert^{2}$ of the curvature of that connection [14, page 417], [27, page 44].
The stationary points of functional $\mathcal{YM}$ are the Yang-Mills gauge fields on ${\mathscr{F}}$. The set of critical points of this functional will be denoted by ${\sf YM}({\mathscr{F}})$.
The reflexive sheaves might be thought as “vector bundles with singularities” [16, page 121], and these singularities can in some way be avoided. From this property, and denoting $m:={\rm dim}\,{\rm Hom}({\mathscr{F}},\,\Omega^{1}({\mathscr{F}})),$ we will prove that if the reflexive sheaf ${\mathscr{F}}$ admits holomorphic gauge fields, then the set ${\sf YM}({\mathscr{F}})$ is in bijective correspondence with the points of an algebraic subset of ${\mathbb{C}}^{m}$ defined by $m$ algebraic equations of degree $\leq 3$ (Theorem 1). Thus, in general, the set of Yang-Mills fields on ${\mathscr{F}}$ will be a finite set.
Yang-Mills fields on a $B$-brane. As we said, a $B$-brane on $Y$ is a complex $({\mathscr{F}}^{\bullet},\,\delta^{\bullet})$ of analytic coherent sheaves on $Y.$ According to the preceding paragraphs, it is reasonable to define a gauge field on this brane as an element of ${\rm Hom}_{D^{b}(Y)}\big{(}{\mathscr{F}}^{\bullet},\,{\mathscr{J}}^{1}({\mathscr{F}}^{\bullet})\big{)}$ which lifts the identity on ${\mathscr{F}}^{\bullet}.$
Let $X$ be either the projective space ${\mathbb{P}}^{n}$ or the variety of complete flags in ${\mathbb{C}}^{3}.$ We prove that the cardinal of the set of holomorphic gauge fields on any $B$-brane over $X$ is $\leq 1$ (Theorems 3 and 4). For the proof of that result we use the existence of finite families of objects in $D^{b}(X)$ which generate those categories. A well-known theorem of Beilinson asserts that the set $\{{\mathscr{O}}_{{\mathbb{P}}^{n}}(-n),\,{\mathscr{O}}_{{\mathbb{P}}^{n}}(-n+1),\dots,{\mathscr{O}}_{{\mathbb{P}}^{n}}(-1),\,{\mathscr{O}}_{{\mathbb{P}}^{n}}\}$ is a full exceptional family in $D^{b}({\mathbb{P}}^{n})$ [19, Sect 8.3]. However, for the proof of the above result we need to construct a more appropriate generator set. For the case when $X$ is the mentioned flag variety, we determine a family of generators $D^{b}(X)$ following an idea from [5].
If $Y$ is a Hodge manifold, then it is a projective smooth variety and by the GAGA correspondence the analytic coherent sheaves on $Y$ can be considered as algebraic ones. Moreover, the category $D^{b}(Y)$ is equivalent to the homotopy category $K^{b}(Y)_{\rm coh}$ of complexes of injective sheaves with coherent and bounded cohomology. In this way, a gauge field $\psi$ on ${\mathscr{F}}^{\bullet}$ can be considered as a homotopy class of morphisms between the complexes of abelian sheaves ${\mathscr{F}}^{\bullet}\to\widehat{\Omega^{1}({\mathscr{F}}^{\bullet})}$, where $\widehat{\Omega^{1}({\mathscr{F}}^{\bullet})}$ is a complex in $K^{b}(Y)_{\rm coh}$ quasi-isomorphic to $\Omega^{1}({\mathscr{F}}^{\bullet})$. Thus, that homotopy class determines a unique morphism between each of the cohomology sheaves $\vartheta^{j}:{\mathscr{H}}^{j}({\mathscr{F}}^{\bullet})\to{\mathscr{H}}^{j}(\Omega^{1}({\mathscr{F}}^{\bullet})).$ In summary, the gauge field on ${\mathscr{F}}^{\bullet}$ defines a family of connections $\vartheta^{j}$ on the cohomology sheaves ${\mathscr{H}}^{j}({\mathscr{F}}^{\bullet})$.
When the cohomology sheaves ${\mathscr{H}}^{j}$ are Hermitian, we define the value of the Yang-Mills functional on the above gauge $\psi$ as $\sum_{i}(-1)^{i}\lVert K_{\vartheta_{i}}\rVert^{2}.$ Thus, the Yang-Mills functional is a kind of Euler characteristic of the gauge field. Obviously, this definition of the Yang-Mills functional on branes generalizes the one given for coherent sheaves.
The gauge field $\psi$ is a Yang-Mills field if it is a stationary point of Yang-Mills functional. We will prove that if the $\vartheta^{j}$ are Yang-Mills fields on the ${\mathscr{H}}^{j}$, so is $\psi$ on the brane (Proposition 14). Theorem 6 is the converse to Proposition 14. In Theorem 5, we generalize the result given in Theorem 1, about the cardinal of the set of Yang-Mills fields on sheaves, to a general brane ${\mathscr{F}}^{\bullet}$.
The article is organized in two sections. In Section 2, we define the gauge fields on a coherent sheaf. In the first subsections of this section, we revise the definition of the $1$-jet sheaf of a coherent sheaf ${\mathscr{F}}$ on a complex manifold $Y,$ because, although this is well known in algebraic geometry, it is not so well known in the community of mathematical physicists. The Yang-Mills functional is introduced in Subsection 2.2, where Theorem 1 is proved.
In Section 3 are considered the gauge fields on a general $B$-brane. Subsection 3.1 defines these fields and discusses the obstruction to their existence. Subsection 3.2 is devoted to the study of gauge fields on $B$-branes over ${\mathbb{P}}^{n};$ in this subsection we prove the above Theorem 3 concerning the cardinal of the set of such fields. Theorem 4, about the cardinal of the set of fields on $B$-branes over the above flag variety, is proved in Subsection 3.3. We show, in Subsection 3.4, how a gauge field on a $B$-brane field gives rise to connections in cohomologies. Finally, we define the Yang-Mills functional on gauge fields over general $B$-branes, showing the reasons on which this definition is based. We also prove Theorems 5 and 6 and Proposition 14.
2. Gauge fields on coherent sheaves
2.1. Gauge fields on a sheaf.
The definition of connection on a coherent sheaf, as it is introduced for example in [23], is reformulated in this section, so that it is suitable for translation to objects in the derived category of coherent sheaves.
As explained in the Introduction, the existence of a gauge field on a coherent sheaf over a complex manifold $Y$ should define an isomorphism between the stalks of the sheaf at any “infinitesimally close” points of $Y.$ In case Y is an algebraic variety the neighborhoods in the Zariski topology are too large to express the idea of “infinitesimal proximity”. Instead by means of the first infinitesimal neighborhood $Y^{(1)}$ of $Y,$ [13, page 698], it is possible to formulate this idea in a suitable way also for algebraic varieties.
If $R$ is a ${\mathbb{C}}$-algebra, ${\rm Hom}({\rm Spec}\,R,\,X)$ is the set of points of complex algebraic variety $X$ with values in $R$. Two points $x,y$ are infinitesimally close if the morphism $(x,\,y):{\rm Spec}\,R\to X\times X$ factorizes through $X^{(1)}$ [7, page 6]. Therefore, following Deligne, one defines a gauge field on the coherent ${\mathscr{O}}_{X}$-module ${\mathscr{G}}$ as an element of
(2.1) ${\rm Hom}_{X^{(1)}}(\pi_{1}^{*}{\mathscr{G}},\,\pi_{2}^{*}{\mathscr{G}}),$
which is the identity on $X$, where $\pi_{1},\pi_{2}:X^{(1)}\to X$ are the projections and $\pi_{i}^{*}$ the corresponding inverse image functor. By the adjunction isomorphism, a gauge field on ${\mathscr{G}}$ is a morphism ${\mathscr{G}}\to\pi_{1*}\pi^{*}_{2}{\mathscr{G}}$ which is a right inverse of the projection $\pi:\pi_{1*}\pi^{*}_{2}{\mathscr{G}}\to{\mathscr{G}}.$
As $\pi_{1*}\pi^{*}_{2}{\mathscr{G}}$ is the first jet sheaf of the coherent sheaf ${\mathscr{G}}$, in the following paragraphs we review the definition of the jet sheaf and also that of first neighborhood of the diagonal.
2.1.1. Neighborhood of the diagonal
Let $Y$ be an analytic complex compact connected manifold. We denote by $i:\Delta\hookrightarrow Y\times Y$ the embedding of the diagonal. As a closed subvariety, $\Delta$ is defined by an ideal $\mathscr{I}$ of $\hat{\mathscr{O}}:={\mathscr{O}}_{Y\times Y}.$ We will consider the following ringed spaces
$(Y,\,{\mathscr{O}}),\;\;Y^{(1)}=\Big{(}\Delta,\,{\mathscr{O}}_{Y^{(1)}}:=\big{(}\hat{\mathscr{O}}/{\mathscr{I}}^{2}\big{)}|_{\Delta}\Big{)},\;\;(Y\times Y,\,\hat{\mathscr{O}}).$
$Y^{(1)}$ is the first infinitesimal neighborhood of $\Delta$.
[…]
Given $f\in{\mathscr{O}}$ and $l\in\hat{\mathscr{O}}$, the product
(2.3) $(l+{\mathscr{I}}^{2})\cdot f=l\,(f\circ p_{2})+{\mathscr{I}}^{2}$
defines a right ${\mathscr{O}}$-module structure on $\hat{\mathscr{O}}/{\mathscr{I}}^{2}$. More explicitly, $(l(x,\,y)+{\mathscr{I}}^{2})\cdot f=l(x,\,y)f(y)+{\mathscr{I}}^{2}.$ Analogously, $f\cdot(l+{\mathscr{I}}^{2})=(f\circ p_{1})l+{\mathscr{I}}^{2}$ gives to $\hat{\mathscr{O}}/{\mathscr{I}}^{2}$ a left ${\mathscr{O}}$-module structure. However, the restrictions to ${\mathscr{I}}/{\mathscr{I}}^{2}$ of these left and right ${\mathscr{O}}$-module structures are equivalent.
The isomorphism ${\mathscr{O}}\oplus\Omega^{1}\simeq{\mathscr{O}}_{Y^{(1)}}$. The cotangent sheaf $\Omega^{1}:=\Omega^{1}_{Y}$ can be identified with the pullback $j^{-1}({\mathscr{I}}/{\mathscr{I}}^{2})$ [8, p. 407]. This identification is defined by the correspondence
(2.4) $\xi:j^{-1}({\mathscr{I}}/{\mathscr{I}}^{2})\to\Omega^{1},\;\;\;g+{\mathscr{I}}^{2}\mapsto\big{(}d_{x}g\big{)}|_{Y},$
where $d_{x}g$ is the exterior derivative of $g$ with respect to the variables $x$; i.e. considering $g$ as function of the variables $x$ and keeping the variables $y$ constant.
If $l\in\hat{\mathscr{O}}$, the correspondence $l+{\mathscr{I}}^{2}\mapsto l\circ j\oplus(d_{x}l)|_{Y}$ defines an isomorphism of right ${\mathscr{O}}$-modules
[…]
When the abelian sheaf ${\mathscr{O}}\oplus\Omega^{1}$ is endowed with the left ${\mathscr{O}}$-action
(2.6) $f\cdot(h\oplus\alpha):=fh\oplus(hdf+f\alpha),$
where $\alpha\in\Omega^{1}$, then $m$ in (2.5) is an isomorphism of left ${\mathscr{O}}$-modules. We summarize those well-known results in the following proposition.
Proposition 1.
With the notations above introduced:
  1. (1)
    The ${\mathscr{O}}$-modules $j^{-1}({\mathscr{I}}/{\mathscr{I}}^{2})$ and $\Omega^{1}$ are canonically isomorphic.
  2. (2)
    The correspondence (2.5) defines an isomorphism between the right ${\mathcal{O}}$-modules $j^{-1}(\hat{\mathscr{O}}/{\mathscr{I}}^{2})$ and ${\mathcal{O}}\oplus{\Omega}^{1}$.
  3. (3)
    Equipped ${\mathcal{O}}\oplus{\Omega}^{1}$ with the left ${\mathscr{O}}$-structure defined in (2.6), the correspondence (2.5) is also an isomorphism of left ${\mathscr{O}}$-modules.
The exact sequence of right $\hat{\mathscr{O}}$-modules
$0\to{\mathscr{I}}/{\mathscr{I}}^{2}\to\hat{\mathscr{O}}/{\mathscr{I}}^{2}\to\hat{\mathscr{O}}/{\mathscr{I}}\to 0$
gives rise, by means the functor $j^{-1}$, the exact sequence of right ${\mathscr{O}}$-modules
(2.7) $0\to j^{-1}({\mathscr{I}}/{\mathscr{I}}^{2})\to j^{-1}(\hat{\mathscr{O}}/{\mathscr{I}}^{2})\to j^{-1}(\hat{\mathscr{O}}/{\mathscr{I}})\to 0,$
or in other terms $0\to\Omega^{1}\to{\mathscr{O}}\oplus\Omega^{1}\to{\mathscr{O}}\to 0$
2.1.2. The first jet sheaf
We denote by $\pi_{a}=p_{a}\circ i:\Delta\to Y$, for $a=1,2$. One has the following morphism of sheaves rings over $\Delta$ (see Subsection 2.1.1)
$\pi_{a}^{-1}{\mathscr{O}}\longrightarrow{\mathscr{O}}_{Y^{(1)}}={\hat{\mathscr{O}}}/{\mathscr{I}^{2}},\;\;\;\;h\mapsto h\circ\pi_{a}+{\mathscr{I}}^{2}.$
In particular, one can consider ${\mathscr{O}}_{Y^{(1)}}$ as a right ${\pi_{2}^{-1}{\mathscr{O}}}$-module.
Given ${\mathscr{F}}$ a left ${\mathscr{O}}$-module on $Y,$ its inverse image by $\pi_{2}$ is left ${\mathscr{O}}_{Y^{(1)}}$-module
$\pi_{2}^{*}({\mathscr{F}})={\mathscr{O}}_{Y^{(1)}}\otimes_{\pi_{2}^{-1}{\mathscr{O}}}\pi_{2}^{-1}{\mathscr{F}}.$
And the first jet sheaf ${\mathscr{J}}^{1}({\mathscr{F}})$ of ${\mathscr{F}}$ is left ${\mathscr{O}}$-module defined by
(2.8) ${\mathscr{J}}^{1}({\mathscr{F}})={\pi_{1}}_{*}\pi_{2}^{*}({\mathscr{F}}).$
Since $\pi_{1}^{*}$ is the left adjoint of ${\pi_{1}}_{*}$, one has
(2.9) ${\rm Hom}\,_{{\mathscr{O}}}\big{(}{\mathscr{F}},\,{\mathscr{J}}^{1}({\mathscr{F}})\big{)}={\rm Hom}\,_{{\mathscr{O}}_{Y^{(1)}}}\big{(}\pi_{1}^{*}{\mathscr{F}},\,\pi_{2}^{*}{\mathscr{F}}\big{)}.$
If ${\mathscr{H}}$ is a left ${\mathscr{O}}_{Y^{(1)}}$-module, the ${\mathscr{O}}$-structure on ${\pi_{1}}_{*}{\mathscr{H}}$ is defined by $f\cdot s=(f\circ\pi_{1})s$, where $f\in{\mathscr{O}}$ and $s\in{\mathscr{H}}$. On the other hand, the ${\mathscr{O}}$ action on
$k^{*}{\mathscr{H}}={\mathscr{O}}\otimes_{k^{-1}{\mathscr{O}}_{\Delta}}k^{-1}{\mathscr{H}}$
is determined by $f\cdot(1\otimes s)=f\otimes s=1\otimes gs=1\otimes(f\circ\pi_{1})s$, where $f=g\circ k$. Hence, $k^{*}{\mathscr{H}}$ and ${\pi_{1}}_{*}{\mathscr{H}}$ are isomorphic ${\mathscr{O}}$-modules. Thus, as $\pi_{2}\circ k={\rm id},$
(2.10) ${\mathscr{J}}^{1}({\mathscr{F}})=k^{-1}({\mathscr{O}}_{Y^{(1)}})\otimes_{{\mathscr{O}}}{\mathscr{F}}=j^{-1}(\hat{\mathscr{O}}/{\mathscr{I}}^{2})\otimes_{{\mathscr{O}}}{\mathscr{F}}.$
By Proposition 1, ${\mathscr{J}}^{1}({\mathscr{F}})$ is the abelian sheaf ${\mathscr{F}}\oplus\Omega^{1}({\mathscr{F}})$ endowed with the following left ${\mathscr{O}}$-module structure
(2.11) $f\ast(\sigma\oplus\beta)=f\sigma\oplus(f\beta+df\otimes\sigma).$
One has the morphism of abelian sheaves
$\eta:{\mathscr{F}}\to{\mathscr{J}}^{1}({\mathscr{F}})={\mathscr{F}}\oplus\Omega^{1}({\mathscr{F}}),\;\;\sigma\mapsto\sigma\oplus 0.$
And from (2.11), it follows
(2.12) $f\eta(\sigma)=\eta(f\sigma)+df\otimes\sigma.$
Summarizing, one has the exact sequence of ${\mathscr{O}}$-modules
[…]
Thus, taking into account (2.1) and (2.9), a gauge field on ${\mathscr{F}}$ will be an element $\psi\in{\rm Hom}({\mathscr{F}},\,{\mathscr{J}}^{1}({\mathscr{F}}))$ right inverse of $\pi$. That is, we give the following definition.
Definition 1.
A holomorphic gauge field on the coherent sheaf ${\mathscr{F}}$ is a right inverse of the ${\mathscr{O}}$-module morphism $\pi:{\mathscr{J}}^{1}({\mathscr{F}})\to{\mathscr{F}}.$
We introduce a piece of notation that will be used later on. Let ${\sf Forg}$ denote the forgetful functor from the category of ${\mathscr{O}}$-modules to the category of abelian sheaves over $Y.$ We set
$\boldsymbol{\mathscr{J}}^{1}({\mathscr{F}})={\sf Forg}\big{(}{\mathscr{F}}\oplus\Omega^{1}({\mathscr{F}})\big{)}.$
In this way ${\mathscr{J}}^{1}({\mathscr{F}})$ is the abelian sheaf $\boldsymbol{\mathscr{J}}^{1}({\mathscr{F}})$ endowed with the ${\mathscr{O}}$-module structure defined in (2.11).
2.1.3. The Atiyah class.
The exact sequence (2.13) defines an element $a({\mathscr{F}})$ of ${\rm Ext}^{1}({\mathscr{F}},\,\Omega^{1}({\mathscr{F}}))$, called the Atiyah class of ${\mathscr{F}}$. The exact sequence (2.13) does not split, in general, in the category of ${\mathscr{O}}$-modules. Thus, there will be exist gauge fields on ${\mathscr{F}}$ iff the Atiyah class $a({\mathscr{F}})$ vanishes.
If there exists a right inverse $\psi$ of $\pi$; then $\pi(\psi-\eta)=0$; and thus $\nabla:=\psi-\eta$ takes its values in $\Omega^{1}({\mathscr{F}})$. Moreover, by (2.12), for $f\in{\mathscr{O}}$ and $\sigma\in{\mathscr{F}}$
(2.14) $\nabla(f\sigma)=f\psi(\sigma)-f\eta(\sigma)+df\otimes\sigma=f\nabla(\sigma)+df\otimes\sigma;$
that is, for $\nabla$ holds the Leibniz’s rule satisfied by the connections on vector bundles.
We denote by ${\mathcal{C}}$ the set consisting of all morphisms of ${\mathbb{C}}_{Y}$-modules $\nabla:{\mathscr{F}}\to\Omega^{1}({\mathscr{F}})$ satisfying (2.14). An element $\nabla\in{\mathcal{C}}$ determines the morphism of ${\mathscr{O}}$-modules
$\kappa:{\mathscr{F}}\oplus\Omega^{1}({\mathscr{F}})\to\Omega^{1}({\mathscr{F}}),\;\;(\sigma,\beta)\mapsto\beta-\nabla(\sigma).$
$\kappa$ is a left inverse of $\iota$. Hence, $\kappa$ defines a splitting of the extension (2.13) and thus it determines a right inverse of $\pi$; i. e. a holomorphic gauge field $\psi$ on ${\mathscr{F}}$. We have proved the following proposition.
Proposition 2.
If $a({\mathscr{F}})=0$, then the map $\nabla\mapsto\nabla+\eta$ defines a bijective correspondence between ${\mathcal{C}}$ and the set of holomorphic gauge fields on the coherent sheaf ${\mathscr{F}}$.
Proposition 3.
The set of holomorphic gauge fields on the coherent sheaf ${\mathscr{F}}$, if it is nonempty, is an affine space associated to the finite dimensional vector space $\Gamma\big{(}Y,\,{\mathscr{H}om}({\mathscr{F}},\Omega^{1}({\mathscr{F}}))\big{)}.$
Proof. The exact sequence (2.13) gives rise to corresponding ${\rm Ext}$-sequence
[…]
Since $\mu(\phi)=\pi\circ\phi$, the existence of a holomorphic gauge field $\psi$ on ${\mathscr{F}}$, is equivalent to $1_{\mathscr{F}}\in{\rm im}(\mu)={\ker}(\nu)$. In fact, the Atiyah class $a({\mathscr{F}})$ is the image of $1_{\mathscr{F}}$ by $\nu$.
If $\psi$ and $\psi_{1}$ are gauge fields on the coherent sheaf ${\mathscr{F}}$, then $\mu(\psi_{1}-\psi)=0$; i. e. $\psi_{1}-\psi\in{\rm im}(\lambda)$. Thus, the set of holomorphic gauge fields on ${\mathscr{F}}$, if nonempty, is an affine space with vector space ${\rm Hom}({\mathscr{F}},\Omega^{1}({\mathscr{F}})).$ This vector space is the space of sections $\Gamma(Y,\,{\mathscr{H}om}({\mathscr{F}},\Omega^{1}({\mathscr{F}}))$. As ${\mathscr{H}om}({\mathscr{F}},\Omega^{1}({\mathscr{F}}))$ is a coherent sheaf, the space of holomorphic gauge fields on ${\mathscr{F}}$, is a finite dimensional affine space [13, page 700 ]. ∎
Remark 1.
Taking into account the isomorphism [4, Prop. 2]
${\rm Ext}^{1}({\mathscr{F}},\,\Omega^{1}({\mathscr{F}}))\simeq H^{1}(Y,\,{\mathscr{H}om}({\mathscr{F}},\,\Omega^{1}({\mathscr{F}}))),$
$a({\mathscr{F}})$ can be considered as a cohomology class. In particular, when ${\mathscr{F}}$ is a locally free sheaf, ${\mathscr{H}om}({\mathscr{F}},\,\Omega^{1}({\mathscr{F}}))\simeq\Omega^{1}\otimes_{\mathscr{O}}{\mathscr{E}nd}({\mathscr{F}})$, and $a({\mathscr{F}})$ is defined by a $\check{\rm C}$ech cocycle with coefficients in $\Omega^{1}\otimes_{\mathscr{O}}{\mathscr{E}nd}({\mathscr{F}})$ (see [18, page 180]).
2.2. Holomorphic Yang-Mills fields
2.2.1. Hermitian sheaves
The fiber at $x\in Y$ of a coherent ${\mathscr{O}}$-module ${\mathscr{G}}$ will be denoted by ${\mathscr{G}}_{(x)}:={\mathscr{G}}_{x}/{\mathfrak{m}}_{x}{\mathscr{G}}_{x}$, where ${\mathfrak{m}}_{x}$ is the maximal ideal of ${\mathscr{O}}_{x}$. If ${\mathcal{Z}}$ is a section of ${\mathscr{G}}$ the corresponding vector in ${\mathscr{G}}_{(x)}$ is denoted by $Z(x)$. On the other hand, the singular set ${\mathcal{S}}$ of ${\mathscr{G}}$ is a closed analytic subset of $Y$ whose codimension is greater than or equal to $1$ [25, Chap. V, Thm. 5.8]. Moreover, ${\mathscr{G}}$ is locally free on $Y\setminus{\mathcal{S}}$. We set ${G}$ for the vector bundle over $Y\setminus{\mathcal{S}}$, with fibers $G(x):={\mathscr{G}}_{(x)}$, determined by the locally free sheaf ${\mathscr{G}}|_{Y\setminus{\mathcal{S}}}$.
The following definition generalizes the given in [32, Chap. III, Sect 1] for locally free sheaves.
Definition 2.
A Hermitian metric on the coherent sheaf ${\mathscr{G}}$ is a set $\{\langle\,,\,\rangle_{x}\}_{x\in Y}$ of Hermitian metrics on the fibers of ${\mathscr{G}}$, satisfying the following condition: Given ${\mathcal{Z}}_{1},{\mathcal{Z}}_{2}$ two sections of ${\mathscr{G}}$ on an open $U$ of $Y,$ the map
${\mathcal{Z}}_{1}\cdot{\mathcal{Z}}_{2}:x\in U\mapsto\langle Z_{1}(x),\,Z_{2}(x)\rangle_{x}\in{\mathbb{C}}$
is bounded, and its restriction to $U\setminus{\mathcal{S}}$ is $C^{\infty}.$ A sheaf endowed with a Hermitian metric is called a Hermitian sheaf.
Given ${\mathcal{Z}}_{1},{\mathcal{Z}}_{2}\in\Gamma(Y,\,{\mathscr{G}})$, ${\mathcal{Z}}_{1}\cdot{\mathcal{Z}}_{2}$ is a bounded map with possible discontinuities in the singular locus. If $Y$ is a Kähler manifold, the function ${\mathcal{Z}}_{1}\cdot{\mathcal{Z}}_{2}$ may be non-integrable on $Y$ with respect to the volume form defined by the Kähler metric.
If ${\mathscr{G}}$ is a torsion-free coherent sheaf on the Kähler manifold $Y,$ then ${\rm codim}\,{\mathcal{S}}\geq 2$ [25, Cap. V, Cor. 5.15]. Hence, there exists an analytic closed subset ${\mathcal{W}}$ of $Y,$ containing ${\mathcal{S}},$ and such that $1\leq{\rm codim}\,{\mathcal{W}}<{\rm codim}\,{\mathcal{S}}.$ Let ${\mathcal{N}}$ be a tubular neighborhood of ${\mathcal{S}}$ in ${\mathcal{W}}.$ Then
  1. (1)
    ${\mathcal{Z}}_{1}\cdot{\mathcal{Z}}_{2}$ is a continuous map on the closed space $Y\setminus{\mathcal{N}},$ which is disjoint with the singularity set ${\mathcal{S}}.$
  2. (2)
    ${\mathcal{W}},$ as a subset of $Y$ with codimension $\geq 1,$ has measure zero with respect to the differential form ${\rm d}{\rm vol}.$
Thus, the following integral is well-defined
(2.15) $({\mathcal{Z}}_{1},\,{\mathcal{Z}}_{2}):=\int_{Y\setminus{\mathcal{N}}}{\mathcal{Z}}_{1}\cdot{\mathcal{Z}}_{2}\,{\rm d}{\rm vol}.$
From the boundedness of ${\mathcal{Z}}_{1}\cdot{\mathcal{Z}}_{2}$ together with $(2),$ it follows that this definition does not depend on the choices of the analytic closed set ${\mathcal{W}}$ and the neighborhood ${\mathcal{N}}.$
Remark 2.
[…]
2.2.2. Yang-Mills functional
Given $\nabla$ a holomorphic connection on the coherent sheaf $\mathscr{F},$ it defines a morphism of ${\mathbb{C}}_{Y}$-modules $\nabla^{(k)}:\Omega^{k}({\mathscr{F}}){\to}\Omega^{k+1}({\mathscr{F}})$ in the usual way. The composition ${\mathcal{K}}_{\nabla}:=\nabla^{(1)}\circ\nabla:{\mathscr{F}}\to\Omega^{2}({\mathscr{F}})$ is the curvature of $\nabla$; and as it is well-known
(2.16) ${\mathcal{K}}_{\nabla}\in{\rm Hom}({\mathscr{F}},\,\Omega^{2}({\mathscr{F}}))=\Gamma(Y,\,{\mathscr{H}om}({\mathscr{F}},\,\Omega^{2}({\mathscr{F}}))).$
The connection is said to be flat if ${\mathcal{K}}_{\nabla}=0$.
For each $x\in Y,$ we denote by $\alpha_{x}$ and $\lambda_{x}$ the natural morphisms
$\big{(}{\mathscr{H}om}_{\mathscr{O}}({\mathscr{F}},\,\Omega^{k}({\mathscr{F}}))\big{)}_{x}\overset{\alpha_{x}}{\to}{\rm Hom}_{{\mathscr{O}}_{x}}\big{(}{\mathscr{F}}_{x},\,\Omega^{k}_{x}\otimes_{{\mathscr{O}}_{x}}{\mathscr{F}}_{x}\big{)}\overset{\lambda_{x}}{\leftarrow}\Omega^{k}_{x}\otimes_{{\mathscr{O}}_{x}}{\rm End}_{{\mathscr{O}}_{x}}({\mathscr{F}}_{x}).$
As ${\mathscr{F}}$ is coherent, $\alpha_{x}$ is isomorphism [12, page 239]. Furthermore, if ${\mathscr{F}}_{x}$ is free, then $\lambda_{x}$ is bijective.
If ${\mathcal{S}}$ is the singularity set of ${\mathscr{F}},$ it is also the singular locus of ${\mathscr{H}om}({\mathscr{F}},\,\Omega^{2}({\mathscr{F}})).$ Hence, for each point $x$ outside of ${\mathcal{S}},$ the fiber of ${\mathscr{H}om}({\mathscr{F}},\,\Omega^{k}({\mathscr{F}}))$ at $x$ can be identified with the vector space $\Omega^{k}{(x)}\otimes{\rm End}({F}{(x)}).$ According to (2.16), the curvature ${\mathcal{K}}_{\nabla}$ of a holomorphic connection $\nabla$ determines the vector
(2.17) $K_{\nabla}(x)\in\Omega^{2}(x)\otimes{\rm End}(F(x))$
for each $x\in Y\setminus{\mathcal{S}}$. That is, $K_{\nabla}$ is a $2$-form ${\rm End}(F)$-valued.
Let us assume that ${\mathscr{F}}$ is a Hermitian torsion-free sheaf on the Kähler manifold $Y,$ according to (2.15), one defines
(2.18) $\lVert{\mathcal{K}}_{\nabla}\rVert^{2}=({\mathcal{K}}_{\nabla},\,{\mathcal{K}}_{\nabla})=\int_{Y\setminus{\mathcal{N}}}{\mathcal{K}}_{\nabla}\cdot{\mathcal{K}}_{\nabla}\,{\rm d}{\rm vol}=\int_{Y\setminus{\mathcal{N}}}|K_{\nabla}\wedge\star\,K_{\nabla}|,$
[…]
More concretely, if locally $K_{\nabla}$ can be expressed as $\alpha\otimes A$, with $\alpha$ a $2$-form and $A$ a local section of ${\rm End}(F)$, then the integrand in (2.18) is $(\alpha\wedge\star\alpha)\langle A\circ A\rangle$. In a local unitary frame of ${\rm End}(F)$, if the connection is compatilble with the metric, the matrix $\check{A}$ associated to $A$ is antihermitian and $\langle A\circ A\rangle=-{\rm tr}(\check{A}\check{A})$. That is,
(2.19) $\langle K_{\nabla}\wedge\star\,K_{\nabla}\rangle=-{\rm tr}\big{(}K_{\nabla}\wedge\star\,K_{\nabla}\big{)}.$
On the space of holomorphic gauge fields on the torsion-free sheaf ${\mathscr{F}}$ one can define the following map
$\mathcal{YM}:\nabla\in{\rm Hom}({\mathscr{F}},\,\Omega^{1}({\mathscr{F}}))\mapsto\lVert{\mathcal{K}}_{\nabla}\rVert^{2}=\int_{Y\setminus{\mathcal{N}}}{\mathcal{K}}_{\nabla}\cdot{\mathcal{K}}_{\nabla}\,{\rm d}{\rm vol}.$
It is called the Yang-Mills’ functional. The $\nabla$ on which this functional takes a stationary value are the holomorphic Yang-Mills fields.
Remark 3.
If ${\mathscr{F}}$ admits a flat connection, then ${\mathscr{F}}$ is a locally free sheaf [10, Chap. 8, Prop. 2.3], [20, Thm. 1.4.10], [23, Prop. 8.8]. In this case the minimum value taken by the corresponding Yang-Mills functional is obviously zero. Conversely, if ${\mathscr{F}}$ is locally free, then the minimum of the respective Yang-Mills functional is zero as well.
2.2.3. Yang-Mills fields on reflexive sheaves
If ${\mathscr{G}}$ is a reflexive sheaf on $Y,$ then it is torsion-free and the codimension of the singular set is $\geq 3$ [25, Cap. V, Cor. 5.20], [16, Cor. 1.4]. Furthermore, if $C$ is a closed subset of $Y$ with codimension $\geq 2,$ then the restriction $\Gamma(Y,\,{\mathscr{G}})\to\Gamma(Y\setminus C,\,{\mathscr{G}})$ is an isomorphism [25, Cap.V, Prop. 5.21] [17, Prop. 1.11].
Let us suppose that ${\mathscr{F}}$ is a reflexive sheaf on the Kähler manifold $Y$ endowed with a Hermitian metric. Then ${\mathscr{E}nd}({\mathscr{F}})$ is also a reflexive sheaf [25, Chap. V, Prop. 4.15]. If ${\mathcal{S}}$ is the singular locus of ${\mathscr{F}},$ let ${\mathcal{W}}$ be a closed subspace of $Y$ containing ${\mathcal{S}}$ and such that ${\rm codim}\,{\mathcal{W}}=2.$ Let ${\mathcal{N}}$ denote a tubular neighborhood of ${\mathcal{S}}$ in ${\mathcal{W}}.$ Therefore, the vector space $\Gamma(Y\setminus{\mathcal{N}},\,\Omega^{k}({\mathscr{E}nd}({\mathscr{F}})))\simeq\Gamma(Y,\,\Omega^{k}({\mathscr{E}nd}({\mathscr{F}}))),$ and the latter is finite dimensional. Furthermore, as $(Y\setminus{\mathcal{N}})\cap{\mathcal{S}}=\emptyset,$ ${\mathscr{H}om}({\mathscr{F}},\,\Omega^{k}({\mathscr{F}}))|_{Y\setminus{\mathcal{N}}}$ is a locally free sheaf.
Let $\nabla$ be a holomorphic gauge field on ${\mathscr{F}}$, then
${\mathcal{K}}_{\nabla}\in\Gamma\big{(}Y,\,{\mathscr{H}om}({\mathscr{F}},\,\Omega^{2}({\mathscr{F}}))\big{)}=\Gamma(Y\setminus{\mathcal{N}},\,{\mathscr{H}om}({\mathscr{F}},\,\Omega^{2}({\mathscr{F}}))\big{)}$
can considered as a global section of the locally free sheaf on $Y\setminus{\mathcal{S}}\,$ $\Omega^{2}\otimes{\rm End}(F).$ A similar observation is also valid for $\nabla$.∇ .
Let us assume that the sheaf ${\mathscr{F}}$ supports a holomorphic gauge field $\nabla_{0}$. By Proposition 3, given ${\mathcal{E}}_{1},\dots\mathcal{E}_{m}$, a basis of
${\rm Hom}({\mathscr{F}},\,\Omega^{1}({\mathscr{F}}))=\Gamma\big{(}Y,\,{\mathscr{H}om}({\mathscr{F}},\Omega^{1}({\mathscr{F}}))\big{)}=\Gamma(Y\setminus{\mathcal{S}},\,\Omega^{1}\otimes{\rm End}(F)),$
any holomorphic gauge field can be written $\nabla=\nabla_{0}+\sum\lambda_{i}{\mathcal{E}}_{i},$ with $\lambda_{i}\in{\mathbb{C}}$. The curvature
(2.20) ${\mathcal{K}}_{\nabla}=\nabla\circ\nabla={\mathcal{K}}_{\nabla_{0}}+\sum_{i}\lambda_{i}{\mathcal{B}}_{i}+\sum_{ij}\lambda_{i}\lambda_{j}{\mathcal{B}}_{ij},$
where ${\mathcal{B}}_{i}:=\nabla_{0}({\mathcal{E}}_{i})$ and ${\mathcal{B}}_{ij}:={\mathcal{E}}_{i}\wedge{\mathcal{E}}_{j}$. Thus,
$\lVert{\mathcal{K}}_{\nabla}\rVert^{2}=({\mathcal{K}}_{\nabla},\,{\mathcal{K}}_{\nabla})=P(\lambda_{1},\dots,\lambda_{m}),$
where $P$ is a polynomial of degree $\leq 4$ in the variables $\lambda_{i}$.
The Yang-Mills fields are those $\nabla$ defined by constants $\lambda_{i}$ which satisfy the algebraic equations of degree $\leq 3$
(2.21) $\frac{\partial\,P}{\partial\lambda_{i}}=0,\;\;\;i=1,\dots,m.$
Therefore,
Theorem 1.
If the reflexive sheaf ${\mathscr{F}}$ admits a holomorphic gauge field and
$m={\rm dim}\,{\rm Hom}({\mathscr{F}},\,\Omega^{1}({\mathscr{F}})).$
Then the set ${\sf YM}({\mathscr{F}})$ of holomorphic Yang-Mills fields on ${\mathscr{F}}$ is in bijective correspondence with the points of the algebraic set in ${\mathbb{C}}^{m}$ defined by $m$ algebraic equations.
In particular, if $m=2$ and the cardinal of ${\sf YM}({\mathscr{F}})$ is finite, then #$\#{\sf YM}({\mathscr{F}})\leq 9$.
Proof. The case $m=2$ is a consequence of Bézout’s theorem. ∎
If $\nabla$ is a holomorphic Yang-Mills field, given ${\mathcal{E}}\in{\rm Hom}({\mathscr{F}},\,\Omega^{1}({\mathscr{F}}))$, then $\nabla+{\mathcal{E}}$ is a new connection. Conversely, the difference of two connections on ${\mathscr{F}}$ is a global section of the sheaf ${\mathscr{H}om}({\mathscr{F}},\,\Omega^{1}({\mathscr{F}}))$. Thus, one can consider the following “variation” of $\nabla$, $\nabla_{\epsilon}=\nabla+\epsilon\,{\mathcal{E}},$ with $\epsilon\in{\mathbb{C}}$. Conversely, if $\nabla$ and $\hat{\nabla}$ are holomorphic connections on the Hermitian sheaf ${\mathscr{F}}$, then $\nabla-\hat{\nabla}$ is determined by a vector $\Gamma(Y,\,\Omega^{1}\otimes{\rm End}({F})).$
Considering a variation $\nabla_{\epsilon}$ as above, then $K_{\nabla_{\epsilon}}=K_{\nabla}+\epsilon\nabla E+O(\epsilon^{2})$, where $E$ is the vector of $\Gamma(Y,\,\Omega^{1}\otimes{\rm End}(F)),$ defined by ${\mathcal{E}}$, and $\nabla E$ is the covariant derivative of $E$.
(2.22) $(1/2)\frac{d}{d\epsilon}\Big{|}_{\epsilon=0}||K_{\nabla_{\epsilon}}||^{2}=\int_{Y\setminus{\mathcal{N}}}\langle K_{\nabla},\,\nabla E\rangle\,{\rm dvol}=:(K_{\nabla},\,\nabla{E}).$
Therefore, $\nabla$ is a Yang-Mills field if for any “variation” ${\mathcal{E}}$ of $\nabla$
(2.23) $(K_{\nabla},\,\nabla{E})=0.$
In particular, the flat holomorphic gauge fields are Yang-Mills.
We denote
${}^{(p)}\nabla:\Gamma(Y,\,\Omega^{p}({\mathscr{E}nd}({\mathscr{F}})))\to\Gamma(Y,\,\Omega^{p+1}({\mathscr{E}nd}({\mathscr{F}}))),$
the operator defined by the connection $\nabla$. In this notation Bianchi’s identity is read as
${}^{(2)}\nabla K_{\nabla}=0.$
On the other hand, the orthogonality condition (2.23) which satisfy the Yang-Mills fields gives rise to the following proposition.
Proposition 4.
The holomorphic gauge field $\nabla$ on the reflexive sheaf ${\mathscr{F}}$ is a Yang-Mills field iff its curvature $K_{\nabla}\in\Gamma(Y,\,\Omega^{2}\otimes_{\mathscr{O}}{\rm End}(F))$ is orthogonal to the vector space ${\rm im}\,(^{(1)}\nabla)$.
2.2.4. The case ${\rm rank}\,{\mathscr{F}}=1$.
In this case the reflexive sheaf ${\mathscr{F}}$ is a locally free sheaf [16, Prop. 1.9]. Then ${\mathscr{E}nd}({\mathscr{F}})$ is the sheaf associated to the trivial line bundle ×${\mathbb{C}}\times Y\to Y$. If $s$ is a local frame of the corresponding line bundle $F,$ a holomorphic connection $\nabla$ on ${\mathscr{F}}$ on this frame is determined by a ${\mathbb{C}}$-valued $1$-form $A$, $\nabla s=As$. In this frame ${}^{(p)}\nabla(\beta)=\partial\beta+A\wedge\beta-(-1)^{p}\beta\wedge A=\partial\beta,$ for any ${\mathbb{C}}$-valued $p$-form $\beta$. That is,
(2.24) $^{(p)}\nabla=\partial.$
In this case (2.20) reduces to $K_{\nabla}=K_{\nabla_{0}}$, since $\nabla_{0}(E_{i})=\partial E_{i}=0.$ In particular, the Yang-Mills functional is constant.
On the other hand, the Bianchi’s identity reduces to $\partial K_{\nabla}=0$. If furthermore, $\nabla$ is a Yang-Mills field, from (2.23) it follows $\partial^{\dagger}(K_{\nabla})=0$, where $\partial^{\dagger}$ is the adjoint of $\partial$. That is, $K_{\nabla}$ is $\partial$-harmonic. As $Y$ is a Kähler manifold, $K_{\nabla}$ is also $d$-harmonic. Hence, the norm of $K_{\nabla}$ minimizes the corresponding norm in its cohomology class. That is, denoting by ${\bf c}$ the cohomology class defined by $K_{\nabla},$
(2.25) $\mathcal{YM}(\nabla)=\lVert K_{\nabla}\rVert^{2}={\min}\{\lVert\beta\rVert^{2}\,|\,\beta\in{\bf c}\}.$
As we are assuming that ${\mathscr{F}}$ supports a holomorphic gauge field, the first Chern class $c_{1}({\mathscr{F}})=[(2\pi)^{-1}K_{\nabla}]$ vanishes [4]. Consequently, ${\bf c}=0$ and $\mathcal{YM}$ is the functional zero.
3. Fields on a brane
3.1. Gauge fields on a $B$-brane.
Let $({\mathscr{F}}^{\bullet},\,\delta^{\bullet})$ be a $B$-brane on the complex manifold $Y$; that is, ${\mathscr{F}}^{\bullet}$ is an object of the category $D^{b}(Y)$, the bounded derived category of coherent sheaves over $Y$. According to the observation at the beginning of Subsection 2.1 (see (2.1)), we define a gauge field on the brane ${\mathscr{F}}^{\bullet}$ as an element of ${\rm Hom}_{D^{b}(Y^{(1)})}(L\pi_{1}^{*}{\mathscr{F}}^{\bullet},\,L\pi_{2}^{*}{\mathscr{F}}^{\bullet})$ which lifts the identity on $({\mathscr{F}}^{\bullet},\,\delta^{\bullet})$. We explain below this condition. By the adjunction relation
(3.1) ${\rm Hom}_{D^{b}(Y^{(1)})}(L\pi_{1}^{*}{\mathscr{F}}^{\bullet},\,L\pi_{2}^{*}{\mathscr{F}}^{\bullet})\simeq{\rm Hom}_{D^{b}(Y)}({\mathscr{F}}^{\bullet},\,{\mathscr{J}}^{1}({\mathscr{F}}^{\bullet})),$
where
${\mathscr{J}}^{1}({\mathscr{F}}^{\bullet}):=R\pi_{1*}L\pi_{2}^{*}{\mathscr{F}}^{\bullet}\simeq{\mathscr{O}}_{Y^{(1)}}\otimes^{L}{\mathscr{F}}^{\bullet}.$
Therefore, the gauge fields on ${\mathscr{F}}^{\bullet}$ are elements of the group
${\rm Ext}^{0}({\mathscr{F}}^{\bullet},\,{\mathscr{J}}^{1}({\mathscr{F}}^{\bullet}));$
i.e., open strings between ${\mathscr{F}}^{\bullet}$ and ${\mathscr{J}}^{1}({\mathscr{F}}^{\bullet})$ with ghost number 00 [2, Sect. 5.2], [24].
As ${\mathscr{O}}_{Y^{(1)}}$ is the locally free module ${\mathscr{O}}\oplus\Omega^{1}$, then ${\mathscr{J}}^{1}({\mathscr{F}}^{\bullet})$ is the complex of abelian sheaves $\sf{Forg}\big{(}{\mathscr{F}}^{\bullet}\oplus\Omega^{1}({\mathscr{F}}^{\bullet})\big{)},$ with ${\mathscr{O}}$-structure given by
(3.2) $f\ast(\sigma^{\bullet}\oplus\alpha^{\bullet})=f\sigma^{\bullet}\oplus(df\otimes\sigma^{\bullet}+f\alpha^{\bullet}).$
More precisely, one can consider the complex of abelian sheaves direct sum of ${\mathscr{F}}^{\bullet}$ and $\Omega^{1}({\mathscr{F}}^{\bullet}).$ According to the notation introduced at the end of Subsection 2.1.2, we write
(3.3) $\Big{(}\boldsymbol{\mathscr{J}}^{1}({\mathscr{F}}^{\bullet}):={\sf Forg}\big{(}{\mathscr{F}}^{\bullet}\oplus\Omega^{1}({\mathscr{F}}^{\bullet})\big{)},\,\,\delta_{J}^{\bullet}\,\Big{)},$
where $\delta_{J}^{\bullet}=\delta^{\bullet}\oplus(1\otimes\delta^{\bullet}).$ The complex $\boldsymbol{\mathscr{J}}^{1}({\mathscr{F}}^{\bullet})$ can be equipped with an ${\mathscr{O}}$-module structure (3.2). As $\delta$ is the coboundary operator of a complex ${\mathscr{O}}$-modules, then $\delta_{J}(f\ast(-))=f\ast(\delta_{J}(-)).$ Then ${\mathscr{J}}^{1}({\mathscr{F}}^{\bullet})$ the complex (3.3) endowed with this ${\mathscr{O}}$-module structure.
The exact sequence of complexes of ${\mathscr{O}}$-modules
[…]
determines a distinguished triangle
[…]
in the category $D^{b}(Y)$ [22, page 46], [11, page 157]. As ${\rm Hom}_{D^{b}(Y)}({\mathscr{F}}^{\bullet},\,.\,)$ is a cohomological functor, it follows that
${\rm Hom}_{D^{b}(Y)}({\mathscr{F}}^{\bullet},\,\Omega^{1}({\mathscr{F}}^{\bullet}))\to{\rm Hom}_{D^{b}(Y)}({\mathscr{F}}^{\bullet},\,{\mathscr{J}}^{1}({\mathscr{F}}^{\bullet}))\overset{\pi\circ}{\longrightarrow}{\rm Hom}_{D^{b}(Y)}({\mathscr{F}}^{\bullet},\,{\mathscr{F}}^{\bullet})$
$\to{\rm Ext}^{1}({\mathscr{F}}^{\bullet},\,\Omega^{1}({\mathscr{F}}^{\bullet}))\to$
is an exact sequence. The Atiyah class of ${\mathscr{F}}^{\bullet}$ is the image of $1\in{\rm Hom}_{D^{b}(Y)}({\mathscr{F}}^{\bullet},\,{\mathscr{F}}^{\bullet})$ in ${\rm Ext}^{1}({\mathscr{F}}^{\bullet},\,\Omega^{1}({\mathscr{F}}^{\bullet}))$. Thus, we give the following definition.
Definition 3.
A gauge field on ${\mathscr{F}}^{\bullet}$ is an element $\psi\in{\rm Hom}_{D^{b}(Y)}({\mathscr{F}}^{\bullet},\,{\mathscr{J}}^{1}({\mathscr{F}}^{\bullet}))$, such that ${\pi\circ}\psi=1\in{\rm Hom}_{D^{b}(Y)}({\mathscr{F}}^{\bullet},\,{\mathscr{F}}^{\bullet})$
From the above exact sequence, it follows the following proposition.
Proposition 5.
The vanishing of the Atiyah class of ${\mathscr{F}}^{\bullet}$ is a necessary and sufficient condition for the existence of gauge fields on this brane. Furthermore, the set of gauge fields on ${\mathscr{F}}^{\bullet}$, if is nonempty, is an affine space over the finite dimensional vector space ${\rm Ext}^{0}({\mathscr{F}}^{\bullet},\,\Omega^{1}({\mathscr{F}}^{\bullet}))$.
Remark 4.
The complex ${\mathscr{J}}^{1}({\mathscr{F}}^{\bullet})$ gives rise to the corresponding cohomology ${\mathscr{O}}$-modules ${\mathscr{H}}^{*}({\mathscr{J}}^{1}({\mathscr{F}}^{\bullet})).$
On the other hand, the $B$-brane $({\mathscr{F}},\,{\delta}^{\bullet})$ determines the “complex” ((), 0)superscriptsuperscript $\big{(}{\mathscr{H}}^{*}({\mathscr{F}}^{\bullet}),\,0\big{)}$ consisting of the cohomology ${\mathscr{O}}$-modules. In turn, we can consider ${\mathscr{J}}^{1}({\mathscr{H}}^{*}({\mathscr{F}}^{\bullet})),$ the jet sheaf complex corresponding to this “complex”. It is immediate to check that ${\mathscr{H}}^{*}({\mathscr{J}}^{1}({\mathscr{F}}^{\bullet}))$ and ${\mathscr{J}}^{1}({\mathscr{H}}^{*}({\mathscr{F}}^{\bullet}))$ are canonically isomorphic
(3.5) ${\mathscr{J}}^{1}({\mathscr{H}}^{*}({\mathscr{F}}^{\bullet}))\simeq{\mathscr{H}}^{*}({\mathscr{J}}^{1}({\mathscr{F}}^{\bullet})).$
3.2. Gauge fields on $B$-branes over ${\mathbb{P}}^{n}$.
In this subsection we construct a family of generators for the derived category $D^{b}({\mathbb{P}}^{n}),$ and prove that the set of holomorphic gauge fields on any $B$-brane over ${\mathbb{P}}^{n}$ has at most $1$ element. Some general properties that will be used later are compiled beforehand.
3.2.1. Generators of a derived category
Let ${\mathbf{C}}$ be an abelian category and let ${\sf E}$ be a finite set of generators of the derived category $D^{b}({\mathbf{C}}).$ Given $A$ a complex in ${\mathbf{C}}$, as usual, we denote with ${A}[l]$, with $l\in{\mathbb{Z}}$, the complex ${A}$ shifted $l$ to the left.
Given $E,\tilde{E}$ elements of the generating set ${\sf E},$ let us consider morphisms $h$ between ${E}^{\prime}:={E}[l]$ and ${\tilde{E}}^{\prime}:={\tilde{E}}[l^{\prime}]$. We denote by ${\rm Cone}(h)={E}^{\prime}[1]\oplus{\tilde{E}}^{\prime}$ the mapping cone of $h$ [11, page 154]. We define ${\sf E}^{(1)}$ the set obtained adding to ${\sf E}$ the elements of the form ${\rm Cone}(h)$. Hence, an element of ${\sf E}^{(1)}$ is a complex whose term at a position $p$ is either 00, or $E$, or a direct sum of ${E}\oplus\tilde{E},$ with $E,\tilde{E}\in{\sf E}.$
Repeating the process with the elements of ${\sf E}^{(1)}$ one obtains ${\sf E}^{(2)}$, etc. The objects of the triangulated subcategory generated by the family ${\sf E}$ are elements which belong to some ${\sf E}^{(m)}$. Therefore, an object of the triangulated subcategory of $D^{b}({\mathbf{C}})$ generated by ${\sf E}$ is a complex $(G,\,d)$, where ${G}^{p}$ is of the form
(3.6) $G^{p}=\bigoplus_{i\in I_{p}}E_{pi},$
with $E_{pi}\in{\sf E}$ and $i$ varying in a finite set $I_{p}$. (When $i$ “runs over the empty set”, the direct sum is taken to be 00). As ${\sf E}$ is a set generating $D^{b}({\mathbf{C}}),$ each object of this category is isomorphic to one of the form (3.6).
Given two complexes $(A,\,\partial_{A})$ and $(B,\,\partial_{B})$ in an additive category ${\mathbf{C}}$, the complex ${\rm Hom}^{\bullet}(A,\,B)$ is defined by (see [21, page 17])
(3.7) ${\rm Hom}^{m}(A,\,B)=\prod_{q\in{\mathbb{Z}}}{\rm Hom}_{\mathbf{C}}\big{(}A^{q},\,B^{q+m}\big{)},$
with the differential $\delta_{H}$.
(3.8) $(\delta_{H}^{m}g)^{q}=\partial_{B}^{m+q}g^{q}+(-1)^{m+1}g^{q+1}\partial_{A}^{q},$
with $g\in{\rm Hom}^{m}(A,\,B).$
Denoting by ${\rm Com}({\mathbf{C}})$ the abelian category of complexes on ${\mathbf{C}}$ and fixed the object $A,$ one has the functor $F_{A}:{\rm Com}(\mathbf{C})\to{\rm Com}(\mathbf{Ab})$ defined by
$F_{A}(B)={\rm Hom}^{\bullet}(A,\,B),\;\;F_{A}(f)^{m}(g)=\big{(}f^{m+q}g^{q}\big{)},$
$f$ a morphism from $(B,\,\partial_{B})\to(C,\,\partial_{C}).$ Assumed that ${\mathbf{C}}$ has enough injectives, as $F_{A}$ is a left exact functor, one has the corresponding derived functor $RF_{A}:D^{+}{(\mathbf{C}})\to D^{+}(\mathbf{Ab}).$
Since the 00th derived functor of a left exact functor is isomorphic to the functor itself, $F_{A}=R^{0}F_{A}={\rm Ext}^{0}(A,\,.\,).$ On the other hand, ${\rm Hom}_{D({\mathbf{C}})}\big{(}{A},\,B\big{)}={\rm Ext}^{0}(A,\,{B})$ [31, Sect. 10.7], so
(3.9) ${\rm Hom}_{D({\mathbf{C}})}\big{(}{A},\,B\big{)}={\rm Hom}^{\bullet}({A},\,{B}).$
One has the following lemma.
Lemma 1.
Let ${G}$ be the complex in the category ${\mathbf{C}}$ defined in (3.6) and $\hat{G}$ the complex ${\hat{G}}^{q}=\bigoplus_{j\in J_{q}}E_{qj}.$ If ${\rm Hom}_{\mathbf{C}}\big{(}E_{pi},\,E_{qj}\big{)}=0$ for all $p,q,i,j,$ then
${\rm Hom}_{D^{b}({\mathbf{C}})}\big{(}{G},\,\hat{G}\big{)}=0.$
Proof. From (3.7) and (3.9) together with the additivity of the functor ${\rm Hom}_{\mathbf{C}}(\,.\,,\,.\,),$ it follows
${\rm Hom}^{m}(G,\,\hat{G})=\prod_{p}\bigoplus_{ij}{\rm Hom}_{\mathbf{C}}(E_{pi},\,E_{(p+m)j})=0.$
Thus, ${\rm Hom}_{D^{b}({\mathbf{C}})}({G},\,\hat{G})=0.$
We shall deal with the particular case where the category ${\mathbf{C}}$ is a category $\mathbf{Coh}(X)$ of coherent sheaves. More precisely, let $X$ be a smooth complex $n$-dimensional variety, let $\mathbf{Qcoh}(X)$ denote the category of quasi-coherent sheaves on $X.$ One has the derived functor
$RF_{A}:D^{+}(\mathbf{Qcoh}(X))\to D^{+}(\mathbf{Ab}).$
Since ${\rm Ext}^{i}(A,\,B)=0$ for $A,B$ complexes of coherent sheaves and $i\notin[0,\,n],$ we have the functor
$RF_{A}:D^{b}(X):=D^{b}(\mathbf{Coh}(X))\to D^{b}(\mathbf{Ab}).$
From Lemma 1 applied to the category $D^{b}(X)$, it follows the following corollary.
Corollary 1.
Let ${\mathscr{G}}^{p}=\bigoplus_{pi}{\mathscr{E}}_{pi}$ be an object of $D^{b}(X)$ and $\hat{\mathscr{G}}$ is an object of this category isomorphic the complex ${\mathscr{G}}^{p}\otimes_{{\mathscr{O}}_{X}}\Omega^{1}_{X},$ then
${\rm Hom}_{D^{b}(X)}\big{(}{{\mathscr{G}}},\,\hat{\mathscr{G}}\big{)}=0,$
if ${\rm Hom}({\mathscr{E}}_{ip},\,\Omega^{1}_{X}({\mathscr{E}}_{jq}))=0$ for all $p,q,i,j.$
3.2.2. Generators of the derived category $D^{b}(X)$
Let $X$ be a locally Noetherian scheme, $i:Z\hookrightarrow X$ a closed subscheme of $X$ and $j:U\hookrightarrow X$ an open part of $X.$ Since $j$ is an open embedding, the functor proper inverse image $j^{!}$ is defined between the categories of modules $\mathbf{Mod}({\mathscr{O}}_{X})\to\mathbf{Mod}({\mathscr{O}}_{U}).$ One has the pair of adjoint functors $(j_{!}\dashv j^{!}=j^{*})$ and the pair $(i^{*}\dashv i_{*}=i_{!}).$
$j^{!}:\mathbf{Mod}({\mathscr{O}}_{X})\leftrightarrows\mathbf{Mod}({\mathscr{O}}_{U}):j_{!}\;\;\;\text{and}\;\;\;i^{*}:\mathbf{Mod}({\mathscr{O}}_{X})\leftrightarrows\mathbf{Mod}({\mathscr{O}}_{Z}):i_{*}.$
For every sheaf ${\mathscr{F}}$ on $X,$ the corresponding adjunction morphisms give rise to the sequence of morphisms of ${\mathscr{O}}_{X}$-modules
(3.10) $j_{!}j^{!}{\mathscr{F}}\to{\mathscr{F}}\to i_{*}i^{*}{\mathscr{F}}.$
If ${\mathscr{F}}$ is a coherent ${\mathscr{O}}_{X}$-module, then sheaf $i^{*}{\mathscr{F}}$ is coherent. Since $i$ is a proper map, $i_{*}i^{*}{\mathscr{F}}$ is a coherent ${\mathscr{O}}_{X}$-module. As $j_{!}j^{!}{\mathscr{F}}$ is a subsheaf of ${\mathscr{F}}$, it is also coherent.
In particular, when $U=X\setminus Z$, for any ${\mathscr{F}}\in\mathbf{Coh}(X)$, one has the exact sequence of ${\mathscr{O}}_{X}$-modules
(3.11) $0\to j_{!}j^{!}{\mathscr{F}}\to{\mathscr{F}}\to i_{*}i^{*}{\mathscr{F}}\to 0$
If ${\mathscr{F}}^{\bullet}$ is an object of the derived category $D^{b}(X)=D^{b}(\mathbf{Coh}(X))$, then we have the distinguished triangle
(3.12) $Rj_{!}j^{!}{\mathscr{F}}^{\bullet}\to{\mathscr{F}}^{\bullet}\to Ri_{*}i^{*}{\mathscr{F}}^{\bullet}\overset{+1}{\longrightarrow},$
where $j^{!}{\mathscr{F}}^{\bullet}\in D^{b}(\mathbf{Coh}(U))$ and $i^{*}{\mathscr{F}}^{\bullet}\in D^{b}(\mathbf{Coh}(Z)).$ Hence, as $j_{!}$ and $i_{*}$ are exact functors, one has the following proposition.
Proposition 6.
If $\{g_{k}\}$ (resp. $\{g^{\prime}_{k^{\prime}}\}$) is a set generators of the derived category $D^{b}(\mathbf{Coh}(U))$ (resp. $D^{b}(\mathbf{Coh}(Z))$) as a triangulated category, then $\{j_{!}g_{k},\,i_{*}g^{\prime}_{k^{\prime}}\}$ is a set that generates $D^{b}(X).$
If $U={\rm Spec}\,R$, where $R$ is is a Noetherian ring, the category $\mathbf{Coh}(U)$ of coherent sheaves on $U$ is equivalent to the category of finitely generated $R$-modules, with $R=\Gamma(U,\,{\mathscr{O}}_{U}).$ The functor $\Gamma(U,\,\cdot\,)$ defines the equivalence.
Moreover, if $R$ is a regular local ring, then its global dimension is finite, and any object $N$ of the category $\mathbf{Mod}_{f}(R)$ (of the $R$-modules of finite type) admits a finite free resolution
(3.13) $0\to S_{m}\overset{f_{m}}{\longrightarrow}S_{m-1}\to\dots\to S_{1}\overset{f_{1}}{\longrightarrow}S_{0}\to N\to 0,$
with $S_{a}$ a finite direct sum $\oplus_{I_{a}}R.$
Denoting by $\mathbf{T}$ the triangulated subcategory of $D^{b}(\mathbf{Mod}_{f}(R))$ generated by $R,$ then the $S_{a}$ are objects of $\mathbf{T}.$ On the other hand, if we have the exact sequence $0\to S_{1}\overset{f_{1}}{\longrightarrow}S_{0}\to N\to 0$ in the category $\mathbf{Mod}_{f}(R),$ then one has the distinguished triangle
$S_{1}\overset{f_{1}}{\longrightarrow}S_{0}\to{\rm Cone}(f_{1})\overset{+1}{\longrightarrow}.$
Thus, $N$ belongs to the category $\mathbf{T}.$ In the case that $N$ has a resolution as (3.13), then one can consider the exact sequences
$0\to S_{m}\overset{f_{m}}{\longrightarrow}S_{m-1}\to{\rm im}(f_{m-1})\to 0,$
and
$0\to{\rm im}(f_{m-1})\to S_{m-2}\to\dots\to N\to 0$
By the previous result, ${\rm im}(f_{m-1})$ belongs to $\mathbf{T}.$ By induction, it follows that $N$ is also an object of $\mathbf{T}.$ Hence $\mathbf{Mod}_{f}(R)\subset{\mathbf{T}}.$ As $D^{b}(\mathbf{Mod}_{f}(R))$ is generated by $\mathbf{Mod}_{f}(R),$ it follows that $R$ generates the derived category $D^{b}(\mathbf{Mod}_{f}(R)).$
From the equivalence between $\mathbf{Coh}(U)$ and $\mathbf{Mod}_{f}(R),$ one deduces the following proposition.
Proposition 7.
If $U$ is a Noetherian affine scheme, then ${\mathscr{O}}_{U}$ generates $D^{b}(U)$, the bounded derived category of coherent sheaves on $U$.
From Proposition 6, it follows the following corollary (see [5, Lem. 2.3.2]).
Corollary 2.
[…]
An Example. Let $X$ be ${\mathbb{P}}^{1},$ $U=\{[1:x]\in{\mathbb{P}}^{1}\,|\,x\in{\mathbb{C}}\}$ and $Z=\{[0:1]\}.$ In this case, ${\mathscr{O}}_{Z}=i_{*}i^{*}{\mathscr{O}}_{X}$ is the skyscraper sheaf at the point $P=[0:1],$ and $j_{!}j^{!}{\mathscr{O}}_{X}={\mathscr{O}}_{X}(-1).$ That is, $j_{!}{\mathscr{O}}_{U}\simeq{\mathscr{O}}_{X}(-1).$
As $U={\rm Spec}\,R$ with $R$ the polynomial ring ${\mathbb{C}}[x]$, then ${\mathscr{O}}_{U}$ is the generator of the corresponding category $D^{b}(U).$ The category $D^{b}(Z)$ is generated by $i^{*}{\mathscr{O}}_{X}$, obviously. By Corollary 2), $\{{\mathscr{O}}_{X}(-1),\,{\mathscr{O}}_{Z}\}$ is a set of generators of $D^{b}(X).$
In this case, the exact sequence (3.11) is
$0\to{\mathscr{O}}_{X}(-1)\to{\mathscr{O}}_{X}\to{\mathscr{O}}_{Z}\to 0$
Therefore, the set $\{{\mathscr{O}}_{X}(-1),\,{\mathscr{O}}_{X}\}$ is also a set of generators of $D^{b}(X)$ (which is in agreement with Beilinson’s theorem). Hence, any object of $D^{b}(X)$ is isomorphic to one complex ${\mathscr{G}}^{\bullet}$ of the form (3.6), where ${\mathscr{E}}_{pi}\in\{{\mathscr{O}_{X}},\,{\mathscr{O}}_{X}(-1)\}.$
On the other hand, for ,${\mathscr{E}},\,{\mathscr{E}}^{\prime}\in\{{\mathscr{O}_{X}},\,{\mathscr{O}}_{X}(-1)\}$, one has
${\rm Hom}({\mathscr{E}},\,{\mathscr{E}}^{\prime}\otimes_{\mathscr{O}}\Omega^{1}_{X})={\rm Hom}({\mathscr{O}}_{X},\,({{\mathscr{E}}}^{\vee}\otimes{\mathscr{E}}^{\prime})\otimes_{\mathscr{O}}\Omega^{1}_{X})=H^{0}(X,\,\Omega^{1}(\epsilon)),$
where $\epsilon\in\{0,+1,-1\}$ and $\Omega^{1}(\epsilon)$ is the twisted sheaf $\Omega^{1}_{X}\otimes_{{\mathscr{O}}_{X}}{\mathscr{O}}(\epsilon).$ But $H^{0}(X,\,\Omega^{1}(\epsilon))=0$ [28, page 4]. Thus, by Corollary 1 together with Proposition 5, the set of of holomorphic gauge fields over any brane on ${\mathbb{P}}^{1}$ has cardinal $\leq 1.$
Remark 5.
The family ${\sf E}=\{{\mathscr{O}}_{{\mathbb{P}}^{n}}(-n),\,{\mathscr{O}}_{{\mathbb{P}}^{n}}(-n+1),\dots,{\mathscr{O}}_{{\mathbb{P}}^{n}}\}$ generates $D^{b}({\mathbb{P}}^{n}).$ If ${\mathscr{E}},\,{\mathscr{E}}^{\prime}\in{\sf E},$ then
${\rm Hom}\big{(}{\mathscr{E}},\,\Omega^{1}_{{\mathbb{P}}^{n}}({\mathscr{E}}^{\prime})\big{)}=H^{0}\big{(}X,\,\Omega^{1}_{{\mathbb{P}}^{n}}(\epsilon)\big{)},$
with $\epsilon\in\{-n,\,-n+1,\dots,n\}.$
If $n>1,$ the cohomology group $H^{0}(X,\,\Omega^{1}_{{\mathbb{P}}^{n}}(\epsilon))$ is not necessarily zero (see [28, page 4]). Thus, Corollary 1 is not applicable in order to extend to ${\mathbb{P}}^{n}$ the above result obtained for the $B$-branes over ${\mathbb{P}}^{1}.$ To achieve such an extension it will be necessary to construct another set of generators.
3.2.3. Generators of $D^{b}({\mathbb{P}}^{n})$
We will generalize the result of the above Example. Let $X$ be a smooth a complex $n$-dimensional variety. Let $Z_{1}$ be an irreducible subvariety of $X$ of codimension $1,$ and $i_{1}:Z_{1}\hookrightarrow X$ the corresponding inclusion. We set $U_{1}:=X\setminus Z_{1}$ and $j_{1}:U_{1}\hookrightarrow X.$ Let us assume that $U_{1}$ is isomorphic to an affine variety. The exact sequence
$0\to j_{1!}j_{1}^{!}{\mathscr{O}}_{X}\to{\mathscr{O}}_{X}\to i_{1*}i^{*}_{1}{\mathscr{O}}_{X}\to 0,$
can be written $0\to{\mathscr{M}}_{1}\to{\mathscr{O}}_{X}\to{\mathscr{O}}_{Z_{1}}\to 0,$ where ${\mathscr{M}}_{1}$ is the invertible sheaf ${\mathscr{O}}_{X}(-Z_{1}),$ which is supported in $U_{1},$ i.e. the stalk $({\mathscr{M}}_{1})_{x}=0$ for all $x\in Z_{1}.$
As $j_{1}^{!}{\mathscr{O}}_{X}={\mathscr{O}}_{U_{1}}$, then ${\mathscr{M}}_{1}=j_{1!}{\mathscr{O}}_{U_{1}}$. From Corollary 2, it follows that the derived category $D^{b}(X)$ is generated by ${\mathscr{M}}_{1}$ together with the elements of $i_{1*}\{{\rm generators\;of}\,D^{b}(Z_{1})\}.$
If $Z_{2}\overset{i_{2}}{\hookrightarrow}Z_{1}$ is an irreducible hypersurface of $Z_{1}$ and $U_{2}:=Z_{1}\setminus Z_{2}$ is isomorphic to an affine variety, then we have an exact sequence
$0\to{\mathscr{M}}_{2}\to{\mathscr{O}}_{Z_{1}}\to i_{2*}i_{2}^{*}{\mathscr{O}}_{Z_{1}}\to 0,$
where ${\mathscr{M}}_{2}$ is the sheaf on $Z_{1},$ ${\mathscr{O}}_{{Z}_{1}}(-Z_{2}).$ Hence, $D^{b}(Z_{1})$ is generated by ${\mathscr{M}}_{2}$ together with $i_{2*}\{{\rm generators\,of}\,D^{b}(Z_{2})\}.$ Consequently, $D^{b}(X)$ is generated by
$\{{\mathscr{M}}_{1},\,i_{1*}{\mathscr{M}}_{2}\}\cup i_{1*}i_{2*}\big{(}\{{\rm generators\,of}\,D^{b}(Z_{2})\}\big{)}.$
Let us assume that there exists a tower $X=Z_{0}\supset Z_{1}\supset Z_{2}\supset\dots\supset Z_{n},$ where $Z_{k}\overset{i_{k}}{\hookrightarrow}Z_{k-1}$ is an irreducible hypersurface of $Z_{k-1}$ with $i=1,\dots,n.$ According to the above argument, ${\mathscr{M}}_{k}={\mathscr{O}}_{Z_{k-1}}(-Z_{k})$ is a sheaf supported on $U_{k}:=Z_{k-1}\setminus Z_{k}$. We set $\iota_{k}=i_{1}\circ\dots\circ i_{k}:Z_{k}\hookrightarrow X$ and $\iota_{0}:={\rm id}:Z_{0}\to X.$ For $k=1,\dots,n$ we denote ${\mathscr{L}}_{k}:=\iota_{k-1*}({\mathscr{M}}_{k}).$
We set ${\mathscr{L}}_{n+1}$ for denoting the sheaf on $X$ direct sum of the skyscraper sheaves on the points of the finite set $Z_{n}.$ Then
${\mathscr{L}}_{n+1}=\iota_{n*}\big{(}\{{\rm generators\,of}\,D^{b}(Z_{n})\}).$
We have proved the following proposition.
Proposition 8.
With the above notations, the derived category $D^{b}(X)$ is generated by the set of ${\mathscr{O}}_{X}$-modules
${\sf E}=\{{\mathscr{L}}_{1},\dots,{\mathscr{L}}_{n},\,{\mathscr{L}}_{n+1}\}.$
  • When $j\neq k$ the $U_{j}\cap U_{k}=\emptyset.$ Hence, ${\mathscr{L}}_{j}$ and $\Omega^{1}({\mathscr{L}}_{k})$ have disjoint supports.; thus,
    (3.14) ${\rm Hom}_{O_{X}}\big{(}{\mathscr{L}}_{j},\,\Omega^{1}_{X}\otimes_{{\mathscr{O}}_{X}}{\mathscr{L}}_{k}\big{)}=0.$
  • Since ${\mathscr{L}}_{1}$ is the invertible ${\mathscr{O}}_{X}(-Z_{1}),$ tensoring by the dual ${\mathscr{L}}_{1}^{\vee}$
    ${\rm Hom}({\mathscr{L}}_{1},\,{\mathscr{L}}_{1}\otimes_{{\mathscr{O}}_{X}}\Omega^{1}_{X})={\rm Hom}({\mathscr{O}}_{X},\,\Omega^{1}_{X})=H^{0}(X,\,\Omega^{1}_{X})=H^{1,0}(X).$
  • In this item, when $k>1,$ we set ${\mathscr{L}}:={\mathscr{L}}_{k},$ ${\mathscr{M}}:={\mathscr{M}}_{k},$ and $\iota:=\iota_{k}$ for the inclusion $Z:=Z_{k}\hookrightarrow X;$ thus ${\mathscr{L}}=\iota_{*}{\mathscr{M}}.$ We will also delete the subscript $X$ in the sheaves ${\mathscr{O}}_{X}=:{\mathscr{O}}$ and $\Omega^{1}_{X}=:\Omega^{1}.$ By ${\mathscr{H}om}\,_{\mathscr{R}}(\,.\,,\,.\,)$ we denote the sheaf of linear homomorphisms between the corresponding ${\mathscr{R}}$-modules.
    By [22, Prop. 2.5.13] and [22, Cor. 2.3.4]
    ${\mathscr{H}om}\,_{\mathscr{O}}({\mathscr{L}},\,{\mathscr{L}}\otimes_{\mathscr{O}}\Omega^{1})\simeq\iota_{*}{\mathscr{H}om}\,_{{\mathscr{O}}_{Z}}({\mathscr{M}},\,{\mathscr{M}}\otimes_{{\mathscr{O}}_{Z}}\Omega^{1}_{Z}).$
    Hence,
    (3.15) ${\mathscr{H}om}\,_{\mathscr{O}}({\mathscr{L}},\,{\mathscr{L}}\otimes_{\mathscr{O}}\Omega^{1})\simeq\Gamma\big{(}X,\,\iota_{*}{\mathscr{H}om}\,_{{\mathscr{O}}_{Z}}({\mathscr{M}},\,{\mathscr{M}}\otimes_{{\mathscr{O}}_{Z}}\Omega^{1}_{Z})\big{)}=$
    $\Gamma\big{(}Z,\,{\mathscr{H}om}\,_{{\mathscr{O}}_{Z}}({\mathscr{M}},\,{\mathscr{M}}\otimes_{{\mathscr{O}}_{Z}}\Omega^{1}_{Z})\big{)}\simeq{\rm Hom}_{{\mathscr{O}}_{Z}}({\mathscr{O}}_{Z},\Omega^{1}_{Z})=H^{1,0}(Z).$
From the above items, together with Corollary 1, Proposition 8 and Proposition 5, it follows the theorem.
Theorem 2.
Let $X$ be an $n$-dimensional smooth variety and suppose that
  1. (1)
    There exists $X=Z_{0}\supset Z_{1}\supset\dots\supset Z_{n},$ where $Z_{i+1}$ is a smooth irreducible subvariety of $Z_{i},$ with ${\rm codim}_{Z_{i}}Z_{i+1}=1$.
  2. (2)
    $U_{j}=Z_{j-1}\setminus Z_{i}$ is isomorphic to an affine variety.
If $H^{1,0}(Z_{j})=0$ for all $j,$ then the set of holomorphic gauge fields on any $B$-brane over $X$ has cardinal $\leq 1.$
3.2.4. Holomorphic gauge fields on $B$-branes over ${\mathbb{P}}^{n}$
Let $X={\mathbb{P}}^{n}$, we define
$Z_{1}:=\{[0:x_{1}:\dots:x_{n}]\simeq{\mathbb{P}}^{n-1},\;U_{1}={\mathbb{P}}^{n}\setminus Z_{1}\simeq{\mathbb{C}}^{n}.$
Analogously, one defines $Z_{2}=\{[0:0:x_{2}:\dots:x_{n}]\}$ and $U_{2}=Z_{1}\setminus Z_{2},$ etc. Moreover, $H^{1,0}(Z_{j})=0.$ By Theorem 2, one has the following theorem.
Theorem 3.
The set of holomorphic gauge fields on any $B$-brane on ${\mathbb{P}}^{n}$ has cardinal $\leq 1.$
Remark 6.
For the $B$-brane on $X={\mathbb{P}}^{n}$ consisting only of the line bundle ${\mathscr{O}}(k),$
${\rm Hom}({\mathscr{O}}(k),\,\Omega^{1}_{X}\otimes_{{\mathscr{O}}_{X}}{\mathscr{O}}(k))={\rm Hom}({\mathscr{O}},\,\Omega^{1}_{X})=H^{0}(X,\,\Omega^{1}_{X})=0.$
This fact is in agreement with the previous general result.
On the other hand, if a line bundle admits a holomorphic gauge field then its Chern class vanishes [4]. Thus, among the elements of the set $\{{\mathscr{O}}_{X}(k)\}_{k\in{\mathbb{Z}}}$, only ${\mathscr{O}}_{X}$ can support a holomorphic gauge field, and this field is precisely $\partial:{\mathscr{O}}_{X}\to\Omega^{1}_{X}.$
3.3. Flag varieties
In this section, we will first recall some properties of the Schubert varieties. We denote by $X:={\rm GL}(n,\,{\mathbb{C}})/B$, where $B$ is the Borel subgroup of ${\rm GL}(n,\,{\mathbb{C}})$ consisting of the upper triangular matrices. $X$ is the manifold of complete flags in ${\mathbb{C}}^{n}.$
Let $T$ denote the maximal torus of ${\rm GL}(n,\,{\mathbb{C}})$ formed by the diagonal matrices. We set $W:=N(T)/T$, where $N(T)$ is the normalizer of $T$ in the group ${\rm GL}(n,\,{\mathbb{C}}).$ As it is well-known, the Weyl group $W$ is isomorphic to the symmetric group $S_{n}$. The Bruhat partial order [9, Sect 10.5] in $S_{n}$ will be denoted by $\leq$.
Given $w\in W,$ we set $C^{0}_{w}$ for the Schubert cell of $X$ associated to $w$ [9]. The closure (in Zariski topology) of this cell is the corresponding Schubert subvariety of $X$
(3.16) $C_{w}:=\bigcup_{v\leq w}C_{v}^{0}.$
The Poincaré polynomial of $C_{w}$ given in [1] shows that the Betti number $b_{1}(C_{w})$ vanishes, if $C_{w}$ is any smooth Schubert variety.
Theorem 1 of [26] gives a criterion to determine whether a Schubert variety in a flag variety is singular. Let $w$ be the permutation $w=(t_{1},\dots,t_{n})$. The variety $C_{w}$ is singular iff there are integers $i,j,k,l$ with $1\leq i<j<k<l\leq n$ such that, either
$t_{k}<t_{l}<t_{i}<t_{j},\;\;\text{or}\;\;t_{l}<t_{j}<t_{k}<t_{i}.$
In particular, all the Schubert varieties of $X={\rm GL}(3,\,{\mathbb{C}})/B$ are smooth. Thus, if $C_{w}$ is such a variety, as the $b_{1}(C_{w})=0$, the Hodge numbers $h^{1,0}(C_{w})$ and $h^{0,1}(C_{w})$ vanish
(3.17) $h^{1,0}(C_{w})=h^{0,1}(C_{w})=0.$
From now on in this section, $X$ will denote the variety of complete flags in ${\mathbb{C}}^{3}.$ The Bruhat order in the group $S_{3}$ is showed in the following diagram [6].
${w_{1}=(321)}$${w_{3}=(231)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${w_{2}=(312)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${w_{5}=(213)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${w_{4}=(132)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${w_{6}=(123)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
Hence, there are only two pairs of permutations that are not comparable in the Bruhat order, namely, $(w_{2},\,w_{3})$ and $(w_{4},\,w_{5}).$ For any other pair $(w,\,w^{\prime})$ the intersection $C_{w}\cap C_{w^{\prime}}=C_{v}$, where $v={\rm min}\{w,\,w^{\prime}\}.$
The length $l(w)$ of the permutation $w$ is defined by
$l(w)=\#\{a<b\,|\,w(a)>w(b)\}.$
We will set $C^{0}_{k}$ for the cell $C^{0}_{w_{k}}$ and $C_{k}:=C_{w_{k}}.$ The cell $C^{0}_{w}$ is an affine variety isomorphic to ${\mathbb{C}}^{l(\omega)}$ [9, page 157]. Hence, ${\rm dim}\,C_{4}=1={\rm dim}\,C_{5}$ and ${\rm dim}\,C_{2}=2={\rm dim}\,C_{3}.$ Furthermore, $X$ is the disjoint union of the cells
$X=\bigsqcup_{k=1}^{6}C^{0}_{k}.$
On the other hand, the inclusion relations among the Schubert subvarieties of $X$ can be obtained from (3.16) by means of the Bruhat order depicted above.
We consider the following tower of subspaces of $X$
$Z_{0}:=X\supset Z_{1}:=C_{2}\cup C_{3}\supset Z_{2}:=C_{4}\cup C_{5}\supset Z_{3}:=C_{6}.$
That is, $Z_{i}$ is the union of the Schubert varieties of dimension $3-i.$ We set $U_{i}:=Z_{i-1}\setminus Z_{i}$. From (3.16), it follows the expression of each $U_{i}$ in terms of Schubert cells. The corresponding inclusion relations are showed in the diagram.
[…]
A family of generators for $D^{b}(X)$ will then be constructed by successive applications of Corollary 2.
$U_{1}$ open part of $Z_{0}=X.$ In the flag variety $X$, we have the divisor $Z_{1}=X\setminus C^{0}_{1}=C_{2}\cup C_{3},$ and the partition of $X$ defined by the affine variety $U_{1}=C^{0}_{1}$ and $Z_{1}.$ We denote by $j_{1}$ and $i_{1}$ the respective inclusions
$U_{1}\overset{j_{1}}{\hookrightarrow}X\overset{i_{1}}{\hookleftarrow}Z_{1}.$
On the other hand, we have the exact sequence of ${\mathscr{O}}_{X}$-modules
$0\to{\mathscr{M}}_{1}:=j_{1!}j^{!}_{1}{\mathscr{O}}_{X}\to{\mathscr{O}}_{X}\to i_{1*}i^{*}_{1}{\mathscr{O}}_{X}\to 0$
The nontrivial term on the left satisfies
  1. (1)
    ${\mathscr{M}}_{1}$ is the invertible sheaf ${\mathscr{O}}_{X}(-C_{2}-C_{3})$ associated to the divisor $-C_{2}-C_{3}.$
  2. (2)
    ${\rm Supp}\,{\mathscr{M}}_{1}\subset U_{1}.$
  3. (3)
    As $j^{!}_{1}{\mathscr{O}}_{X}={\mathscr{O}}_{U_{1}}$, from Corollary 2, it follows that $D^{b}(X)$ is generated by ${\mathscr{M}}_{1}$ together with $i_{1*}\{{\rm generators\;of}D^{b}(Z_{1})\}.$ We write
    (3.18) $D^{b}(X)=\langle{\mathscr{M}}_{1},\,i_{1*}D^{b}(Z_{1})\rangle.$
$U_{2}$ open subset of $Z_{1}.$ Next we will consider a divisor $Z_{2}=C_{4}+C_{5}$ of $Z_{1}$ and we apply to it the above analysis. We will first state the following simple lemma.
Lemma 2.
Let $Y$ be a complex variety, and $Y_{1},\,Y_{2}$ affine subvarieties of $Y$, such that $Y=Y_{1}\cup Y_{2}.$ If $Y_{k}$, for $k=1,2,$ is isomorphic to an affine variety, then the derived category $D^{b}(Y)$ is generated by
$\{\alpha_{1!}{\mathscr{O}}_{Y_{1}},\,\alpha_{2!}{\mathscr{O}}_{Y_{2}}\}\cup\alpha_{!}\{{\rm generators\,of}\,D^{b}(Y_{1}\cap Y_{2})\},$
where $\alpha$ is the inclusion $\alpha:Y_{1}\cap Y_{2}\hookrightarrow Y,$ and $\alpha_{k}:Y_{k}\hookrightarrow Y$ denotes the corresponding embedding.
Proof. Given ${\mathscr{F}}\in\mathbf{Coh}(Y),$ one has the coherent sheaves ${\mathscr{F}}_{k}:=\alpha_{k!}\alpha_{k}^{!}{\mathscr{F}}$ and ${\mathscr{F}}_{Y_{1}\cap Y_{2}}=\alpha_{!}\alpha^{!}{\mathscr{F}}$ and the exact sequence of ${\mathscr{O}}_{Y}$-modules [22, Prop. 2.3.6 (vii)]
$0\to{\mathscr{F}}_{Y_{1}\cap Y_{2}}\to{\mathscr{F}}_{1}\oplus{\mathscr{F}}_{2}\to{\mathscr{F}}\to 0$
By Proposition 7, $D(Y_{k})$ is generated by ${\mathscr{O}}_{Y_{k}}.$ From the above exact sequence, it follows that the sheaf ${\mathscr{F}}$ belongs to the subcategory of $D^{b}(Y)$ generated by $\alpha_{1!}{\mathscr{O}}_{Y_{1}},\,$ $\alpha_{2!}{\mathscr{O}}_{Y_{2}}$ together with the elements of form $\alpha_{!}(g),$ where $g$ belongs to a set of generators of $D^{b}(Y_{1}\cap Y_{2}).$
Going back to the variety $X$. Next we will consider a divisor $Z_{2}=C_{4}+C_{5}$ of $Z_{1}.$ Then $U_{2}:=Z_{1}\setminus Z_{2}=C_{2}^{0}\cup C_{3}^{0}.$ For $k=2,3$ we have the inclusions
$C^{0}_{k}\overset{\alpha_{k}}{\hookrightarrow}U_{2}\overset{j_{2}}{\hookrightarrow}Z_{1}\overset{i_{2}}{\hookleftarrow}Z_{2}.$
As $C^{0}_{2}\cap C^{0}_{3}=\emptyset,$ from Lemma 2 together with Corollary 2, it follows that the set $\{\alpha_{k!}{\mathscr{O}}_{C^{0}_{k}}\}_{k=2,3}$ is a set of generators of $D^{b}(U_{2}).$
For $k=2,3$ the ${\mathscr{O}}_{Z_{1}}$-module $(j_{2}\alpha_{k})_{!}{\mathscr{O}}_{C^{0}_{k}},$ will be denoted ${\mathscr{M}}_{k}.$
  1. (1)
[…]
  2. (2)
    ${\rm Supp}\,{\mathscr{M}}_{k}\subset C^{0}_{k}.$
  3. (3)
    From Corollary 2, we conclude
    $D^{b}(Z_{1})=\langle{\mathscr{M}}_{2},\,{\mathscr{M}}_{3},\,i_{2*}D^{b}(Z_{2})\rangle.$
From (3.18), it follows
(3.19) $D^{b}(X)=\langle{\mathscr{M}}_{1},\,i_{1*}{\mathscr{M}}_{2},\,i_{1*}{\mathscr{M}}_{3},\,i_{1*}i_{2*}D^{b}(Z_{2})\rangle.$
$U_{3}$ open subset of $Z_{2}.$ In a similar way, we define $U_{2}=C^{0}_{4}\cup C^{0}_{5}$ and $Z_{3}=Z_{2}\setminus U_{2}=C_{6}.$ For $k=4,5$ one has the embeddings
$C^{0}_{k}\overset{\alpha_{k}}{\hookrightarrow}U_{3}\overset{j_{3}}{\hookrightarrow}Z_{2}\overset{i_{3}}{\hookleftarrow}Z_{3}.$
Since $C^{0}_{4}\cap C^{0}_{5}=\emptyset,$ from Lemma 2 together with Corollary 2, it follows that the set $\{\alpha_{k!}{\mathscr{O}}_{C^{0}_{k}}\}_{k=4,5}$ is a set of generators of $D^{b}(U_{3}).$
For $k=4,5$ the ${\mathscr{O}}_{Z_{2}}$-module $(j_{3}\alpha_{k})_{!}{\mathscr{O}}_{C^{0}_{k}},$ will be denoted ${\mathscr{M}}_{k}.$
  1. (1)
[…]
  2. (2)
    ${\rm Supp}\,{\mathscr{M}}_{k}\subset C^{0}_{k}.$
  3. (3)
    By Corollary 2, $D^{b}(Z_{2})=\langle{\mathscr{M}}_{3},\,{\mathscr{M}}_{4},\,i_{3*}D^{b}(Z_{3})\rangle.$
From (3.19), it follows
(3.20) $D^{b}(X)=\langle{\mathscr{M}}_{1},\,i_{1*}{\mathscr{M}}_{2},\,i_{1*}{\mathscr{M}}_{3},\,(i_{1}i_{2})_{*}{\mathscr{M}}_{4},\,(i_{1}i_{2})_{*}{\mathscr{M}}_{5},\,(i_{1}i_{2}i_{3})_{*}D^{b}(Z_{3})\rangle.$
For the sake of simplicity in the notation we set
${\mathscr{L}}_{1}:={\mathscr{M}}_{1},\,{\mathscr{L}}_{2}:=i_{1*}{\mathscr{M}}_{2},\,{\mathscr{L}}_{3}:=i_{1*}{\mathscr{M}}_{3},\,{\mathscr{L}}_{4}:=(i_{1}i_{2})_{*}{\mathscr{M}}_{4},\,{\mathscr{L}}_{5}:=(i_{1}i_{2})_{*}{\mathscr{M}}_{5}.$
The point $Z_{3}.$ ${\mathscr{L}}_{6}$ will denote the skyscraper sheaf on $X$ at the point $C_{6}.$ Thus,
Proposition 9.
The derived category $D^{b}(X)$ is generated by the set $\{{\mathscr{L}}_{1},\dots,{\mathscr{L}}_{6}\}.$
Since ${\rm Supp}({\mathscr{L}}_{k})\subset C^{0}_{k}$ and $C^{0}_{r}\cap C^{0}_{s}=\emptyset.$ for $r\neq s.$ Then
(3.21) ${\rm Hom}_{{\mathscr{O}}_{X}}\big{(}{\mathscr{L}}_{r},\,{\mathscr{L}}_{s}\otimes_{{\mathscr{O}}_{X}}\Omega^{1}_{X}\big{)}=0.$
Proposition 10.
For $r=1,\dots,6$
${\rm Hom}_{{\mathscr{O}}_{X}}({\mathscr{L}}_{r},\,{\mathscr{L}}_{r}\otimes_{{\mathscr{O}}_{X}}\Omega^{1}_{X})=0.$
Proof. As ${\mathscr{L}}_{1}={\mathscr{M}}_{1}$ is the invertible sheaf ${\mathscr{O}}_{X}(-Z_{1}),$
${\rm Hom}_{{\mathscr{O}}_{X}}({\mathscr{L}}_{1},\,{\mathscr{L}}_{1}\otimes_{{\mathscr{O}}_{X}}\Omega^{1}_{X})={\rm Hom}_{{\mathscr{O}}_{X}}({\mathscr{O}}_{X},\,\Omega^{1}_{X})=H^{1,0}(X).$
By (3.17), $H^{1,0}(X)=0.$
The remaining cases can be treated similarly. To simplify the notation we will write ${\mathscr{X}}$ for ${\mathscr{O}}_{X}$ and ${\mathscr{Z}}_{1}$ for ${\mathscr{O}}_{Z_{1}};$ we also write $i:=i_{1}.$
As $\Omega^{1}_{X}$ is a locally free ${\mathscr{X}}$-module, by [22, Prop. 2.5.13]
${\mathscr{L}}_{2}\otimes_{\mathscr{X}}\Omega^{1}_{X}\simeq i_{!}\big{(}{\mathscr{M}}_{2}\otimes_{{\mathscr{Z}}_{1}}\Omega^{1}_{Z_{1}}\big{)}.$
Hence,
${\mathscr{H}om}_{\mathscr{X}}\big{(}{\mathscr{L}}_{2},\,{\mathscr{L}}_{2}\otimes_{\mathscr{X}}\Omega^{1}_{X}\big{)}\simeq{\mathscr{H}om}_{\mathscr{X}}\big{(}i_{*}{\mathscr{M}}_{2},\,i_{!}({\mathscr{M}}_{2}\otimes_{{\mathscr{Z}}_{1}}\Omega^{1}_{Z_{1}})\big{)}$
$\simeq i_{*}{\mathscr{H}om}_{{\mathscr{Z}}_{1}}\big{(}{\mathscr{M}}_{2},\,{\mathscr{M}}_{2}\otimes_{{\mathscr{Z}}_{1}}\Omega^{1}_{Z_{1}}\big{)}.$
Since ${\mathscr{M}}_{2}$ is an invertible ${\mathscr{Z}}_{1}$-module
${\rm Hom}_{\mathscr{X}}\big{(}{\mathscr{L}}_{2},\,{\mathscr{L}}_{2}\otimes_{\mathscr{X}}\Omega^{1}_{X}\big{)}\simeq\Gamma\big{(}Z_{1},\,{\mathscr{H}om}_{{\mathscr{Z}}_{1}}\big{(}{\mathscr{M}}_{2},\,{\mathscr{M}}_{2}\otimes_{{\mathscr{Z}}_{1}}\Omega^{1}_{Z_{1}}\big{)}\big{)}$
$={\rm Hom}_{{\mathscr{Z}}_{1}}\big{(}{\mathscr{M}}_{2},\,{\mathscr{M}}_{2}\otimes_{{\mathscr{Z}}_{1}}\Omega^{1}_{Z_{1}}\big{)}=H^{0}(Z_{1},\,\Omega^{1}_{Z_{1}})$
On the other hand, $Z_{1}=C_{2}\cup C_{3}$, and $H^{1,0}(C_{2})=H^{1,0}(C_{3})=0.$ Thus, $H^{0}(Z_{1},\,\Omega^{1}_{Z_{1}})=0.$ That is, the proposition is proved for the case $r=2.$ The proof of the remaining cases is analogous. ∎
Theorem 4.
The set of holomorphic gauge fields on any $B$-brane over the variety of complete flags in ${\mathbb{C}}^{3}$ is $\leq 1$.
Proof. An object of $D^{b}(X)$ is isomorphic to one of the triangulated category generated by the family ${\mathscr{L}}_{1},\dots,{\mathscr{L}}_{6}$ in $D^{b}(X).$ Let $({G}^{\bullet},\,d^{\bullet})$ be an object of this triangulated category. Then ${G}^{p}$ is a sheaf of the form
(3.22) ${G}^{p}=\bigoplus_{i\in S_{p}}{\mathscr{L}}_{pi},$
$S_{p}$ being a finite set. From Corollary 1 together with (3.21) and Proposition 10, it follows
${\rm Hom}_{D^{b}(X)}\big{(}{G}^{\bullet},\,\Omega^{1}(G^{\bullet})\big{)}=0.$
The theorem is a consequence of Proposition 5. ∎
3.4. The homotopy category
Let $X$ be a smooth projective variety. The category $\mathbf{Coh}(X)$ of coherent sheaves over $X$ has not enough injectives, for this reason it is convenient to regard $D^{b}(X)$ as a subcategory of $D^{b}(\mathbf{Mod}({\mathscr{O}_{X}}))$, the bounded derived category of the ${\mathscr{O}}_{X}$-modules. In fact, $D^{b}(X)$ is equivalent to $D^{b}(\mathbf{Mod}({\mathscr{O}_{X}}))_{\rm coh}$, the full subcategory of $D^{b}(\mathbf{Mod}({\mathscr{O}_{X}}))$ consisting of the complexes with coherent cohomology [29, Exp II] [15, Chapter II] (for a brief description [19, page 63]). Thus, $D^{b}(X)$ can be identified with the homotopy category whose objects are the complexes ${\mathscr{G}}^{\bullet}$ of injective ${\mathscr{O}}_{X}$-modules, such that its cohomogy is bounded and coherent; i.e. ${\mathscr{H}}^{j}({\mathscr{G}}^{\bullet})$ is coherent and vanishes for $|j|>>0$. We will denote this homotopy category by $K^{b}(X)_{\rm coh}$.
In the identification of $D^{b}(X)$ with the homotopy category $K^{b}(X)_{\rm coh},$ the quasi-isomorphisms correspond to homotopy equivalences. Thus, the independence of quasi-isomorphisms in the frame of $K(X)_{\rm coh}$ means independence with respect to homotopy equivalences. In particular, if ${\mathscr{F}}^{\bullet}$ and ${\mathscr{G}}^{\bullet}$ are objects of $K^{b}(X)_{\rm coh}$ homotopically equivalent, the respective cohomology sheaves ${\mathscr{H}}^{i}({\mathscr{F}}^{\bullet})$ and ${\mathscr{H}}^{i}({\mathscr{G}}^{\bullet})$ are isomorphic.
If $Y$ is a Hodge manifold, then it is a smooth projective variety, according to a well-known Kodaira’s theorem. By the GAGA correspondence, the coherent analytic sheaves on $Y$ can be identified with the algebraic ones. Hence, the $B$-branes on $Y$ can be considered as objects of $K^{b}(Y)_{\rm coh}$.
From now on, we assume that $Y$ is a Hodge manifold. In accordance with the preceding paragraph, one can assume that the $B$-brane $({\mathscr{F}}^{\bullet},\delta^{\bullet})$ over $Y$ is a complex of injective ${\mathscr{O}}$-modules with coherent cohomology modules satisfying ${\mathscr{H}}^{i}({\mathscr{F}}^{\bullet})=0$ for $|i|>>0$. Moreover,
${\rm Hom}_{D^{b}(Y)}({\mathscr{F}}^{\bullet},\,{\mathscr{J}}^{1}({\mathscr{F}}^{\bullet}))\simeq{\rm Hom}_{K^{b}(Y)_{\rm coh}}({\mathscr{F}}^{\bullet},\,\widehat{{\mathscr{J}}^{1}({\mathscr{F}}^{\bullet}})),$
where $\widehat{{\mathscr{J}}^{1}({\mathscr{F}}^{\bullet}})$ is an object of $K^{b}(Y)_{\rm coh}$ $q$-isomorphic to ${\mathscr{J}}^{1}({\mathscr{F}}^{\bullet})$. Thus, the elements of the space (3.1) can be identified with morphisms in the homotopyy category $K^{b}(Y)_{\rm coh}$. Given a gauge $\psi\in{\rm Hom}_{D^{b}(Y)}({\mathscr{F}}^{\bullet},\,{\mathscr{J}}^{1}({\mathscr{F}}^{\bullet})),$ we denote by $\hat{\psi}$ the corresponding morphism in the category $K^{b}(Y)_{\rm coh}.$ The exact sequence (3.4) gives rise to the following sequence of arrows in the additive category $K^{b}(Y)_{\rm coh}$
$\widehat{\Omega^{1}({\mathscr{F}}^{\bullet})}\overset{\hat{\iota}}{\rightarrow}\widehat{{\mathscr{J}}^{1}({\mathscr{F}}^{\bullet})}\overset{\hat{\pi}}{\rightarrow}{\mathscr{F}}^{\bullet}$
where $\widehat{\Omega^{1}({\mathscr{F}}^{\bullet})}$ is quasi-isomorphic to $\Omega^{1}({\mathscr{F}}^{\bullet})$ and $\hat{\pi}\circ\hat{\psi}=1.$ We will take to the cohomology of those complexes taking into account that:
  • The cohomology of an object with a hat is isomorphic to the cohomology of the corresponding complex without hat.
  • A morphism of the homotopy category, determines a well-defined morphism between the cohomologies; i. e. a morphism of complexes of ${\mathscr{O}}$-modules.
  • ${\mathscr{H}}^{*}(\Omega^{1}({\mathscr{F}}^{\bullet}))\simeq\Omega^{1}({\mathscr{H}}^{*}({\mathscr{F}}^{\bullet})).$
Then, one has
(3.23) $0\to\Omega^{1}({\mathscr{H}}^{j}({\mathscr{F}}^{\bullet}))\overset{\boldsymbol{\iota}^{j}}{\longrightarrow}{\mathscr{H}}^{j}({\mathscr{J}}^{1}({\mathscr{F}}^{\bullet}))\overset{\boldsymbol{\pi}^{j}}{\longrightarrow}{\mathscr{H}}^{j}({\mathscr{F}}^{\bullet})\to 0,$
where the maps induced between the cohomologies are denoted in bold. In turn, $\hat{\psi}$ determines a well-defined morphism of ${\mathscr{O}}$-modules between the cohomologies
$\boldsymbol{\psi}^{j}:{\mathscr{H}}^{j}({\mathscr{F}}^{\bullet})\to{\mathscr{H}}^{j}({\mathscr{J}}^{1}({\mathscr{F}}^{\bullet})),$
that obviously satisfies $\boldsymbol{\pi}^{j}\circ\boldsymbol{\psi}^{j}=1.$
Hence, the sequence (3.23) of ${\mathscr{O}}$-modules splits. Furthermore, by (3.5), ${\mathscr{J}}^{1}({\mathscr{H}}^{j}({\mathscr{F}}^{\bullet}))\simeq{\mathscr{H}}^{j}({\mathscr{J}}^{1}({\mathscr{F}}^{\bullet})).$ Thus, we have a splitting, in the category of ${\mathscr{O}}$-modules, of ${\mathscr{J}}^{1}({\mathscr{H}}^{j}({\mathscr{F}}^{\bullet}))$ as direct sum ${\mathscr{H}}^{j}({\mathscr{F}}^{\bullet})$ and $\Omega^{1}({\mathscr{H}}^{j}({\mathscr{F}}^{\bullet})).$
According to Definition 1, $\boldsymbol{\psi}^{j}$ defines a holomorphic connection on the ${\mathscr{O}}$-module ${\mathscr{H}}^{j}({\mathscr{F}}^{\bullet}).$
We set $\boldsymbol{\eta}^{j}$ for the morphism of abelian sheaves defined by the inclusion in the direct sum
$\boldsymbol{\eta}^{j}:\sf{Forg}\big{(}{\mathscr{H}}^{j}({\mathscr{F}}^{\bullet})\big{)}\rightarrow\sf{Fog}\big{(}{\mathscr{H}}^{j}({\mathscr{J}}^{1}({\mathscr{F}}^{\bullet}))\big{)}=\sf{Forg}\big{(}{\mathscr{H}}^{j}({\mathscr{F}}^{\bullet})\oplus\Omega^{1}({\mathscr{H}}^{j}({\mathscr{F}}^{\bullet}))\big{)},$
Hence, $\boldsymbol{\pi}^{j}(\boldsymbol{\psi}^{j}-\boldsymbol{\eta}^{j})=0$, and thus $\boldsymbol{\psi}^{j}-\boldsymbol{\eta}^{j}$ defines a morphism of abelian sheaves
(3.24) $\vartheta^{j}:{\mathscr{H}}^{j}({\mathscr{F}})\to{\Omega^{1}}({\mathscr{H}}^{j}({\mathscr{F}}^{\bullet})),$
which, by (3.2), satisfies the Leibniz’s rule (2.14). That is,
Proposition 11.
The gauge field $\psi$ on the brane ${\mathscr{F}}^{\bullet}$ determines on each ${\mathscr{O}}$-module ${\mathscr{H}}^{j}({\mathscr{F}}^{\bullet})$ a holomorphic connection $\vartheta^{j}$.
Definition 4.
The gauge field $\psi$ is called flat, if the curvature of the connection $\vartheta^{j}$ vanishes, for all $j$.
Remark 7.
Let $\psi,\phi$ be two gauge fields on ${\mathscr{F}}$, we set
$\xi:=\phi-\psi\in{\rm Hom}_{D^{b}(Y)}\big{(}{\mathscr{F}}^{\bullet},\,{\mathscr{J}}^{1}({\mathscr{F}}^{\bullet})\big{)},$
and denote by $\hat{\xi}$ the corresponding morphism ${\mathscr{F}}^{\bullet}\to\widehat{{\mathscr{J}}^{1}({\mathscr{F}}^{\bullet})}$ in the category $K^{b}(Y)_{\rm coh}.$ Thus, $\hat{\xi}$ determines a well defined morphism of ${\mathscr{O}}$-modules between the cohomologies, $\boldsymbol{\xi}^{j}:{\mathscr{H}}^{j}({\mathscr{F}}^{\bullet})\to{\mathscr{H}}^{j}({\mathscr{J}}^{1}({\mathscr{F}}^{\bullet})).$
On the other hand, as $\phi$ and $\psi$ are gauge fields, $\hat{\pi}\hat{\xi}=0$. Hence, $\boldsymbol{\xi}^{j}$ defines morphisms of ${\mathscr{O}}$-modules $\zeta^{j}({\xi}):{\mathscr{H}}^{j}({\mathscr{F}}^{\bullet})\to\Omega^{1}({\mathscr{H}}^{j}({\mathscr{F}}^{\bullet})).$ We denote by $\vartheta^{j}$ and $\chi^{j}$ the connections on ${\mathscr{H}}^{j}({\mathscr{F}})$ determined by $\hat{\psi}$ and $\hat{\phi}$, respectively. Since $\boldsymbol{\xi}^{j}=(\boldsymbol{\phi}^{j}-\boldsymbol{\eta}^{j})-(\boldsymbol{\psi}^{j}-\boldsymbol{\eta}^{j})$, it follows that $\chi^{j}=\vartheta^{j}+{\zeta}^{j}(\xi)$. In short, $\zeta^{j}(\xi)$ is the “variation” on the connection $\vartheta^{j}$ induced by the “variation” $\xi$ of the gauge field $\psi$.
3.5. Yang-Mills fields on a brane.
The result deduced in the following paragraph gives us a suggestion for the definition of the Yang-Mills functional over the gauge fields on a brane over a Hodge manifold.
3.5.1. An Euler-Poincaré mapping.
Let ${\mathscr{A}}$ be a coherent sheaf on the Hodge manifold $Y,$ and $\alpha:{\mathscr{A}}\to\Omega^{1}({\mathscr{A}})$ a holomorphic connection on ${\mathscr{A}}$. Denoting by ${\mathcal{S}}_{\mathscr{A}}$ the singular set of ${\mathscr{A}}$, on $Y\setminus{\mathcal{S}}_{\mathscr{A}}$ we define differential form
$\Phi({\mathscr{A}},\alpha):={\rm tr}\big{(}K_{\alpha}\wedge\star K_{\alpha}\big{)}\in\Gamma(Y\setminus{\mathcal{S}}_{\mathscr{A}},\,\Omega^{\rm top}\big{)},$
$K_{\alpha}$ being the curvature of $\alpha$ considered as an ${\rm End}(A)$-valued $2$-form.
By $\mathfrak{C}$, we denote the category whose objects are pairs $({\mathscr{A}},\,\alpha)$. A morphism $f:({\mathscr{A}},\,\alpha)\to({\mathscr{B}},\,\beta)$ is a morphism of coherent sheaves compatible with the connections; i.e. such that $(1\otimes f)\circ\alpha=\beta\circ f$.
Proposition 12.
[…]
$\Phi({\mathscr{B}},\beta)=\Phi({\mathscr{A}},\alpha)+\Phi({\mathscr{C}},\gamma),$
where ${\mathcal{S}}$ is the union of the singular sets of ${\mathscr{A}},$ ${\mathscr{B}},$ and ${\mathscr{C}}$.
Proof. Let $y_{0}\in Y\setminus{\mathcal{S}}$. As the exact sequence splits locally on $Y\setminus{\mathcal{S}}$, there exists an open neighborhood $U$ of $y_{0}$ such that $g|_{U}$, in the sequence of locally free modules $0\to{\mathscr{A}}|_{U}\overset{f|_{U}}{\to}{\mathscr{B}}|_{U}\overset{g|_{U}}{\to}{\mathscr{C}}|_{U}\to 0,$ has a right inverse $h$.
Let $a$ be a frame for ${\mathscr{A}}|_{U}$, then $\alpha(a)={\sf A}\cdot a$, where ${\sf A}$ is a matrix of $1$-forms on $U$. Furthermore, $a$ can be chosen so that ${\sf A}(y_{0})=0$. Similarly, let $c$ be a frame for ${\mathscr{C}}|_{U}$, then $\gamma(c)={\sf C}\cdot c$ and we choose $c$ so that ${\sf C}(y_{0})=0$. From the splitting, it follows that $\{f(a),\,h(c)\}$ is a frame for ${\mathscr{B}}|_{U}$. By the compatibility of the connections with $f$ and $g$,
$\beta(f(a))=(1\otimes f)(\alpha(a))=(1\otimes f)({\sf A}\cdot a)={\sf A}\cdot f(a).$
On the other hand, $\beta(h(c))={\sf R}\cdot f(a)+{\sf S}\cdot h(c)$, with ${\sf R}$ and ${\sf S}$ matrices of $1$-forms. But,
${\sf C}\cdot c=\gamma(c)=\gamma(gh(c))=(1\otimes g)(\beta(h(c))=(1\otimes g)\big{(}{\sf R}\cdot f(a)+{\sf S}\cdot h(c)\big{)}.$
As $g\circ f=0$ and $g\circ h=1$, it follows that ${\sf C}={\sf S}.$ That is, the matrix of the connection $\beta$ in the frame $\{f(a),\,h(c)\}$ is
(3.25) ${\sf M}:=\begin{pmatrix}{\sf A}&{\sf R}\\ 0&{\sf C}\end{pmatrix}$
Since ${\sf A}(y_{0})=0$ and ${\sf C}(y_{0})=0$, the matrix of $K_{\alpha}(y_{0})$, of the curvature of $\alpha$ at the point $y_{0}$, is $d{\sf A}$. Analogously, the matrix of $K_{\gamma}(y_{0})$ is $d{\sf C}$. The one of $K_{\beta}(y_{0})$ is the exterior derivative of (3.25), since ${\sf M}\wedge{\sf M}=0$ at $y_{0}$. Then
${\rm tr}\big{(}K_{\beta}(y_{0})\wedge\star K_{\beta}(y_{0})\big{)}$ $={\rm tr}(d{\sf A}\wedge\star d{\sf A})+{\rm tr}(d{\sf C}\wedge\star d{\sf C})$
$={\rm tr}\big{(}K_{\alpha}(y_{0})\wedge\star K_{\alpha}(y_{0})\big{)}+{\rm tr}\big{(}K_{\gamma}(y_{0})\wedge\star K_{\gamma}(y_{0})\big{)}.$
As $y_{0}$ is an arbitrary point of $Y\setminus{\mathcal{S}}$, it follows the proposition. ∎
Let $({\mathscr{G}}^{\bullet},\,{\delta}^{\bullet})$ be a bounded complex of coherent sheaves on the manifold $Y$. Let $\nabla^{\bullet}$ be a family of holomorphic connections, compatible with the operators $\delta^{\bullet}$. That is, $\nabla^{i}:{\mathscr{G}}^{i}\rightarrow\Omega^{1}({\mathscr{G}}^{i})$ is a holomorphic gauge field on the coherent sheaf ${\mathscr{G}}^{i}$ such that
$(1\otimes\delta^{i})\nabla^{i}=\nabla^{i+1}\delta^{i}.$
Hence, $\nabla^{i}({\mathscr{K}er}({\delta}^{i}))\subset{\mathscr{K}er}(1\otimes{\delta}^{i})$ and a similar relation for the image ${\mathscr{I}m}(\delta^{i-1})$. It follows that $\nabla^{i}$ induces a connection $\theta^{i}$ on the cohomology
$\theta^{i}:{\mathscr{H}}^{i}({\mathscr{G}}^{\bullet})\to\Omega^{1}({\mathscr{H}}^{i}({\mathscr{G}}^{\bullet})).$
Obviously, the restrictions of $\nabla^{i}$ determine connections on ${\mathscr{K}er}(\delta^{i})$ and ${\mathscr{I}m}(\delta^{i+1})$, respectively. One has the exact sequence
(3.26) $0\to\big{(}{\mathscr{K}er}(\delta^{i}),\,\nabla^{i}\big{)}\to\big{(}{\mathscr{G}}^{i},\,\nabla^{i}\big{)}\to\big{(}{\mathscr{I}m}(\delta^{i}),\,\nabla^{i+1}\big{)}\to 0$
in the category $\mathfrak{C}$. Similarly, we have the exact sequence
(3.27) $0\to\big{(}{\mathscr{I}m}(\delta^{i-1}),\,\nabla^{i}\big{)}\to\big{(}{\mathscr{K}er}(\delta^{i}),\,\nabla^{i}\big{)}\to\big{(}{\mathscr{H}}^{i},\,\theta^{i}\big{)}\to 0$
Corollary 3.
Denoting with ${\mathcal{S}}$ the union of the singular sets of the sheaves ${\mathscr{G}}^{i}$, then on $Y\setminus{\mathcal{S}}$
(3.28) $\sum_{i}(-1)^{i}{\rm tr}\big{(}K_{\nabla^{i}}\wedge\star K_{\nabla^{i}}\big{)}=\sum_{i}(-1)^{i}{\rm tr}\big{(}K_{\theta^{i}}\wedge\star K_{\theta^{i}}\big{)}.$
Proof. From Proposition 12 together with (3.26), it follows
$\Phi({\mathscr{G}}^{i},\,\nabla^{i})=\Phi({\mathscr{K}er}(\delta^{i}),\,\nabla^{i})+\Phi({\mathscr{I}m}(\delta^{i}),\,\nabla^{i+1}).$
From (3.27), one obtains an analogous relation. Taking the alternate sums
$\sum_{i}(-1)^{i}\Phi({\mathscr{G}}^{i},\,\nabla^{i})=\sum_{i}(-1)^{i}\Phi({\mathscr{H}}^{i},\,\theta^{i}).$
3.5.2. The Yang-Mills functional.
We propose a definition for the Yang-Mills functional over gauge fields on a brane. This proposal is based on the following considerations:
  1. (1)
    It is reasonable to require that this definition generalizes the one for coherent sheaves.
  2. (2)
    As a gauge field is a homotopy class of a morphism of complexes, it seems convenient to move on the cohomology of these complexes.
  3. (3)
    Let $E^{\bullet}$ be a bounded complex of Hermitian vector bundles over the Kähler manifold $Y,$ and $\nabla^{\bullet}$ a family of connections compatible with the Hermitian metrics and the coboundary operators. Denoting by $H^{i}(E^{\bullet})$ the cohomology bundles, there exist connections $\theta^{i}$ on those bundles, induced by the family $\nabla^{\bullet}$. By Corollary 3 together with (2.19), one has the following equality of Euler-Poincaré type.
    $\sum_{i}(-1)^{i}\lVert K_{\nabla^{i}}\rVert^{2}=\sum_{i}(-1)^{i}\lVert(K_{\theta^{i}}\rVert^{2}.$
On the basis of the above considerations, it seems appropriate to define the value of the Yang-Mills functional on the gauge $\psi$ on the brane ${\mathscr{F}}^{\bullet}$ as $\sum_{i}(-1)^{i}\lVert K_{\vartheta_{i}}\rVert^{2}.$
More precisely, taking into account Proposition 11, we adopt the following definitions.
Definition 5.
The brane $({\mathscr{F}}^{\bullet},\,\delta^{\bullet})$ is called a Hermitian brane, if the cohomology sheaves ${\mathscr{H}}^{j}$ are Hermitian.
Let $({\mathscr{F}}^{\bullet},\,{\delta}^{\bullet})$ be a Hermitian brane on the Hodge manifold $Y$. Given a gauge field $\psi$ on the brane $({\mathscr{F}}^{\bullet},\,\delta^{\bullet})$, by Proposition 11, one has the family of curvatures ${\mathcal{K}}_{\vartheta^{i}}$ of the connections induced on the cohomologies. Whose norms $\lVert{\mathcal{K}}_{\vartheta^{i}}\rVert$ are defined in accordance with (2.18).
Definition 6.
Given a gauge field $\psi$ on the Hermitian $B$-brane $({\mathscr{F}}^{\bullet},\,\delta^{\bullet})$, if the sheaves ${\mathscr{H}}^{i}({\mathscr{F}}^{\bullet})$, are reflexive, we define the value of the Yang-Mills functional at $\psi$ by
(3.29) $\mathcal{YM}(\psi)=\sum_{i}(-1)^{i}\lVert{\mathcal{K}}_{\vartheta_{i}}\rVert^{2}.$
Thus, $\mathcal{YM}(\psi)$ is a kind of Euler characteristic of the gauge field. The Yang-Mills fields on the brane $({\mathscr{F}}^{\bullet},\,{\delta}^{\bullet})$ are the critical points of the functional $\psi\mapsto\mathcal{YM}(\psi).$
Note that, if $({\mathscr{F}}^{\bullet},\,\delta^{\bullet})$ is an acyclic complex, then the Yang-Mills functional for this complex is identically zero.
Let ${\mathscr{A}}^{\bullet}:=\big{(}{\mathscr{A}}^{\bullet},\,d_{A}^{\bullet},\,\alpha^{\bullet}\big{)}$ be a complex in the category ${\mathfrak{C}}$; i.e, a complex of coherent sheaves with a family of holomorphic connections compatible with the coboundary operator $d_{A}$. Let $f:=(f^{\bullet})$ a morphism $f^{\bullet}:{\mathscr{A}}^{\bullet}\to{\mathscr{B}}^{\bullet}$ in ${\mathfrak{C}}$; that is, $f$ is a morphism of complexes compatible with the connections. Let ${\mathscr{C}}^{\bullet}$ denote the mapping cone of $f$. Thus, ${\mathscr{C}}^{\bullet}=\big{(}{\mathscr{A}}^{\bullet}[1]\oplus{\mathscr{B}}^{\bullet},d_{C}^{\bullet},\nabla^{\bullet},\big{)}$, with $d_{C}(a,\,b)=\big{(}d(a),\,(-1)^{{\rm degree}\,a}f(a)+db\big{)}$ and $\nabla(a,\,b)=(\alpha(a),\,\beta(b)).$ In fact, $(1\otimes d_{C})\circ\nabla=\nabla\circ d_{C}$ and thus ${\mathscr{C}}^{\bullet}$ is a complex of the category ${\mathfrak{C}}$
For each $i$ one has the following exact sequence in the category ${\mathfrak{C}}$
$0\to{\mathscr{B}}^{i}\to{\mathscr{C}}^{i}\to{\mathscr{A}}^{i}[1]\to 0$
From Proposition 12, $\Phi({\mathscr{B}}^{i})+\Phi({\mathscr{A}}^{i+1})=\Phi({\mathscr{C}}^{i})$. Multiplying by $(-1)^{i}$ and summing
(3.30) $\sum_{i}(-1)^{i}{\rm tr}(K_{\beta^{i}}\wedge\star K_{\beta^{i}})+$ $\sum_{i}(-1)^{i}{\rm tr}(K_{\alpha^{i+1}}\wedge\star K_{\alpha^{i+1}})$
$=\sum_{i}(-1)^{i}{\rm tr}(K_{\nabla^{i}}\wedge\star K_{\nabla^{i}}).$
Let us assume that
  • ${\mathscr{A}}^{i}$ and ${\mathscr{B}}^{i}$ Hermitian sheaves for all $i$.
  • $\alpha^{i}$ and $\beta^{i}$ are Hermitian gauge fields (i.e., compatible with the metric) on ${\mathscr{A}}^{i}$ and ${\mathscr{B}}^{i}$, respectively.
Then one defines on ${\mathscr{C}}^{i}$ the metric $\langle(a,b),\,(a^{\prime},b^{\prime})\rangle:=\langle a,\,a^{\prime}\rangle+\langle b,\,b^{\prime}\rangle$. The connection $\nabla^{i}$ is compatible with this metric. From the equality (3.30) together with (2.19), one deduces the following proposition.
Proposition 13.
With the above notations and under the above hypotheses, $\alpha$ and $\beta$ determine in a natural way a gauge field $\nabla$ on the mapping cone of $f^{\bullet}$ satisfying
(3.31) $\mathcal{YM}(\beta)-\mathcal{YM}(\alpha)=\mathcal{YM}(\nabla).$
On the other hand, in the context of the branes theory, the fact that the branes ${\mathscr{A}}^{\bullet},\,{\mathscr{B}}^{\bullet}$ and ${\mathscr{C}}^{\bullet}$ are the members of the distinguished triangle ${\mathscr{A}}^{\bullet}\to{\mathscr{B}}^{\bullet}\to{\mathscr{C}}^{\bullet}\overset{+1}{\to}$ means that ${\mathscr{A}}^{\bullet}$ and ${\mathscr{C}}^{\bullet}$ can potentially bind together to form the membrane ${\mathscr{B}}^{\bullet}$ [2, Section 6.2.1]. Thus, the additive nature of equation (3.31) is consistent with this interpretation.
From now on, we assume that ${\mathscr{F}}^{\bullet}$ is a Hermitian $B$-brane such that the cohomology sheaves ${\mathscr{H}}^{i}({\mathscr{F}}^{\bullet})$ are reflexive.
Let us suppose that the set of gauge fields on the brane ${\mathscr{F}}^{\bullet}$ is nonempty. Let $m$ be the dimension of the vector space ${\rm Ext}^{0}({\mathscr{F}}^{\bullet},\,\Omega^{1}({\mathscr{F}}^{\bullet})).$ We denote by $\xi_{1},\dots,\xi_{m}$ a basis of this vector space. According to Proposition 5, any gauge field $\psi$ on the brane can be expressed
$\psi=\tilde{\psi}+\sum_{a}\lambda_{a}\xi_{a}$
$\tilde{\psi}$ being a fixed gauge field and $\lambda_{a}\in{\mathbb{C}}.$ Hence, the connections on the cohomology sheaves ${\mathscr{H}}^{i}$ can be written in the form (see Remark 7)
$\vartheta^{i}=\tilde{\vartheta}^{i}+\sum_{a}\lambda_{a}\zeta^{i}_{a},$
with $\zeta^{i}_{a}\in{\rm Hom}({\mathscr{H}}^{i},\,\Omega^{1}({\mathscr{H}}^{i}))$. The corresponding curvatures satisfy
$K_{\vartheta^{i}}=K_{\tilde{\vartheta}^{i}}+\sum_{a}\lambda_{a}\tilde{\vartheta}^{i}(\zeta^{i}_{a})+\sum_{a,b}\lambda_{a}\lambda_{b}\zeta^{i}_{a}\wedge\zeta^{i}_{b}.$
Therefore, $\lVert K_{\vartheta^{i}}\rVert^{2}$ is a polynomial $P^{i}(\lambda_{1},\dots,\lambda_{m})$ of degree $\leq 4.$ Thus, the critical points of the Yang-Mills functional correspond to the points $(\lambda_{1},\dots,\lambda_{m})\in{\mathbb{C}}^{m}$ which satisfy the equations $\frac{\partial P}{\partial\lambda_{a}}=0,$ where $P$ is the polynomial $\sum_{i}(-1)^{i}P^{i}(\lambda_{1},\dots,\lambda_{m})$. We have the following result, which generalizes Theorem 1.
Theorem 5.
Assumed the set of gauge fields on the brane ${\mathscr{F}}^{\bullet}$ is nonempty and $m={\rm dim}\,{\rm Ext}^{0}({\mathscr{F}}^{\bullet},\,\Omega^{1}({\mathscr{F}}^{\bullet}))$. Then the set of Yang-Mills fields on ${\mathscr{F}}^{\bullet}$ are in bijective correspondence with the points of a subvariety of ${\mathbb{C}}^{m}$ defined by $m$ polynomials of degree $\leq 3$.
Let $({\mathscr{F}}^{\bullet},\,\delta^{\bullet})$ be a $B$-brane, such that the ${\mathscr{H}}^{i}({\mathscr{F}}^{\bullet})$ are reflexive sheaves. If $\phi$ and $\psi$ are gauge fields on ${\mathscr{F}}^{\bullet}$ and $\xi=\phi-\psi$, using the notations introduced in Remark 7, the connections on the cohomologies induced by $\phi$ and $\psi$ satisfy $\chi^{j}(\xi)=\vartheta^{j}+\zeta^{j}(\xi)$, with
$\zeta^{j}(\xi)\in\Gamma(Y\setminus{\mathcal{S}},\,\Omega^{1}\otimes_{\mathscr{O}}{\mathscr{E}nd}(H^{j})),$
${\mathcal{S}}$ being the union of the singularity sets of the reflexive sheaves ${\mathscr{H}}^{j}.$
With the mentioned notation, an infinitesimal variation $\psi_{\epsilon}$ of $\psi$ is an element of the form $\epsilon\xi,$ where $\epsilon\in{\mathbb{C}}$ and $\xi\in{\rm Ext}^{0}({\mathscr{F}}^{\bullet},\,\Omega^{1}({\mathscr{F}}^{\bullet}).$ The relation $\psi_{\epsilon}=\psi+\epsilon\xi$ gives rise to the following equality between the connections on the cohomologies
$\vartheta^{j}_{\epsilon}=\vartheta^{j}+\epsilon\zeta^{j}.$
Furthermore, on $Y\setminus{\mathcal{S}}$ the curvatures satisfy
$K_{\vartheta^{j}_{\epsilon}}=K_{\vartheta^{j}}+\epsilon\vartheta^{j}({\zeta^{j}})+O((\epsilon)^{2}),$
$\vartheta^{j}({\zeta^{j}})$ being the covariant derivative of ${\zeta^{j}}$ considered as a section of $\Omega^{1}\otimes_{\mathscr{O}}{\mathscr{E}nd}(H^{j})$.
From (2.22) together with Definition 6, the functional $\mathcal{YM}$ takes at the gauge field ${\psi}$ a stationary value if
(3.32) $(1/2)\frac{d}{d\epsilon}\mathcal{YM}(\psi_{\epsilon})\Big{|}_{\epsilon=0}=\sum_{j}(-1)^{j}\langle K_{\vartheta^{j}},\,\vartheta^{j}(\zeta^{j})\rangle=0,$
for all any variation of $\psi$.
In particular, if $\vartheta^{i}$ es a Yang-Mills field for all $i,$ then by (2.23)
(3.33) $\langle K_{\vartheta^{i}},\,\vartheta^{i}(\zeta^{i})\rangle=0,$
and we have the following proposition.
Proposition 14.
Let $({\mathscr{F}}^{\bullet},\,\delta^{\bullet})$ be a Hermitian brane, such that the sheaves ${\mathscr{H}}^{i}({\mathscr{F}}^{\bullet})$ are reflexive, and let $\psi$ gauge field on ${\mathscr{F}}^{\bullet}$. If $\vartheta^{i}$ is a Yang-Mills field on ${\mathscr{H}}^{i}$ for all $i$, then $\psi$ is a stationary point of the Yang-Mills functional; i.e. $\psi$ is a Yang-Mills field on the brane.
The following theorem is a converse to Proposition 14.
Theorem 6.
Let ${\mathscr{F}}^{\bullet}$ be a Hermitian $B$-brane as in Proposition 14. If $\nabla^{\bullet}$ is a Yang-Mills field on ${\mathscr{F}}^{\bullet},$ then the connection $\vartheta^{j}$ induced on ${\mathscr{H}}^{j}$ is a Yang-Mills field on this sheaf.
Proof. We will consider $({\mathscr{F}}^{\bullet},\,\delta^{\bullet})$ as an object of the category $K^{b}_{\rm coh}(Y)$; that is, we assume that $({\mathscr{F}}^{\bullet},\,\delta^{\bullet})$ is a complex of injective ${\mathscr{O}}$-modules such that its cohomology is bounded and coherent.
As ${\mathscr{K}er}(\delta^{i})$ is a submodule of the injective ${\mathscr{O}}$-module ${\mathscr{F}}^{i}$, then ${\mathscr{K}er}(\delta^{i})$ is a retract of ${\mathscr{F}}^{i}$ and thus it is also an injective ${\mathscr{O}}$-module. Therefore, the following short exact sequence of coherent ${\mathscr{O}}$-modules
$0\to{\mathscr{K}er}(\delta^{i})\to{\mathscr{F}}^{i}\to{\mathscr{C}oim}(\delta^{i})\to 0$
splits. That is, ${\mathscr{F}}^{i}\simeq{\mathscr{K}er}(\delta^{i})\oplus{\mathscr{C}oim}(\delta^{i}).$
Since ${\mathscr{I}m}({\delta}^{i-1})$ is a retract of the injective ${\mathscr{O}}$-module ${\mathscr{K}er}(\delta^{i})$, the following short exact sequence also splits
$0\to{\mathscr{I}m}({\delta}^{i-1})\to{\mathscr{K}er}(\delta^{i})\to{\mathscr{H}}^{i}\to 0$
Thus,
(3.34) ${\mathscr{F}}^{i}\simeq{\mathscr{H}}^{i}\oplus{\mathscr{G}}^{i},$
where ${\mathscr{G}}^{i}$ is isomorphic to the direct sum of ${\mathscr{C}oim}(\delta^{i})$ and ${\mathscr{I}m}(\delta^{i-1})$. As ${\mathscr{H}}^{i}$ and ${\mathscr{G}}^{i}$ are summands in a direct sum decomposition of an injective ${\mathscr{O}}$-module, they are also injective.
On the other hand, the coboundary operator $\delta^{i}:{\mathscr{F}}^{i}\to{\mathscr{F}}^{i+1}$ induces via the isomorphisms (3.34) to the morphism
(3.35) $\delta^{i}:{\mathscr{H}}^{i}\oplus{\mathscr{G}}^{i}\to{\mathscr{H}}^{i+1}\oplus{\mathscr{G}}^{i+1},\;\;\;(a,\,b)\mapsto(0,\,\delta^{i}b).$
Given $\xi\in{\rm Hom}_{K^{b}_{\rm coh}(Y)}\big{(}{\mathscr{F}}^{\bullet},\,\Omega^{1}\otimes_{\mathscr{O}}{\mathscr{F}}^{\bullet}\big{)}$, according to Remark 7, it determines $\zeta^{i}\in{\rm Hom}\big{(}{\mathscr{H}}^{i},\,\Omega^{1}\otimes_{\mathscr{O}}{\mathscr{H}}^{i}\big{)}.$ As $\nabla^{\bullet}$ is, by hypothesis, a Yang-Mills field then (3.32) is satisfied.
Given $j$, a general “variation” of $\vartheta^{j}$ is defined by an element $\tau\in{\rm Hom}\big{(}{\mathscr{H}}^{j},\,\Omega^{1}\otimes_{\mathscr{O}}{\mathscr{H}}^{j}\big{)}.$ Under the hypotheses of the proposition, we need to prove that
$\langle K_{\vartheta^{j}},\,\vartheta^{j}(\tau)\rangle=0,$
for any variation $\tau$.
The morphism $\tau$ can be extended
$C^{i}:{\mathscr{H}}^{i}\oplus{\mathscr{G}}^{i}\to\Omega^{1}\otimes_{\mathscr{O}}\big{(}{\mathscr{H}}^{i}\oplus{\mathscr{G}}^{i}\big{)},$
where
$C^{i}(a,\,b)=\begin{cases}(\tau(a),\,0),\;\;\text{if}\;\,i=j\\ (0,\,0),\;\;\text{if}\;\,i\neq j\end{cases}$
[…]
$0=\sum_{i}(-1)^{i}\langle K_{\vartheta^{i}},\,\vartheta^{i}(\zeta^{i})\rangle=(-1)^{j}\langle K_{{\vartheta}^{j}},\,\vartheta^{j}(\zeta^{j})\rangle=(-1)^{j}\langle K_{{\vartheta}^{j}},\,\vartheta^{j}(\tau)\rangle.$
This holds for any “variation” $\tau$ of $\vartheta^{j}$. That is, by (2.23), $\vartheta^{j}$ is a Yang-Mills field on ${\mathscr{H}}^{j}$. ∎
From Theorem 6 together with Proposition 14, it follows the corollary.
Corollary 4.
Let $({\mathscr{F}}^{\bullet},\,\delta^{\bullet})$ be a Hermitian $B$-brane on the Hodge manifold $Y,$ such that sheaves ${\mathscr{H}}^{i}({\mathscr{F}}^{\bullet})$ are reflexive. A gauge field $\psi$ on the brane is a Yang-Mills field iff $\vartheta^{i}$ is a Yang-Mills field on ${\mathscr{H}}^{i}({\mathscr{F}}^{\bullet})$ for all $i$.