### v2

# Holomorphic Yang-Mills fields on $B$-branes

Abstract.
Considering $B$-branes over a complex manifold $X$ as objects of the bounded derived category of coherent sheaves over $X$, we define holomorphic gauge fields on $B$-branes and introduce the Yang-Mills functional for these fields. These definitions extend well-known concepts in the context of vector bundles to the setting of $B$-branes.
For a given $B$-brane, we show that its Atiyah class is the obstruction to the existence of gauge fields. When $X$ is the variety of complete flags in a $3$-dimensional complex vector space, we prove that any $B$-brane over $X$ admits at most one holomorphic gauge field.
Furthermore, we establish that the set of Yang-Mills fields on a given $B$-brane, if nonempty, is in bijective correspondence with the points of an algebraic set defined by $m$ complex polynomials of degree less than four in $m$ indeterminates, where $m$ is the dimension of the space of morphisms from the brane to its tensor product with the sheaf of holomorphic one-forms.
Key words and phrases:
Yang-Mills fields, holomorphic connections, coherent reflexive sheaves
MSC 2020: 53C05, 58E15, 18G10
1. Introduction
In this article, we extend the well-known concepts of gauge field and Yang-Mills field on vector bundles to $B$-branes. From the mathematical point of view, a $B$-brane over a complex manifold $X$ is an object of $D^{b}(X)$, the bounded derived category of coherent analytic sheaves over $X$ [2, Sect. 5.4] [3, Sect. 5.3].
The simplest $B$-branes are the holomorphic vector bundles. A gauge field on a holomorphic vector bundle $V\to M,$ in mathematical terms a connection on $V,$ defines a derivation of sections of $V$ along vector fields on $M$ giving rise to sections not necessarily holomorphic. A holomorphic gauge field on $V,$ i. e. a holomorphic connection on $V$ [4], allows us to define a derivative of the holomorphic sections of $V$ along any “direction”, giving rise to holomorphic sections.
Not every holomorphic vector bundle admits a holomorphic connection. The obstruction for this existence is the Atiyah class of the bundle. The vanishing of this class is a necessary and sufficient condition for the existence of such connections on the vector bundle. Assuming that there is no obstruction to the existence of holomorphic connections, the set of these connections on a vector bundle over a compact manifold is a finite dimensional affine space, in contrast to the smooth case. This property is a consequence of the finite dimensionality of the coherent cohomology groups.
The two features discussed above concerning the holomorphic connections on vector bundles: the obstruction to its existence, and the finite dimension of the space of those fields, when it is not empty, are also present in the extension to general branes.
Connections on coherent sheaves. Given a coherent sheaf ${\mathscr{F}}$ over the compact analytic manifold $X,$ a holomorphic connection on ${\mathscr{F}}$ [17] determines isomorphisms between the stalks of ${\mathscr{F}}$ over “close” points of $X;$ i. e. identifications in the directions defined by the tangent vectors of the base.
The idea of being “infinitesimally close” can be formulated by means of the first infinitesimal neighborhood of the diagonal of $X$. In this setting, a connection on ${\mathscr{F}}$ can be regarded as a right inverse of the natural morphism ${\mathscr{J}}^{1}({\mathscr{F}})\to{\mathscr{F}}$, where ${\mathscr{J}}^{1}({\mathscr{F}})$ is the corresponding $1$-jet sheaf [26, Sect. 3]. This inverse determines a morphism of abelian sheaves $\nabla:{\mathscr{F}}\to\Omega^{1}({\mathscr{F}})=\Omega^{1}_{X}\otimes_{{\mathscr{O}}_{X}}{\mathscr{F}},$ which satisfies the Leibniz rule.
In this context, the obstruction to the existence of a holomorphic connection on the sheaf ${\mathscr{F}}$ is an element of the group ${\rm Ext}^{1}({\mathscr{F}},\,\Omega^{1}({\mathscr{F}})).$ Furthermore, when the set of holomorphic connections on ${\mathscr{F}}$ is nonempty, it is an affine space associated to the finite dimensional vector space ${\rm Hom}_{{\mathscr{O}}_{X}}({\mathscr{F}},\,\Omega^{1}({\mathscr{F}}))$.
Yang-Mills fields on sheaves. The reflexive sheaves might be thought as “vector bundles with singularities” [13, p. 121], and these singularities can in some way be avoided. Setting ${\mathcal{S}}$ for the singularity set of the reflexive sheaf ${\mathscr{F}}$, let us assume that on the locally free sheaf ${\mathscr{F}}|_{X\setminus{\mathcal{S}}}$ there is defined a Hermitian metric, we say that ${\mathscr{F}}$ is a Hermitian sheaf. If $X$ is a Kähler manifold, on the set of holomorphic connections over ${\mathscr{F}},$ one defines the Yang-Mills functional $\mathcal{YM}$. Essentially, the value of $\mathcal{YM}$ at a connection $\nabla$ is the squared norm $\lVert K_{\nabla}\rVert^{2}$ of the curvature of that connection [12, p. 417], [22, p. 44], [23, p. 357].
The stationary points of $\mathcal{YM}$ are the Yang-Mills fields on ${\mathscr{F}}$. The set of these points of this functional will be denoted by ${\sf YM}({\mathscr{F}})$. Denoting $m:={\rm dim}\,{\rm Hom}_{{\mathscr{O}}_{X}}({\mathscr{F}},\,\Omega^{1}({\mathscr{F}})),$ we will prove that if the reflexive sheaf ${\mathscr{F}}$ admits holomorphic connections, then the set ${\sf YM}({\mathscr{F}})$ is in bijective correspondence with the points of an algebraic subset of ${\mathbb{C}}^{m}$ defined by $m$ algebraic equations of degree $\leq 3$ (Theorem 3). Thus, in general, the set of Yang-Mills fields on ${\mathscr{F}}$ will be a finite set.
Yang-Mills fields on a $B$-brane. As we said, a $B$-brane on $X$ is a complex $({\mathscr{F}}^{\bullet},\,\delta^{\bullet})$ of analytic coherent sheaves on $X.$ According to the preceding paragraph, it is reasonable to define a holomorphic gauge field on this brane as an element $\psi\in{\rm Hom}_{D^{b}(X)}\big{(}{\mathscr{F}}^{\bullet},\,{\mathscr{J}}^{1}({\mathscr{F}}^{\bullet})\big{)}$ which lifts the identity on ${\mathscr{F}}^{\bullet}.$
When the manifold $X$ admits a stratification, where the strata satisfy certain properties, the space of holomorphic gauge fields on any $B$-brane on $X$ is a set with cardinal $\leq 1$ (Theorem 1). The case where $X$ is the projective space ${\mathbb{P}}^{n}$ has been considered in [27]. Here, we study the case where $X$ is the variety of complete flags in ${\mathbb{C}}^{3}$ (Theorem 2).
The homomorphic gauge field $\psi$ determines a unique morphism between each of the cohomology sheaves $\vartheta^{j}:{\mathscr{H}}^{j}({\mathscr{F}}^{\bullet})\to{\mathscr{H}}^{j}(\Omega^{1}({\mathscr{F}}^{\bullet})),$ which in fact is a holomorphic connection ${\mathscr{H}}^{j}({\mathscr{F}}^{\bullet})$.
When the cohomology sheaves ${\mathscr{H}}^{j}$ are Hermitian, we define the value of the Yang-Mills functional on the above gauge field $\psi$ as $\sum_{i}(-1)^{i}\lVert K_{\vartheta_{i}}\rVert^{2}.$ Thus, the Yang-Mills functional is a kind of Euler characteristic of the gauge field. Obviously, this definition of the Yang-Mills functional in the context of $B$-branes generalizes the one for coherent sheaves.
The gauge field $\psi$ is a Yang-Mills field if it is a stationary point of Yang-Mills functional. In Theorem 4, we generalize the result given in Theorem 3, about the cardinal of the set of Yang-Mills fields on sheaves, to a general brane ${\mathscr{F}}^{\bullet}$. We will prove that if the $\vartheta^{j}$ are Yang-Mills fields on the ${\mathscr{H}}^{j}$, so is $\psi$ on the brane (Proposition 11). Theorem 5 is a partial converse to Proposition 11.
The article is organized in two sections. In Section 2 are considered the holomorphic gauge fields on $B$-branes. Subsection 2.1 concerns with the holomorphic connections on a sheaf. In Subsection 2.2, we define the holomorphic gauge fields on a brane and discuss the obstruction to their existence. Subsection 2.3 discusses branes on a manifold which admits certain stratifications; in this subsection we prove Theorems 1 and 2.
Section 3 is devoted to the holomophic Yang-Mills fields. In Subsection 3.1, we introduce the Yang-Mills functional for holomorphic connections on reflexive sheaves and prove Theorem 3. The Yang-Mills functional for holomorphic gauge fields on a $B$-brane is defined in Subsection 3.2, along with a justification for this definition. In this subsection are also proved Proposition 11 and Theorems 4 and 5.
2. Holomorphic gauge fields on $B$-branes
As we mentioned in the Introduction, our purpose is to define holomorphic gauge fields on $B$-branes, extending the concept of holomorphic connection on vector bundles.
2.1. Holomorphic connections on a sheaf.
The definition of connection on a coherent sheaf, as it is introduced for example in [17], is reformulated in this section, so that it is suitable for translation to objects in the derived category of coherent sheaves.
As explained in the Introduction, given a coherent sheaf ${\mathscr{F}}$ over the complex manifold $X$, the existence of a homomorphic connection on ${\mathscr{F}}$ should define an isomorphism between the stalks of the sheaf at any two “infinitesimally close” points of $X.$ By means of the first infinitesimal neighborhood $X^{(1)}$ of diagonal of $X$ [11, p. 698], it is possible to formulate this idea in a suitable way also for algebraic varieties.
The outline of a new definition for a connection on a sheaf ${\mathscr{G}}$ over a complex variety $X$, inspired by an idea from crystalline cohomology, is as follows. If $R$ is a ${\mathbb{C}}$-algebra, ${\rm Hom}({\rm Spec}\,R,\,X)$ is the set of points of complex algebraic variety $X$ with values in $R$. Two points $x,y$ are infinitesimally close if the morphism $(x,\,y):{\rm Spec}\,R\to X\times X$ factorizes through the infinitesimal neighborhood $X^{(1)}.$ Hence, a connection on the sheaf ${\mathscr{G}}$ should define an identification of the pullbacks $x^{*}{\mathscr{G}}$ and $y^{*}{\mathscr{G}}$ for any two infinitesimally close points.
Therefore, following Deligne [7, p. 6], one defines a connection on the coherent ${\mathscr{O}}_{X}$-module ${\mathscr{F}}$ as an element of
(2.1) ${\rm Hom}_{X^{(1)}}(\pi_{1}^{*}{\mathscr{F}},\,\pi_{2}^{*}{\mathscr{F}}),$
which is the identity on $X$, where $\pi_{1},\pi_{2}:X^{(1)}\to X$ are the projections and $\pi_{i}^{*}$ the corresponding inverse image functor. By the adjunction isomorphism, a holomorphic connection on ${\mathscr{F}}$ is a morphism ${\mathscr{F}}\to\pi_{1*}\pi^{*}_{2}{\mathscr{F}}$ which is a right inverse of the projection $\pi:\pi_{1*}\pi^{*}_{2}{\mathscr{F}}\to{\mathscr{F}}.$
As $\pi_{1*}\pi^{*}_{2}{\mathscr{F}}$ is the first jet sheaf of the coherent sheaf ${\mathscr{F}}$, in the following paragraph we review the definition of the jet sheaf.
2.1.1. The first jet sheaf
Let $X$ be an analytic complex compact connected manifold. Let $i:\Delta\hookrightarrow X\times X$ be the embedding of the diagonal. As a closed submanifold, $\Delta$ is defined by an ideal $\mathscr{I}$ of $\hat{\mathscr{O}}:={\mathscr{O}}_{X\times X}.$ The first infinitesimal neighborhood of $\Delta$ is the following ringed space
$X^{(1)}=\Big{(}\Delta,\,{\mathscr{O}}_{X^{(1)}}:=\big{(}\hat{\mathscr{O}}/{\mathscr{I}}^{2}\big{)}|_{\Delta}\Big{)}$
[…]
Given ${\mathscr{F}}$ a left ${\mathscr{O}}_{X}$-module, its inverse image by $\pi_{2}$ is the left ${\mathscr{O}}_{X^{(1)}}$-module
$\pi_{2}^{*}({\mathscr{F}})={\mathscr{O}}_{X^{(1)}}\otimes_{\pi_{2}^{-1}{{\mathscr{O}}_{X}}}\pi_{2}^{-1}{\mathscr{F}}.$
And the first jet sheaf ${\mathscr{J}}^{1}({\mathscr{F}})$ of ${\mathscr{F}}$ is the left ${\mathscr{O}}$-module defined by (see [21, p. 505] [25, Sect. 2.4])
(2.2) ${\mathscr{J}}^{1}({\mathscr{F}})={\pi_{1}}_{*}\pi_{2}^{*}({\mathscr{F}}).$
We set $\Omega^{1}({\mathscr{F}}):=\Omega^{1}_{X}\otimes_{{\mathscr{O}}_{X}}{\mathscr{F}},$ where $\Omega^{1}_{X}$ is the sheaf of holomorphic $1$-forms on $X$. The first jet sheaf ${\mathscr{J}}^{1}({\mathscr{F}})$ is the abelian sheaf ${\mathscr{F}}\oplus\Omega^{1}({\mathscr{F}})$ endowed with the following left ${\mathscr{O}}_{X}$-module structure
(2.3) $f\cdot(\sigma\oplus\beta)=f\sigma\oplus(f\beta+df\otimes\sigma).$
One has the Atiyah exact sequence of ${\mathscr{O}}_{X}$-modules
[…]
where $\pi$ is the projection morphism.
Since $\pi_{1}^{*}$ is the left adjoint of ${\pi_{1}}_{*}$, one has
(2.5) ${\rm Hom}\,_{{\mathscr{O}}_{X}}\big{(}{\mathscr{F}},\,{\mathscr{J}}^{1}({\mathscr{F}})\big{)}={\rm Hom}\,_{{\mathscr{O}}_{X^{(1)}}}\big{(}\pi_{1}^{*}{\mathscr{F}},\,\pi_{2}^{*}{\mathscr{F}}\big{)}.$
We introduce a piece of notation that will be used later on. Let ${\sf Forg}$ denote the forgetful functor from the category of ${\mathscr{O}}_{X}$-modules to the category of abelian sheaves over $X.$ We set
$\boldsymbol{\mathscr{J}}^{1}({\mathscr{F}}):={\sf Forg}({\mathscr{F})}\oplus{\sf Forg}(\Omega^{1}({\mathscr{F}})).$
In this way ${\mathscr{J}}^{1}({\mathscr{F}})$ is the abelian sheaf $\boldsymbol{\mathscr{J}}^{1}({\mathscr{F}})$ endowed with the ${\mathscr{O}}_{X}$-module structure defined in (2.3). In short, we will write $\,{\mathscr{J}}^{1}({\mathscr{F}})={\mathscr{F}}\,\tilde{\oplus}\,\Omega^{1}({\mathscr{F}}).$
One has the morphism of ${\mathbb{C}}_{X}$-modules
(2.6) $\eta:{\mathscr{F}}\to\boldsymbol{\mathscr{J}}^{1}({\mathscr{F}})={\mathscr{F}}\oplus\Omega^{1}({\mathscr{F}}),\;\;\sigma\mapsto\sigma\oplus 0.$
And from (2.3), it follows
(2.7) $f\eta(\sigma)=\eta(f\sigma)+df\otimes\sigma,$
where $f\in\mathscr{O}_{X}$ is a function of the structure sheaf of $X$ and $\sigma$ a holomorphic section of ${\mathscr{F}}.$
Given $\phi\in{\rm Hom}_{\mathscr{O}_{X}}({\mathscr{F}},\,{\mathscr{J}}^{1}({\mathscr{F}}))$ a right inverse of $\pi$; that is, such that $\pi\circ\phi={\rm id}$. Then $p\circ(\phi-\eta)=0,$ where $p:\boldsymbol{\mathscr{J}}^{1}({\mathscr{F}})\to{\mathscr{F}}$ is the projection morphism. Thus $\phi-\eta$ factors uniquely through ${\rm Ker}(p)=\Omega^{1}({\mathscr{F}})$, defining the morphism $\nabla$ in the following diagram in the category of ${\mathbb{C}}_{X}$-modules.
[…]
From (2.7), together with the fact that $\phi$ is an ${\mathscr{O}}_{X}$-morphism, it follows that $\nabla$ satisfies the Leibniz rule $\nabla(f\sigma)=\partial f\otimes\sigma+f\nabla(\sigma).$ Therefore, one can give a new definition of holomorphic connection equivalent to the one given in [17].
Definition 1.
The holomorphic connections on the coherent ${\mathscr{O}}_{X}$-module ${\mathscr{F}}$ are the elements of the following set
(2.8) $\{\phi\in{\rm Hom}_{\mathscr{O}_{X}}({\mathscr{F}},\,{\mathscr{J}}^{1}({\mathscr{F}}))\;|\;\pi\circ\phi={\rm id}\}.$
That is, a holomorphic connection $\phi$ is a splitting of the corresponding Atiyah exact sequence. On the other hand, by (2.5), $\phi$ is an element of ${\rm Hom}_{{\mathscr{O}}_{X^{(1)}}}\big{(}\pi_{1}^{*}{\mathscr{F}},\,\pi_{2}^{*}{\mathscr{F}}\big{)}$ which is the identity on $X.$ Thus, we recover (2.1).
2.2. Gauge fields on a $B$-brane
Let $({\mathscr{F}}^{\bullet},\,\delta^{\bullet})$ be a $B$-brane on the complex manifold $X$; that is, ${\mathscr{F}}^{\bullet}$ is an object of the category $D^{b}(X)$, the bounded derived category of coherent sheaves over $X$. The corresponding first jet complex ${\mathscr{J}}^{1}({\mathscr{F}}^{\bullet})$ is defined by
${\mathscr{J}}^{1}({\mathscr{F}}^{\bullet}):=R\pi_{1*}L\pi_{2}^{*}{\mathscr{F}}^{\bullet}\simeq{\mathscr{O}}_{X^{(1)}}\otimes^{L}{\mathscr{F}}^{\bullet}.$
As ${\mathscr{O}}_{X^{(1)}}$ is the locally free module ${\mathscr{O}}_{X}\oplus\Omega^{1}_{X}$, then ${\mathscr{J}}^{1}({\mathscr{F}}^{\bullet})$ is the complex of abelian sheaves $\sf{Forg}({\mathscr{F}}^{\bullet})\,\oplus\,\sf{Forg}(\Omega^{1}({\mathscr{F}}^{\bullet})),$ with ${\mathscr{O}}_{X}$-structure given by
(2.9) $f\cdot(\sigma^{\bullet}\oplus\beta^{\bullet})=f\sigma^{\bullet}\oplus(df\otimes\sigma^{\bullet}+f\beta^{\bullet}).$
More precisely, one can consider the complex of abelian sheaves direct sum of ${\mathscr{F}}^{\bullet}$ and $\Omega^{1}({\mathscr{F}}^{\bullet}).$ According to the notation introduced at the end of Subsection 2.1.1, we write
$\,\boldsymbol{\mathscr{J}}^{1}({\mathscr{F}}^{\bullet}):=\big{(}{\sf Forg}({\mathscr{F}}^{\bullet})\oplus{\sf Forg}(\Omega^{1}({\mathscr{F}}^{\bullet})),\,\,\delta_{J}^{\bullet}\,\big{)},$
where $\delta_{J}^{\bullet}=\delta^{\bullet}\oplus(1\otimes\delta^{\bullet}).$ The complex $\boldsymbol{\mathscr{J}}^{1}({\mathscr{F}}^{\bullet})$ can be equipped with an ${\mathscr{O}}_{X}$-module structure (2.9). As $\delta$ is the coboundary operator of a complex ${\mathscr{O}}_{X}$-modules, then $\delta_{J}(f\cdot(-))=f\cdot(\delta_{J}(-)).$ Thus, ${\mathscr{J}}^{1}({\mathscr{F}}^{\bullet})$ is the complex $\boldsymbol{\mathscr{J}}^{1}({\mathscr{F}}^{\bullet})$ endowed with this ${\mathscr{O}}_{X}$-module structure (2.9).
One has the exact sequence of complexes of ${\mathscr{O}}_{X}$-modules
[…]
The morphisms in the derived category $D^{b}(X)$ defined by the arrows in the sequence (2.10) will also be denoted by $\iota$ and $\pi.$
According to Definition 1, we propose the following definition.
Definition 2.
A gauge field on ${\mathscr{F}}^{\bullet}$ is an element $\psi\in{\rm Hom}_{D^{b}(X)}({\mathscr{F}}^{\bullet},\,{\mathscr{J}}^{1}({\mathscr{F}}^{\bullet}))$, such that ${\pi\circ}\psi={\rm id}\in{\rm Hom}_{D^{b}(X)}({\mathscr{F}}^{\bullet},\,{\mathscr{F}}^{\bullet}).$
Remark 1.
The inclusion functor from the category $\mathbf{Coh}(X)$ of coherent sheaves on $X$ to its derived category $D^{b}(X)$ is fully faithful [9, p. 164]. Therefore, Definition 2 coincides with Definition 1 when it is applied to a brane consisting of only one nontrivial term ${\mathscr{F}}.$
Remark 2.
By the adjunction isomorphism
${\rm Hom}_{D^{b}(X^{(1)})}\big{(}L\pi_{1}^{*}{\mathscr{F}}^{\bullet},\,L\pi_{2}^{*}{\mathscr{F}}^{\bullet}\big{)}\simeq{\rm Hom}_{D^{b}(X)}({\mathscr{F}}^{\bullet},\,{\mathscr{J}}^{1}({\mathscr{F}}^{\bullet}))$
a gauge field on ${\mathscr{F}}^{\bullet}$ can be considered as an element of
${\rm Hom}_{D^{b}(X^{(1)})}\big{(}L\pi_{1}^{*}{\mathscr{F}}^{\bullet},\,L\pi_{2}^{*}{\mathscr{F}}^{\bullet}\big{)},$
which lifts the identity on $({\mathscr{F}}^{\bullet},\,\delta^{\bullet})$, in accordance with the observation at the beginning of Subsection 2.1 (see (2.1)).
Remark 3.
The gauge fields on ${\mathscr{F}}^{\bullet}$ are elements ${\rm Ext}^{0}({\mathscr{F}}^{\bullet},\,{\mathscr{J}}^{1}({\mathscr{F}}^{\bullet}));$ i.e., the are particular open strings between ${\mathscr{F}}^{\bullet}$ and ${\mathscr{J}}^{1}({\mathscr{F}}^{\bullet})$ with ghost number 00 [2, Sect. 5.2], [18].
The exact sequence (2.10) determines a distinguished triangle
[…]
in the category $D^{b}(X)$ [16, p. 46], [9, p. 157]. As ${\rm Hom}_{D^{b}(X)}({\mathscr{F}}^{\bullet},\,.\,)$ is a cohomological functor, it follows that
[…]
is an exact sequence.
Since $\mu(\psi)=\pi\circ\psi$, the existence of a holomorphic gauge field $\psi$ on ${\mathscr{F}}^{\bullet}$, is equivalent to $1_{{\mathscr{F}}^{\bullet}}\in{\rm Im}(\mu)={\rm Ker}(\nu)$. If $\psi$ and $\psi_{1}$ are gauge fields on the coherent sheaf ${\mathscr{F}^{\bullet}}$, then $\mu(\psi_{1}-\psi)=0$; i. e. $\psi_{1}-\psi\in{\rm Im}(\lambda)$. Hence, the set of holomorphic gauge fields on ${\mathscr{F}}^{\bullet}$, if nonempty, is an affine space with vector space ${\rm Hom}_{D^{b}(X)}({\mathscr{F}}^{\bullet},\Omega^{1}({\mathscr{F}}^{\bullet})).$ Thus, the space of holomorphic gauge fields on ${\mathscr{F}}^{\bullet}$, is a finite dimensional affine space.
The Atiyah class $a({\mathscr{F}}^{\bullet})$ of ${\mathscr{F}}^{\bullet}$ is the image of $1\in{\rm Hom}_{D^{b}(X)}({\mathscr{F}}^{\bullet},\,{\mathscr{F}}^{\bullet})$ in ${\rm Ext}^{1}({\mathscr{F}}^{\bullet},\,\Omega^{1}({\mathscr{F}}^{\bullet}))$. Hence, we have the following proposition.
Proposition 1.
The vanishing of $a({\mathscr{F}}^{\bullet})$ is a necessary and sufficient condition for the existence of gauge fields on the brane ${\mathscr{F}}^{\bullet}.$ Furthermore, the set of gauge fields on ${\mathscr{F}}^{\bullet},$ if is nonempty, is an affine space over the finite dimensional vector space ${\rm Ext}^{0}({\mathscr{F}}^{\bullet},\,\Omega^{1}({\mathscr{F}}^{\bullet})).$
2.3. $B$-branes on stratified manifolds
When the manifold $X$ admits a certain type of stratification, the derived category $D^{b}(X)$ is generated by a finite set of coherent sheaves, and the space of gauge fields on any brane over $X$ has cardinal $\leq 1.$ In this section, we will address this issue. In the next subsection, we review some properties of the generators of $D^{b}(X).$ Finally, we consider the particular case where $X$ is the variety of complete flags in ${\mathbb{C}}^{3},$ showing that the cardinality of the set of holomorphic gauge fields on any brane over $X$ is $<2.$
2.3.1. Generators of a derived category
Let ${\mathbf{C}}$ be an abelian category. If $A$ a complex in ${\mathbf{C}}$, as usual, we denote by ${A}[l]$, with $l\in{\mathbb{Z}}$, the complex ${A}$ shifted $l$ to the left. Let ${\sf E}$ be a finite set of objects of ${\mathbf{C}}$ which generates the derived category $D^{b}({\mathbf{C}}).$ Given $E,\tilde{E}$ elements of the generating set ${\sf E},$ let us consider morphisms $h$ between ${E}^{\prime}:={E}[l]$ and ${\tilde{E}}^{\prime}:={\tilde{E}}[l^{\prime}]$. We denote by ${\rm Cone}(h)={E}^{\prime}[1]\oplus{\tilde{E}}^{\prime}$ the mapping cone of $h$ [9, p. 154]. We define ${\sf E}^{(1)}$ the set obtained adding to ${\sf E}$ the elements of the form ${\rm Cone}(h)$. Hence, an element of ${\sf E}^{(1)}$ is a complex whose term at a position $p$ is either 00, or $E$, or a direct sum of ${E}\oplus\tilde{E},$ with $E,\tilde{E}\in{\sf E}.$
Repeating the process with the elements of ${\sf E}^{(1)}$ one obtains ${\sf E}^{(2)}$, etc. The objects of the triangulated subcategory generated by the family ${\sf E}$ are elements which belong to some ${\sf E}^{(m)}$. Therefore, an object of the triangulated subcategory of $D^{b}({\mathbf{C}})$ generated by ${\sf E}$ is a complex $(G,\,d)$, where ${G}^{p}$ is of the form
(2.11) $G^{p}=\bigoplus_{i\in I_{p}}E_{pi},$
with $E_{pi}\in{\sf E}$ and $i$ varying in a finite set $I_{p}$. (When $i$ “runs over the empty set”, the direct sum is taken to be 00). As ${\sf E}$ is a set generating $D^{b}({\mathbf{C}}),$ each object of this category is isomorphic to one of the form (2.11).
Given two complexes $(A,\,\partial_{A})$ and $(B,\,\partial_{B})$ in an additive category ${\mathbf{C}}$, the complex ${\rm Hom}^{\bullet}(A,\,B)$ is defined by (see [15, p. 17])
(2.12) ${\rm Hom}^{m}(A,\,B)=\prod_{q\in{\mathbb{Z}}}{\rm Hom}_{\mathbf{C}}\big{(}A^{q},\,B^{q+m}\big{)},$
with the differential $\delta_{H}$.
(2.13) $(\delta_{H}^{m}g)^{q}=\partial_{B}^{m+q}g^{q}+(-1)^{m+1}g^{q+1}\partial_{A}^{q},$
with $g\in{\rm Hom}^{m}(A,\,B).$
Denoting by ${\rm Com}({\mathbf{C}})$ the abelian category of complexes on ${\mathbf{C}}$ and fixed the object $A,$ one has the functor $F_{A}:{\rm Com}(\mathbf{C})\to{\rm Com}(\mathbf{Ab}),$ where $\mathbf{Ab}$ is the category of abelian groups, defined by
$F_{A}(B)={\rm Hom}^{\bullet}(A,\,B),\;\;F_{A}(f)^{m}(g)=\big{(}f^{m+q}g^{q}\big{)},$
$f$ a morphism from $(B,\,\partial_{B})\to(C,\,\partial_{C}).$ Assumed that ${\mathbf{C}}$ has enough injectives, as $F_{A}$ is a left exact functor, one has the corresponding derived functor $RF_{A}:D^{+}{(\mathbf{C}})\to D^{+}(\mathbf{Ab}).$
Since the 00th derived functor of a left exact functor is isomorphic to the functor itself, $F_{A}=R^{0}F_{A}={\rm Ext}^{0}(A,\,.\,).$ On the other hand, ${\rm Hom}_{D({\mathbf{C}})}\big{(}{A},\,B\big{)}={\rm Ext}^{0}(A,\,{B})$ [28, Sect. 10.7], so
(2.14) ${\rm Hom}_{D({\mathbf{C}})}\big{(}{A},\,B\big{)}={\rm Hom}^{\bullet}({A},\,{B}).$
One has the following lemma.
Lemma 1.
Let ${G}$ be the complex in the category ${\mathbf{C}}$ defined in (2.11) and $\hat{G}$ the complex ${\hat{G}}^{q}=\bigoplus_{j\in J_{q}}E_{qj}.$ If ${\rm Hom}_{\mathbf{C}}\big{(}E_{pi},\,E_{qj}\big{)}=0$ for all $p,q,i,j,$ then
${\rm Hom}_{D^{b}({\mathbf{C}})}\big{(}{G},\,\hat{G}\big{)}=0.$
Proof. From (2.12) together with the additivity of the functor ${\rm Hom}_{\mathbf{C}}(\,.\,,\,.\,),$ it follows
${\rm Hom}^{m}(G,\,\hat{G})=\prod_{p}\bigoplus_{ij}{\rm Hom}_{\mathbf{C}}(E_{pi},\,E_{(p+m)j})=0.$
By (2.14), ${\rm Hom}_{D^{b}({\mathbf{C}})}({G},\,\hat{G})=0.$
We shall deal with the particular case where the category ${\mathbf{C}}$ is a category $\mathbf{Coh}(X)$ of coherent sheaves. More precisely, let $X$ be a smooth complex $n$-dimensional variety, let $\mathbf{Qcoh}(X)$ denote the category of quasi-coherent sheaves on $X.$ One has the derived functor $RF_{A}:D^{+}(\mathbf{Qcoh}(X))\to D^{+}(\mathbf{Ab}).$ Since ${\rm Ext}^{i}(A,\,B)=0$ for $A,B$ complexes of coherent sheaves and $i\notin[0,\,n],$ we have the functor
$RF_{A}:D^{b}(X):=D^{b}(\mathbf{Coh}(X))\to D^{b}(\mathbf{Ab}).$
From Lemma 1 applied to the category $D^{b}(X)$, it follows the following corollary.
Corollary 1.
Let ${\mathscr{G}}^{p}=\bigoplus_{pi}{\mathscr{E}}_{pi}$ be an object of $D^{b}(X),$ where ${\mathscr{E}}_{pi}\in\mathbf{Coh}(X),$ and $\hat{\mathscr{G}}$ is an object of $D^{b}(X)$ category isomorphic the complex $\Omega^{1}_{X}\otimes_{{\mathscr{O}}_{X}}{\mathscr{G}}^{p},$ then ${\rm Hom}_{D^{b}(X)}\big{(}{{\mathscr{G}}},\,\hat{\mathscr{G}}\big{)}=0,$ if ${\rm Hom}({\mathscr{E}}_{ip},\,\Omega^{1}({\mathscr{E}}_{jq}))=0$ for all $p,q,i,j.$
Generators of the derived category $D^{b}(X)$.
Let $X$ be a locally Noetherian scheme, $i:Z\hookrightarrow X$ a closed subscheme of $X$ and $j:U\hookrightarrow X$ an open part of $X.$ Since $j$ is an open embedding, the functor proper inverse image $j^{!}$ is defined between the categories of modules $\mathbf{Mod}({\mathscr{O}}_{X})\to\mathbf{Mod}({\mathscr{O}}_{U}).$ One has the pair of adjoint functors $(j_{!}\dashv j^{!}=j^{*})$ and the pair $(i^{*}\dashv i_{*}=i_{!}).$
$\mathbf{Mod}({\mathscr{O}}_{X})\overset{j^{!}}{\to}\mathbf{Mod}({\mathscr{O}}_{U})\overset{j_{!}}{\longrightarrow}\mathbf{Mod}({\mathscr{O}}_{X})\overset{i^{*}}{\to}\mathbf{Mod}({\mathscr{O}}_{Z})\overset{i_{*}}{\longrightarrow}\mathbf{Mod}({\mathscr{O}}_{X}).$
For every sheaf ${\mathscr{F}}$ on $X,$ the corresponding adjunction morphisms give rise to the sequence of morphisms of ${\mathscr{O}}_{X}$-modules $\,j_{!}j^{!}{\mathscr{F}}\to{\mathscr{F}}\to i_{*}i^{*}{\mathscr{F}}.$ If ${\mathscr{F}}$ is a coherent ${\mathscr{O}}_{X}$-module, then sheaf $i^{*}{\mathscr{F}}$ is coherent. Since $i$ is a proper map, $i_{*}i^{*}{\mathscr{F}}$ is a coherent ${\mathscr{O}}_{X}$-module. As $j_{!}j^{!}{\mathscr{F}}$ is a subsheaf of ${\mathscr{F}}$, it is also coherent.
In particular, when $U=X\setminus Z$, for any ${\mathscr{F}}\in\mathbf{Coh}(X)$, one has the exact sequence of ${\mathscr{O}}_{X}$-modules $0\to j_{!}j^{!}{\mathscr{F}}\to{\mathscr{F}}\to i_{*}i^{*}{\mathscr{F}}\to 0$ If ${\mathscr{F}}^{\bullet}$ is an object of the derived category $D^{b}(X),$ then we have the distinguished triangle $Rj_{!}j^{!}{\mathscr{F}}^{\bullet}\to{\mathscr{F}}^{\bullet}\to Ri_{*}i^{*}{\mathscr{F}}^{\bullet}\overset{+1}{\longrightarrow},$ where $j^{!}{\mathscr{F}}^{\bullet}\in D^{b}(\mathbf{Coh}(U))$ and $i^{*}{\mathscr{F}}^{\bullet}\in D^{b}(\mathbf{Coh}(Z)).$ Hence, as $j_{!}$ and $i_{*}$ are exact functors, one has the following proposition.
Proposition 2.
If $\{g_{k}\}$ (resp. $\{g^{\prime}_{k^{\prime}}\}$) is a set generators of the derived category $D^{b}(\mathbf{Coh}(U))$ (resp. $D^{b}(\mathbf{Coh}(Z))$) as a triangulated category, then $\{j_{!}g_{k},\,i_{*}g^{\prime}_{k^{\prime}}\}$ is a set that generates $D^{b}(X).$
If $U={\rm Spec}\,R$, where $R$ is a Noetherian ring, the category $\mathbf{Coh}(U)$ of coherent sheaves on $U$ is equivalent to the category of finitely generated $R$-modules, with $R=\Gamma(U,\,{\mathscr{O}}_{U}).$ The functor $\Gamma(U,\,\cdot\,)$ defines the equivalence.
Moreover, if $R$ is a regular local ring, then its global dimension is finite, and any object $N$ of the category $\mathbf{Mod}_{f}(R)$ (of the $R$-modules of finite type) admits a finite free resolution
(2.15) $0\to S_{m}\overset{f_{m}}{\longrightarrow}S_{m-1}\to\dots\to S_{1}\overset{f_{1}}{\longrightarrow}S_{0}\to N\to 0,$
with $S_{a}$ a finite direct sum $\oplus R.$
Denoting by $\mathbf{T}$ the triangulated subcategory of $D^{b}(\mathbf{Mod}_{f}(R))$ generated by $R,$ then the $S_{a}$ are objects of $\mathbf{T}.$ On the other hand, if (2.15) reduces to $0\to S_{1}\overset{f_{1}}{\longrightarrow}S_{0}\to N\to 0$ in the category $\mathbf{Mod}_{f}(R),$ then one has the distinguished triangle
$S_{1}\overset{f_{1}}{\longrightarrow}S_{0}\to{\rm Cone}(f_{1})\overset{+1}{\longrightarrow}.$
Thus, $N$ belongs to the category $\mathbf{T}.$ In the general case where case $N$ has a resolution as (2.15), then one can consider the exact sequences
$0\to S_{m}\overset{f_{m}}{\longrightarrow}S_{m-1}\to{\rm Im}(f_{m-1})\to 0,$
and
$0\to{\rm Im}(f_{m-1})\to S_{m-2}\to\dots\to N\to 0$
By the previous result, ${\rm Im}(f_{m-1})$ belongs to $\mathbf{T}.$ By induction, it follows that $N$ is also an object of $\mathbf{T}.$ Hence $\mathbf{Mod}_{f}(R)\subset{\mathbf{T}}.$ As $D^{b}(\mathbf{Mod}_{f}(R))$ is generated by $\mathbf{Mod}_{f}(R),$ it follows that $R$ generates the derived category $D^{b}(\mathbf{Mod}_{f}(R)).$
From the equivalence between $\mathbf{Coh}(U)$ and $\mathbf{Mod}_{f}(R),$ one deduces the following proposition.
Proposition 3.
If $U$ is a Noetherian affine scheme, then ${\mathscr{O}}_{U}$ generates $D^{b}(U)$, the bounded derived category of coherent sheaves on $U$.
From Proposition 2, it follows the following corollary (see [5, Lem. 2.3.2]).
Corollary 2.
[…]
Let $X$ be a smooth a complex $n$-dimensional variety. Let $Z_{1}$ be an irreducible subvariety of $X$ of codimension $1,$ and $i_{1}:Z_{1}\hookrightarrow X$ the corresponding inclusion. We set $U_{1}:=X\setminus Z_{1}$ and $j_{1}:U_{1}\hookrightarrow X.$ Let us assume that $U_{1}$ is isomorphic to an affine variety. The exact sequence
$0\to j_{1!}j_{1}^{!}{\mathscr{O}}_{X}\to{\mathscr{O}}_{X}\to i_{1*}i^{*}_{1}{\mathscr{O}}_{X}\to 0,$
can be written $0\to{\mathscr{M}}_{1}\to{\mathscr{O}}_{X}\to{\mathscr{O}}_{Z_{1}}\to 0,$ where ${\mathscr{M}}_{1}$ is the invertible sheaf ${\mathscr{O}}_{X}(-Z_{1}),$ which is supported in $U_{1},$ i.e. the stalk $({\mathscr{M}}_{1})_{x}=0$ for all $x\in Z_{1}.$
As $j_{1}^{!}{\mathscr{O}}_{X}={\mathscr{O}}_{U_{1}}$, then ${\mathscr{M}}_{1}=j_{1!}{\mathscr{O}}_{U_{1}}$. From Corollary 2, it follows that the derived category $D^{b}(X)$ is generated by ${\mathscr{M}}_{1}$ together with the elements of $i_{1*}\{{\rm generators\;of}\,D^{b}(Z_{1})\}.$
If $Z_{2}\overset{i_{2}}{\hookrightarrow}Z_{1}$ is an irreducible hypersurface of $Z_{1}$ and $U_{2}:=Z_{1}\setminus Z_{2}$ is isomorphic to an affine variety, then we have an exact sequence
$0\to{\mathscr{M}}_{2}\to{\mathscr{O}}_{Z_{1}}\to i_{2*}i_{2}^{*}{\mathscr{O}}_{Z_{1}}\to 0,$
where ${\mathscr{M}}_{2}$ is the sheaf on $Z_{1},$ ${\mathscr{O}}_{{Z}_{1}}(-Z_{2}).$ Hence, $D^{b}(Z_{1})$ is generated by ${\mathscr{M}}_{2}$ together with $i_{2*}\{{\rm generators\,of}\,D^{b}(Z_{2})\}.$ Consequently, $D^{b}(X)$ is generated by
$\{{\mathscr{M}}_{1},\,i_{1*}{\mathscr{M}}_{2}\}\cup i_{1*}i_{2*}\big{(}\{{\rm generators\,of}\,D^{b}(Z_{2})\}\big{)}.$
Let us assume that there exists a tower $X=Z_{0}\supset Z_{1}\supset Z_{2}\supset\dots\supset Z_{n},$ where $Z_{k}\overset{i_{k}}{\hookrightarrow}Z_{k-1}$ is an irreducible hypersurface of $Z_{k-1}$ with $i=1,\dots,n.$ According to the above argument, ${\mathscr{M}}_{k}={\mathscr{O}}_{Z_{k-1}}(-Z_{k})$ is a sheaf supported on $U_{k}:=Z_{k-1}\setminus Z_{k}$. We set $\iota_{k}=i_{1}\circ\dots\circ i_{k}:Z_{k}\hookrightarrow X$ and $\iota_{0}:={\rm id}:Z_{0}\to X.$ For $k=1,\dots,n$ we denote ${\mathscr{L}}_{k}:=\iota_{k-1*}({\mathscr{M}}_{k}).$ We set ${\mathscr{L}}_{n+1}$ for denoting the sheaf on $X$ direct sum of the skyscraper sheaves on the points of the finite set $Z_{n}.$ Then ${\mathscr{L}}_{n+1}=\iota_{n*}\big{(}\{{\rm generators\,of}\,D^{b}(Z_{n})\}).$ We have proved the following proposition.
Proposition 4.
With the above notations, the derived category $D^{b}(X)$ is generated by the set of ${\mathscr{O}}_{X}$-modules $\,{\sf E}=\{{\mathscr{L}}_{1},\dots,{\mathscr{L}}_{n},\,{\mathscr{L}}_{n+1}\}.$
Remark 4.
When $j\neq k$ the intersection $U_{j}\cap U_{k}=\emptyset.$ Hence, ${\mathscr{L}}_{j}$ and $\Omega^{1}({\mathscr{L}}_{k})$ have disjoint supports; thus,
(2.16) ${\rm Hom}\big{(}{\mathscr{L}}_{j},\,\Omega^{1}_{X}\otimes_{{\mathscr{O}}_{X}}{\mathscr{L}}_{k}\big{)}=0.$
Lemma 2.
With the notation introduced above, ${\rm Hom}({\mathscr{L}}_{k},\,{\mathscr{L}}_{k}\otimes_{{\mathscr{O}}_{X}}\Omega^{1}_{X})=H^{1,0}(Z_{k}),$ for $k=1,\dots,n.$
Proof. We set in this proof ${\mathscr{L}}:={\mathscr{L}}_{k},$ ${\mathscr{M}}:={\mathscr{M}}_{k},$ and $\iota:=\iota_{k-1}$ for the inclusion $Z:=Z_{k-1}\hookrightarrow X;$ thus ${\mathscr{L}}=\iota_{*}{\mathscr{M}}.$ We will also delete the subscript $X$ in the sheaves ${\mathscr{O}}_{X}=:{\mathscr{O}}$ and $\Omega^{1}_{X}=:\Omega^{1}.$ By ${\mathscr{H}om}\,_{\mathscr{R}}(\,.\,,\,.\,)$ we denote the sheaf of linear homomorphisms between the corresponding ${\mathscr{R}}$-modules.
By [16, Prop. 2.5.13] and [16, Cor. 2.3.4]
${\mathscr{H}om}\,_{\mathscr{O}}({\mathscr{L}},\,{\mathscr{L}}\otimes_{\mathscr{O}}\Omega^{1})\simeq\iota_{*}{\mathscr{H}om}\,_{{\mathscr{O}}_{Z}}({\mathscr{M}},\,{\mathscr{M}}\otimes_{{\mathscr{O}}_{Z}}\Omega^{1}_{Z}).$
By [16, Prop. 2.5.13] and [16, Cor. 2.3.4]
${\mathscr{H}om}\,_{\mathscr{O}}({\mathscr{L}},\,{\mathscr{L}}\otimes_{\mathscr{O}}\Omega^{1})\simeq\iota_{*}{\mathscr{H}om}\,_{{\mathscr{O}}_{Z}}({\mathscr{M}},\,{\mathscr{M}}\otimes_{{\mathscr{O}}_{Z}}\Omega^{1}_{Z}).$
Hence, as ${\mathscr{M}}$ is an invertible ${\mathscr{O}}_{Z}$-module,
${\rm Hom}({\mathscr{L}},\,{\mathscr{L}}\otimes_{\mathscr{O}}\Omega^{1})\simeq\Gamma\big{(}X,\,\iota_{*}{\mathscr{H}om}\,_{{\mathscr{O}}_{Z}}({\mathscr{M}},\,{\mathscr{M}}\otimes_{{\mathscr{O}}_{Z}}\Omega^{1}_{Z})\big{)}=$
$\Gamma\big{(}Z,\,{\mathscr{H}om}\,_{{\mathscr{O}}_{Z}}({\mathscr{M}},\,{\mathscr{M}}\otimes_{{\mathscr{O}}_{Z}}\Omega^{1}_{Z})\big{)}\simeq{\rm Hom}_{{\mathscr{O}}_{Z}}({\mathscr{O}}_{Z},\Omega^{1}_{Z})=H^{1,0}(Z).$
Theorem 1.
Let $X$ be an $n$-dimensional smooth variety and suppose that:
  1. (1)
    There exists $X=Z_{0}\supset Z_{1}\supset\dots\supset Z_{n},$ where $Z_{i+1}$ is a smooth irreducible subvariety of $Z_{i},$ with ${\rm codim}_{Z_{i}}Z_{i+1}=1$.
  2. (2)
    $U_{j}=Z_{j-1}\setminus Z_{i}$ is isomorphic to an affine variety.
If $H^{1,0}(Z_{j})=0$ for all $j,$ then the set of holomorphic gauge fields on any $B$-brane over $X$ has cardinal $\leq 1.$
Proof. From Lemma 2 together with Remark 4, it follows that ${\rm Hom}({\mathscr{L}}_{j},\,\Omega^{1}({\mathscr{L}}_{k}))=0,$ for all $j,k.$ Given a brane ${\mathscr{F}}^{\bullet}$ over $X,$ from Corollary 1 and Proposition 4, we deduce that ${\rm Hom}_{D^{b}(X)}({\mathscr{F}}^{\bullet},\,\Omega^{1}({\mathscr{F}}^{\bullet}))=0.$ Now the theorem is consequence of Proposition 1. ∎
Using this theorem when $X$ is the projective space, it has been proven in [27] that the set of holomorphic gauge fields on any $B$-brane on ${\mathbb{P}}^{n}$ has cardinal $\leq 1.$
2.3.2. Flag varieties
In this section, we will first recall some properties of the Schubert varieties. We denote by $X:={\rm GL}(n,\,{\mathbb{C}})/B$, where $B$ is the Borel subgroup of ${\rm GL}(n,\,{\mathbb{C}})$ consisting of the upper triangular matrices. $X$ is the manifold of complete flags in ${\mathbb{C}}^{n}.$
Let $T$ denote the maximal torus of ${\rm GL}(n,\,{\mathbb{C}})$ formed by the diagonal matrices. We set $W:=N(T)/T$, where $N(T)$ is the normalizer of $T$ in the group ${\rm GL}(n,\,{\mathbb{C}}).$ As it is well-known, the Weyl group $W$ is isomorphic to the symmetric group $S_{n}$. The Bruhat partial order [8, Sect. 10.5] in $S_{n}$ will be denoted by $\leq$.
Given $w\in W,$ we set $C^{0}_{w}$ for the Schubert cell of $X$ associated to $w$ [8]. The closure (in Zariski topology) of this cell is the corresponding Schubert subvariety of $X$
(2.17) $C_{w}:=\bigcup_{v\leq w}C_{v}^{0}.$
The Poincaré polynomial of $C_{w}$ given in [1] shows that the Betti number $b_{1}(C_{w})$ vanishes, if $C_{w}$ is any smooth Schubert variety.
Theorem 1 of [20] gives a criterion to determine whether a Schubert variety in a flag variety is singular. Let $w$ be the permutation $w=(t_{1},\dots,t_{n})$. The variety $C_{w}$ is singular iff there are integers $i,j,k,l$ with $1\leq i<j<k<l\leq n$ such that, either
$t_{k}<t_{l}<t_{i}<t_{j}\quad\text{or}\quad t_{l}<t_{j}<t_{k}<t_{i}.$
In particular, all the Schubert varieties of $X={\rm GL}(3,\,{\mathbb{C}})/B$ are smooth. Thus, if $C_{w}$ is such a variety, as the $b_{1}(C_{w})=0$, the Hodge numbers $h^{1,0}(C_{w})$ and $h^{0,1}(C_{w})$ vanish
(2.18) $h^{1,0}(C_{w})=h^{0,1}(C_{w})=0.$
From now on in this section, $X$ will denote the variety of complete flags in ${\mathbb{C}}^{3}.$ The Bruhat order in the group $S_{3}$ is showed in the following diagram [6].
${w_{1}=(321)}$${w_{3}=(231)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${w_{2}=(312)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${w_{5}=(213)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${w_{4}=(132)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${w_{6}=(123)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
Hence, there are only two pairs of permutations that are not comparable in the Bruhat order, namely, $(w_{2},\,w_{3})$ and $(w_{4},\,w_{5}).$ For any other pair $(w,\,w^{\prime})$ the intersection $C_{w}\cap C_{w^{\prime}}=C_{v}$, where $v={\rm min}\{w,\,w^{\prime}\}.$
The length $l(w)$ of the permutation $w$ is defined by
$l(w)=\#\{a<b\,|\,w(a)>w(b)\}.$
We will set $C^{0}_{k}$ for the cell $C^{0}_{w_{k}}$ and $C_{k}:=C_{w_{k}}.$ The cell $C^{0}_{w}$ is an affine variety isomorphic to ${\mathbb{C}}^{l(\omega)}$ [8, p. 157]. Hence, ${\rm dim}\,C_{4}=1={\rm dim}\,C_{5}$ and ${\rm dim}\,C_{2}=2={\rm dim}\,C_{3}.$ Furthermore, $X$ is the disjoint union of the cells
$X=\bigsqcup_{k=1}^{6}C^{0}_{k}.$
On the other hand, the inclusion relations among the Schubert subvarieties of $X$ can be obtained from (2.17) by means of the Bruhat order depicted above.
We consider the following tower of subspaces of $X$
$X=Z_{0}=C_{1}\overset{i_{1}}{\hookleftarrow}Z_{1}=C_{2}\cup C_{3}\overset{i_{2}}{\hookleftarrow}Z_{2}=C_{4}\cup C_{5}\overset{i_{3}}{\hookleftarrow}Z_{3}=C_{6}$
That is, $Z_{i}$ is the union of the Schubert varieties of dimension $3-i.$ We set $U_{i}:=Z_{i-1}\setminus Z_{i}$. From (2.17), it follows the expression of each $U_{i}$ in terms of Schubert cells. One has the the following inclusion relations
$Z_{0}\overset{j_{1}}{\hookleftarrow}U_{1}=C^{0}_{1},\;\;\;Z_{1}\overset{j_{2}}{\hookleftarrow}U_{2}=C^{0}_{2}\cup C^{0}_{3},\;\;\;Z_{2}\overset{j_{3}}{\hookleftarrow}U_{3}=C^{0}_{4}\cup C^{0}_{5}.$
A family of generators for $D^{b}(X)$ will then be constructed by successive applications of Corollary 2.
$U_{1}$ open part of $Z_{0}=X.$ In the flag variety $X$, we have the divisor $Z_{1}=X\setminus C^{0}_{1}=C_{2}\cup C_{3},$ and the partition of $X$ defined by the affine variety $U_{1}=C^{0}_{1}$ and $Z_{1}.$ We denote by $j_{1}$ and $i_{1}$ the respective inclusions
$U_{1}\overset{j_{1}}{\hookrightarrow}X\overset{i_{1}}{\hookleftarrow}Z_{1}.$
On the other hand, we have the exact sequence of ${\mathscr{O}}_{X}$-modules
$0\to{\mathscr{M}}_{1}:=j_{1!}j^{!}_{1}{\mathscr{O}}_{X}\to{\mathscr{O}}_{X}\to i_{1*}i^{*}_{1}{\mathscr{O}}_{X}\to 0$
The nontrivial term on the left satisfies
  1. (1)
    ${\mathscr{M}}_{1}$ is the invertible sheaf ${\mathscr{O}}_{X}(-C_{2}-C_{3})$ associated to the divisor $-C_{2}-C_{3}.$
  2. (2)
    ${\rm Supp}\,{\mathscr{M}}_{1}\subset U_{1}.$
  3. (3)
    As $j^{!}_{1}{\mathscr{O}}_{X}={\mathscr{O}}_{U_{1}}$, from Corollary 2, it follows that $D^{b}(X)$ is generated by ${\mathscr{M}}_{1}$ together with $i_{1*}\{{\rm generators\;of}D^{b}(Z_{1})\}.$ We write
    (2.19) $D^{b}(X)=\langle{\mathscr{M}}_{1},\,i_{1*}D^{b}(Z_{1})\rangle.$
$U_{2}$ open subset of $Z_{1}.$ Next we will consider a divisor $Z_{2}=C_{4}+C_{5}$ of $Z_{1}$ and we apply to it the above analysis. We will first state the following simple lemma.
Lemma 3.
Let $Y$ be a complex variety, and $Y_{1},\,Y_{2}$ disjoint affine subvarieties of $Y,$ then the derived category $D^{b}(Y)$ is generated by $\{\alpha_{1!}{\mathscr{O}}_{Y_{1}},\,\alpha_{2!}{\mathscr{O}}_{Y_{2}}\},$ where $\alpha_{k}:Y_{k}\hookrightarrow Y$ denotes the corresponding embedding.
Proof. Given ${\mathscr{F}}\in\mathbf{Coh}(Y),$ one has the coherent sheaves ${\mathscr{F}}_{k}:=\alpha_{k!}\alpha_{k}^{!}{\mathscr{F}}$ and the exact sequence of ${\mathscr{O}}_{Y}$-modules [16, Prop. 2.3.6 (vii)]
$0\to{\mathscr{F}}_{Y_{1}\cap Y_{2}}=0\to{\mathscr{F}}_{1}\oplus{\mathscr{F}}_{2}\to{\mathscr{F}}\to 0$
By Proposition 3, $D(Y_{k})$ is generated by ${\mathscr{O}}_{Y_{k}}.$ From the above exact sequence, it follows that the sheaf ${\mathscr{F}}$ belongs to the subcategory of $D^{b}(Y)$ generated by $\alpha_{1!}{\mathscr{O}}_{Y_{1}},\,$ $\alpha_{2!}{\mathscr{O}}_{Y_{2}}.$
Going back to the variety $X$. Next we will consider a divisor $Z_{2}=C_{4}+C_{5}$ of $Z_{1}.$ Then $U_{2}:=Z_{1}\setminus Z_{2}=C_{2}^{0}\cup C_{3}^{0}.$ For $k=2,3$ we have the inclusions
$C^{0}_{k}\overset{\alpha_{k}}{\hookrightarrow}U_{2}\overset{j_{2}}{\hookrightarrow}Z_{1}\overset{i_{2}}{\hookleftarrow}Z_{2}.$
As $C^{0}_{2}\cap C^{0}_{3}=\emptyset,$ from Lemma 3 together with Corollary 2, it follows that the set $\{\alpha_{k!}{\mathscr{O}}_{C^{0}_{k}}\}_{k=2,3}$ is a set of generators of $D^{b}(U_{2}).$
For $k=2,3$ the ${\mathscr{O}}_{Z_{1}}$-module $(j_{2}\alpha_{k})_{!}{\mathscr{O}}_{C^{0}_{k}},$ will be denoted ${\mathscr{M}}_{k}.$
  1. (1)
[…]
  2. (2)
    ${\rm Supp}\,{\mathscr{M}}_{k}\subset C^{0}_{k}.$
  3. (3)
    From Corollary 2, we conclude
    $D^{b}(Z_{1})=\langle{\mathscr{M}}_{2},\,{\mathscr{M}}_{3},\,i_{2*}D^{b}(Z_{2})\rangle.$
From (2.19), it follows
(2.20) $D^{b}(X)=\langle{\mathscr{M}}_{1},\,i_{1*}{\mathscr{M}}_{2},\,i_{1*}{\mathscr{M}}_{3},\,i_{1*}i_{2*}D^{b}(Z_{2})\rangle.$
$U_{3}$ open subset of $Z_{2}.$ In a similar way, we define $U_{2}=C^{0}_{4}\cup C^{0}_{5}$ and $Z_{3}=Z_{2}\setminus U_{2}=C_{6}.$ For $k=4,5$ one has the embeddings
$C^{0}_{k}\overset{\alpha_{k}}{\hookrightarrow}U_{3}\overset{j_{3}}{\hookrightarrow}Z_{2}\overset{i_{3}}{\hookleftarrow}Z_{3}.$
The ${\mathscr{O}}_{Z_{2}}$-module $(j_{3}\alpha_{k})_{!}{\mathscr{O}}_{C^{0}_{k}},$ will be denoted ${\mathscr{M}}_{k}.$
  1. (1)
[…]
  2. (2)
    ${\rm Supp}\,{\mathscr{M}}_{k}\subset C^{0}_{k}.$
  3. (3)
    By Corollary 2, $D^{b}(Z_{2})=\langle{\mathscr{M}}_{3},\,{\mathscr{M}}_{4},\,i_{3*}D^{b}(Z_{3})\rangle.$
From (2.20), it follows
(2.21) $D^{b}(X)=\langle{\mathscr{M}}_{1},\,i_{1*}{\mathscr{M}}_{2},\,i_{1*}{\mathscr{M}}_{3},\,(i_{1}i_{2})_{*}{\mathscr{M}}_{4},\,(i_{1}i_{2})_{*}{\mathscr{M}}_{5},\,(i_{1}i_{2}i_{3})_{*}D^{b}(Z_{3})\rangle.$
For the sake of simplicity in the notation we set
${\mathscr{L}}_{1}:={\mathscr{M}}_{1},\,{\mathscr{L}}_{2}:=i_{1*}{\mathscr{M}}_{2},\,{\mathscr{L}}_{3}:=i_{1*}{\mathscr{M}}_{3},\,{\mathscr{L}}_{4}:=(i_{1}i_{2})_{*}{\mathscr{M}}_{4},\,{\mathscr{L}}_{5}:=(i_{1}i_{2})_{*}{\mathscr{M}}_{5}.$
The point $Z_{3}.$ ${\mathscr{L}}_{6}$ will denote the skyscraper sheaf on $X$ at the point $C_{6}.$ Thus,
Proposition 5.
The derived category $D^{b}(X)$ is generated by the set $\{{\mathscr{L}}_{1},\dots,{\mathscr{L}}_{6}\}.$
Since ${\rm Supp}({\mathscr{L}}_{k})\subset C^{0}_{k}$ and $C^{0}_{r}\cap C^{0}_{s}=\emptyset.$ for $r\neq s.$ Then
(2.22) ${\rm Hom}_{{\mathscr{O}}_{X}}\big{(}{\mathscr{L}}_{r},\,{\mathscr{L}}_{s}\otimes_{{\mathscr{O}}_{X}}\Omega^{1}_{X}\big{)}=0.$
Proposition 6.
${\rm Hom}_{{\mathscr{O}}_{X}}({\mathscr{L}}_{r},\,{\mathscr{L}}_{r}\otimes_{{\mathscr{O}}_{X}}\Omega^{1}_{X})=0,\,$ for $r=1,\dots,6.$
Proof. As in the proof of Lemma 2, ${\rm Hom}_{{\mathscr{O}}_{X}}({\mathscr{L}}_{1},\,{\mathscr{L}}_{1}\otimes_{{\mathscr{O}}_{X}}\Omega^{1}_{X})=H^{1,0}(Z_{1}).$ On the other hand, $Z_{1}=C_{2}\cup C_{3}$, and by (2.18) $H^{1,0}(C_{2})=H^{1,0}(C_{3})=0.$ Thus, $H^{0}(Z_{1},\,\Omega^{1}_{Z_{1}})=0.$ That is, the proposition is proved for the case $r=1.$ The proof of the remaining cases is analogous. ∎
Theorem 2.
The set of holomorphic gauge fields on any $B$-brane over the variety of complete flags in ${\mathbb{C}}^{3}$ is $\leq 1$.
Proof. An object of $D^{b}(X)$ is isomorphic to one of the triangulated category generated by the family ${\mathscr{L}}_{1},\dots,{\mathscr{L}}_{6}$ in $D^{b}(X).$ Let $({G}^{\bullet},\,d^{\bullet})$ be an object of this triangulated category. Then ${G}^{p}$ is a sheaf of the form $G^{p}=\oplus_{i}{\mathscr{L}}_{pi},$ with $pi\in\{1,\dots,6\}.$ From Corollary 1 together with (2.22) and Proposition 6, it follows ${\rm Hom}_{D^{b}(X)}\big{(}{G}^{\bullet},\,\Omega^{1}(G^{\bullet})\big{)}=0.$ The theorem is a consequence of Proposition 1. ∎
3. Holomorphic Yang-Mills fields
3.1. Hermitian sheaves
The fiber at $x\in X$ of a coherent ${\mathscr{O}}_{X}$-module ${\mathscr{G}}$ will be denoted by ${\mathscr{G}}_{(x)}:={\mathscr{G}}_{x}/{\mathfrak{m}}_{x}{\mathscr{G}}_{x}$, where ${\mathfrak{m}}_{x}$ is the maximal ideal of ${\mathscr{O}}_{x}$. If ${\mathcal{Z}}$ is a section of ${\mathscr{G}}$ the corresponding vector in ${\mathscr{G}}_{(x)}$ is denoted by $Z(x)$. On the other hand, the singular set ${\mathcal{S}}$ of ${\mathscr{G}}$ is a closed analytic subset of $X$ whose codimension is greater than or equal to $1$ [19, Chap. V, Thm. 5.8]. Moreover, ${\mathscr{G}}$ is locally free on $X\setminus{\mathcal{S}}$. We set ${G}$ for the vector bundle over $X\setminus{\mathcal{S}}$, with fibers $G(x):={\mathscr{G}}_{(x)}$, determined by the locally free sheaf ${\mathscr{G}}|_{X\setminus{\mathcal{S}}}$.
The following definition generalizes the given in [29, Chap. III, Sect. 1] for locally free sheaves.
Definition 3.
A Hermitian metric on the coherent sheaf ${\mathscr{G}}$ is a set $\{\langle\,,\,\rangle_{x}\}_{x\in X}$ of Hermitian metrics on the fibers of ${\mathscr{G}}$, satisfying the following condition: Given ${\mathcal{Z}}_{1},{\mathcal{Z}}_{2}$ two sections of ${\mathscr{G}}$ on an open $U$ of $X,$ the map
${\mathcal{Z}}_{1}\cdot{\mathcal{Z}}_{2}:x\in U\mapsto\langle Z_{1}(x),\,Z_{2}(x)\rangle_{x}\in{\mathbb{C}}$
is bounded, and its restriction to $U\setminus{\mathcal{S}}$ is $C^{\infty}.$ A sheaf endowed with a Hermitian metric is called a Hermitian sheaf.
Given ${\mathcal{Z}}_{1},{\mathcal{Z}}_{2}\in\Gamma(X,\,{\mathscr{G}}),\,$ ${\mathcal{Z}}_{1}\cdot{\mathcal{Z}}_{2}$ is a bounded map with possible discontinuities in the singular locus. If $X$ is a Kähler manifold, the function ${\mathcal{Z}}_{1}\cdot{\mathcal{Z}}_{2}$ may be non-integrable on $X$ with respect to the volume form defined by the Kähler metric.
If ${\mathscr{G}}$ is a torsion-free coherent sheaf on the Kähler manifold $X,$ then ${\rm codim}\,{\mathcal{S}}\geq 2$ [19, Cap. V, Cor. 5.15]. Hence, there exists an analytic closed subset ${\mathcal{W}}$ of $X,$ containing ${\mathcal{S}},$ and such that $1\leq{\rm codim}\,{\mathcal{W}}<{\rm codim}\,{\mathcal{S}}.$ Let ${\mathcal{N}}$ be a tubular neighborhood of ${\mathcal{S}}$ in ${\mathcal{W}}.$ Then
  1. (1)
    ${\mathcal{Z}}_{1}\cdot{\mathcal{Z}}_{2}$ is a continuous map on the closed space $X\setminus{\mathcal{N}},$ which is disjoint with the singularity set ${\mathcal{S}}.$
  2. (2)
    ${\mathcal{W}},$ as a subset of $X$ with codimension $\geq 1,$ has measure zero with respect to the differential form ${\rm d}{\rm vol}.$
Thus, the following integral is well-defined
(3.1) $({\mathcal{Z}}_{1},\,{\mathcal{Z}}_{2}):=\int_{X\setminus{\mathcal{N}}}{\mathcal{Z}}_{1}\cdot{\mathcal{Z}}_{2}\,{\rm d}{\rm vol}.$
From the boundedness of ${\mathcal{Z}}_{1}\cdot{\mathcal{Z}}_{2}$ together with $(2),$ it follows that this definition does not depend on the choices of the analytic closed set ${\mathcal{W}}$ and the neighborhood ${\mathcal{N}}.$
Remark 5.
[…]
3.1.1. Yang-Mills functional
Given $\nabla$ a holomorphic connection on the coherent sheaf $\mathscr{F}.$ Then the corresponding operator $\nabla$ defines a morphism of ${\mathbb{C}}_{X}$-modules $\nabla^{(k)}:\Omega^{k}({\mathscr{F}}){\to}\Omega^{k+1}({\mathscr{F}})$ in the usual way. The composition ${\mathcal{K}}_{\nabla}:=\nabla^{(1)}\circ\nabla:{\mathscr{F}}\to\Omega^{2}({\mathscr{F}})$ is the curvature of $\nabla$; and as it is well-known
(3.2) ${\mathcal{K}}_{\nabla}\in{\rm Hom}({\mathscr{F}},\,\Omega^{2}({\mathscr{F}}))=\Gamma(X,\,{\mathscr{H}om}_{{\mathscr{O}}_{X}}({\mathscr{F}},\,\Omega^{2}({\mathscr{F}}))).$
The connection is said to be flat if ${\mathcal{K}}_{\nabla}=0$.
For each $x\in X,$ we denote by $\alpha_{x}$ and $\lambda_{x}$ the natural morphisms
$\big{(}{\mathscr{H}om}_{{\mathscr{O}}_{X}}({\mathscr{F}},\,\Omega^{k}({\mathscr{F}}))\big{)}_{x}\overset{\alpha_{x}}{\to}{\rm Hom}_{{\mathscr{O}}_{x}}\big{(}{\mathscr{F}}_{x},\,\Omega^{k}_{x}\otimes_{{\mathscr{O}}_{x}}{\mathscr{F}}_{x}\big{)}\overset{\lambda_{x}}{\leftarrow}\Omega^{k}_{x}\otimes_{{\mathscr{O}}_{x}}{\rm End}_{{\mathscr{O}}_{x}}({\mathscr{F}}_{x}).$
As ${\mathscr{F}}$ is coherent, $\alpha_{x}$ is isomorphism [10, p. 239]. Furthermore, if ${\mathscr{F}}_{x}$ is free, then $\lambda_{x}$ is bijective.
If ${\mathcal{S}}$ is the singularity set of ${\mathscr{F}},$ it is also the singular locus of the sheaf ${\mathscr{H}om}_{{\mathscr{O}}_{X}}({\mathscr{F}},\,\Omega^{2}({\mathscr{F}})).$ Hence, for each point $x$ outside of ${\mathcal{S}},$ the fiber of ${\mathscr{H}om}_{{\mathscr{O}}_{X}}({\mathscr{F}},\,\Omega^{k}({\mathscr{F}}))$ at $x$ can be identified with the vector space $\Omega^{k}{(x)}\otimes{\rm End}({F}{(x)}).$ According to (3.2), the curvature ${\mathcal{K}}_{\nabla}$ of a holomorphic connection $\nabla$ determines the vector
(3.3) $K_{\nabla}(x)\in\Omega^{2}(x)\otimes{\rm End}(F(x))$
for each $x\in X\setminus{\mathcal{S}}$. That is, $K_{\nabla}$ is a $2$-form ${\rm End}(F)$-valued.
Let us assume that ${\mathscr{F}}$ is a Hermitian torsion-free sheaf on the Kähler manifold $X,$ according to (3.1), one defines
(3.4) $\lVert{\mathcal{K}}_{\nabla}\rVert^{2}=({\mathcal{K}}_{\nabla},\,{\mathcal{K}}_{\nabla})=\int_{X\setminus{\mathcal{N}}}{\mathcal{K}}_{\nabla}\cdot{\mathcal{K}}_{\nabla}\,{\rm d}{\rm vol}=\int_{X\setminus{\mathcal{N}}}|K_{\nabla}\wedge\star\,K_{\nabla}|,$
[…]
More concretely, if locally $K_{\nabla}$ can be expressed as $\alpha\otimes A$, with $\alpha$ a $2$-form and $A$ a local section of ${\rm End}(F)$, then the integrand in (3.4) is $(\alpha\wedge\star\alpha)\langle A\circ A\rangle$. In a local unitary frame of ${\rm End}(F)$, if the connection is compatilble with the metric, the matrix $\check{A}$ associated to $A$ is antihermitian and $\langle A\circ A\rangle=(-1/2){\rm tr}(\check{A}\check{A})$. That is,
(3.5) $|K_{\nabla}\wedge\star\,K_{\nabla}|=-(1/2){\rm tr}\big{(}K_{\nabla}\wedge\star\,K_{\nabla}\big{)}.$
On the space of holomorphic gauge fields on the torsion-free sheaf ${\mathscr{F}}$ one can define the following map $\,\mathcal{YM}:\nabla\in{\rm Hom}({\mathscr{F}},\,\Omega^{1}({\mathscr{F}}))\mapsto\lVert{\mathcal{K}}_{\nabla}\rVert^{2}.\,$ It is called the Yang-Mills’ functional. The $\nabla$ on which this functional takes a stationary value are the holomorphic Yang-Mills fields.
3.1.2. Yang-Mills fields on reflexive sheaves
If ${\mathscr{G}}$ is a reflexive sheaf on $X,$ then it is torsion-free and the codimension of the singular set is $\geq 3$ [19, Cap. V, Cor. 5.20], [13, Cor. 1.4]. Furthermore, if $C$ is a closed subset of $X$ with codimension $\geq 2,$ then the restriction $\Gamma(X,\,{\mathscr{G}})\to\Gamma(X\setminus C,\,{\mathscr{G}})$ is an isomorphism [19, Cap. V, Prop. 5.21] [14, Prop. 1.11].
Let us suppose that ${\mathscr{F}}$ is a reflexive sheaf on the Kähler manifold $X$ endowed with a Hermitian metric. Then ${\mathscr{E}nd}({\mathscr{F}})$ is also a reflexive sheaf [19, Chap. V, Prop. 4.15]. If ${\mathcal{S}}$ is the singular locus of ${\mathscr{F}},$ let ${\mathcal{W}}$ be a closed subspace of $X$ containing ${\mathcal{S}}$ and such that ${\rm codim}\,{\mathcal{W}}=2.$ Let ${\mathcal{N}}$ denote a tubular neighborhood of ${\mathcal{S}}$ in ${\mathcal{W}}.$ Therefore, the vector space $\Gamma(X\setminus{\mathcal{N}},\,\Omega^{k}({\mathscr{E}nd}({\mathscr{F}})))\simeq\Gamma(X,\,\Omega^{k}({\mathscr{E}nd}({\mathscr{F}}))),$ and the latter is finite dimensional. Furthermore, as $(X\setminus{\mathcal{N}})\cap{\mathcal{S}}=\emptyset,\,$ ${\mathscr{H}om}_{{\mathscr{O}}_{X}}({\mathscr{F}},\,\Omega^{k}({\mathscr{F}}))|_{X\setminus{\mathcal{N}}}$ is a locally free sheaf.
Let $\nabla$ be a holomorphic gauge field on ${\mathscr{F}}$, then
${\mathcal{K}}_{\nabla}\in\Gamma\big{(}X,\,{\mathscr{H}om}({\mathscr{F}},\,\Omega^{2}({\mathscr{F}}))\big{)}=\Gamma(X\setminus{\mathcal{N}},\,{\mathscr{H}om}({\mathscr{F}},\,\Omega^{2}({\mathscr{F}}))\big{)}$
can considered as a global section of the locally free sheaf $\big{(}\Omega^{2}\otimes{\rm End}(F)\big{)}|_{X\setminus{\mathcal{S}}}.$ A similar observation is also valid for $\nabla$.∇ .
Let us assume that the sheaf ${\mathscr{F}}$ supports a holomorphic connection $\nabla_{0}$. By Proposition 1, given ${\mathcal{E}}_{1},\dots\mathcal{E}_{m}$, a basis of
(3.6) ${\rm Hom}({\mathscr{F}},\,\Omega^{1}({\mathscr{F}}))=\Gamma\big{(}Y,\,{\mathscr{H}om}_{{\mathscr{O}}_{X}}({\mathscr{F}},\Omega^{1}({\mathscr{F}}))\big{)}=\Gamma(Y\setminus{\mathcal{S}},\,\Omega^{1}\otimes{\rm End}(F)),$
any holomorphic gauge field can be written $\nabla=\nabla_{0}+\sum\lambda_{i}{\mathcal{E}}_{i},$ with $\lambda_{i}\in{\mathbb{C}}$. The curvature
(3.7) ${\mathcal{K}}_{\nabla}=\nabla\circ\nabla={\mathcal{K}}_{\nabla_{0}}+\sum_{i}\lambda_{i}{\mathcal{B}}_{i}+\sum_{ij}\lambda_{i}\lambda_{j}{\mathcal{B}}_{ij},$
where ${\mathcal{B}}_{i}:=\nabla_{0}({\mathcal{E}}_{i})$ and ${\mathcal{B}}_{ij}:={\mathcal{E}}_{i}\wedge{\mathcal{E}}_{j}$. If moreover ${\mathscr{F}}$ hermitian, then
$\lVert{\mathcal{K}}_{\nabla}\rVert^{2}=({\mathcal{K}}_{\nabla},\,{\mathcal{K}}_{\nabla})=P(\lambda_{1},\dots,\lambda_{m}),$
where $P$ is a polynomial of degree $\leq 4$ in the variables $\lambda_{i}$.
The Yang-Mills fields are those $\nabla$ defined by constants $\lambda_{i}$ which satisfy the algebraic equations of degree $\leq 3$
(3.8) $\frac{\partial\,P}{\partial\lambda_{i}}=0,\;\;\;i=1,\dots,m.$
Therefore,
Theorem 3.
If the Hermitian reflexive sheaf ${\mathscr{F}}$ admits a holomorphic gauge field and $m={\rm dim}\,{\rm Hom}({\mathscr{F}},\,\Omega^{1}({\mathscr{F}})).$ Then the set ${\sf YM}({\mathscr{F}})$ of holomorphic Yang-Mills fields on ${\mathscr{F}}$ is in bijective correspondence with the points of the algebraic set in ${\mathbb{C}}^{m}$ defined by $m$ algebraic equations.
In particular, if $m=2$ and the cardinal of ${\sf YM}({\mathscr{F}})$ is finite, then #$\#{\sf YM}({\mathscr{F}})\leq 9$.
Proof. The case $m=2$ is a consequence of Bézout’s theorem. ∎
Taking into account Proposition 1 and (3.6), any “variation” of a holomorphic connection $\nabla$ on ${\mathscr{F}}$ can be written as $\nabla_{\epsilon}=\nabla+\epsilon E,$ with $\epsilon\in{\mathbb{C}}$ and $E\in\Gamma(X\setminus S,\,\Omega^{1}\otimes{\rm End}(F)).$
(3.9) $(1/2)\frac{d}{d\epsilon}\Big{|}_{\epsilon=0}||K_{\nabla_{\epsilon}}||^{2}=\int_{X\setminus{\mathcal{N}}}\langle K_{\nabla},\,\nabla E\rangle\,{\rm dvol}=:(K_{\nabla},\,\nabla{E}).$
Therefore, $\nabla$ is a Yang-Mills field if for any “variation” ${\mathcal{E}}$ of $\nabla$
(3.10) $(K_{\nabla},\,\nabla{E})=0.$
In particular, the flat holomorphic gauge fields are Yang-Mills.
We denote
${}^{(p)}\nabla:\Gamma(X,\,\Omega^{p}({\mathscr{E}nd}({\mathscr{F}})))\to\Gamma(X,\,\Omega^{p+1}({\mathscr{E}nd}({\mathscr{F}}))),$
the operator defined by the connection $\nabla$. In this notation Bianchi’s identity is read as
${}^{(2)}\nabla K_{\nabla}=0.$
On the other hand, the orthogonality condition (3.10) which satisfy the Yang-Mills fields gives rise to the following proposition.
Proposition 7.
The holomorphic gauge field $\nabla$ on the Hermitian reflexive sheaf ${\mathscr{F}}$ is a Yang-Mills field iff its curvature $K_{\nabla}\in\Gamma(Y,\,\Omega^{2}\otimes_{\mathscr{O}}{\rm End}(F))$ is orthogonal to the vector space ${\rm Im}\,(^{(1)}\nabla)$.
3.1.3. The case ${\rm rank}\,{\mathscr{F}}=1$
In this case the reflexive sheaf ${\mathscr{F}}$ is a locally free sheaf [13, Prop. 1.9]. Then ${\mathscr{E}nd}({\mathscr{F}})$ is the sheaf associated to the trivial line bundle ×${\mathbb{C}}\times X\to X$. If $s$ is a local frame of the corresponding line bundle $F,$ a holomorphic connection $\nabla$ on ${\mathscr{F}}$ on this frame is determined by a ${\mathbb{C}}$-valued $1$-form $A$, $\nabla s=As$. In this frame ${}^{(p)}\nabla(\beta)=\partial\beta+A\wedge\beta-(-1)^{p}\beta\wedge A=\partial\beta,$ for any ${\mathbb{C}}$-valued $p$-form $\beta$. That is,
(3.11) $^{(p)}\nabla=\partial.$
In this case (3.7) reduces to $K_{\nabla}=K_{\nabla_{0}}$, since $\nabla_{0}(E_{i})=\partial E_{i}=0.$ In particular, the Yang-Mills functional is constant.
On the other hand, the Bianchi’s identity reduces to $\partial K_{\nabla}=0$. If furthermore, $\nabla$ is a Yang-Mills field, from (3.10) it follows $\partial^{\dagger}(K_{\nabla})=0$, where $\partial^{\dagger}$ is the adjoint of $\partial$. That is, $K_{\nabla}$ is $\partial$-harmonic. As $X$ is a Kähler manifold, $K_{\nabla}$ is also $d$-harmonic. Hence, the norm of $K_{\nabla}$ minimizes the corresponding norm in its cohomology class. That is, denoting by ${\bf c}$ the cohomology class defined by $K_{\nabla},$
(3.12) $\mathcal{YM}(\nabla)=\lVert K_{\nabla}\rVert^{2}={\min}\{\lVert\beta\rVert^{2}\,|\,\beta\in{\bf c}\}.$
As we are assuming that ${\mathscr{F}}$ supports a holomorphic gauge field, the first Chern class $c_{1}({\mathscr{F}})=[(2\pi)^{-1}K_{\nabla}]$ vanishes [4]. Consequently, ${\bf c}=0$ and $\mathcal{YM}$ is the functional zero.
3.2. Yang-Mills fields on $B$-branes
In this subsection we assume that $X$ be a smooth projective variety. Then any object of $D^{b}(X)$ is isomorphic to a bounded complex consisting of locally free sheaves [24, Sect. 36.36]. Thus, we assume that the $B$-brane ${\mathscr{F}}^{\bullet}$ on $X$ is a bounded complex of locally free sheaves on $X.$ In this case, the jet sheaf ${\mathscr{J}}^{1}({\mathscr{F}}^{\bullet})$ is the complex of vector bundles ${\mathscr{F}}^{\bullet}\oplus\Omega^{1}({\mathscr{F}}^{\bullet})$ endowed with ${\mathscr{O}}_{X}$-module structure (2.9).
[…]
${\mathscr{H}}^{i}({\mathscr{M}}^{\bullet})\overset{(s^{i})^{-1}}{\rightarrow}{\mathscr{H}}^{i}({\mathscr{G}}^{\bullet})\overset{f^{i}}{\rightarrow}{\mathscr{H}}^{i}({\mathscr{N}}^{\bullet}).$
[…]
Thus, a gauge field on ${\mathscr{F}}^{\bullet}$, i.e. a morphism $\psi\in{\rm Hom}_{D^{b}(X)}\big{(}{\mathscr{F}}^{\bullet},\,{\mathscr{J}}^{1}({\mathscr{F}}^{\bullet})\big{)}$ satisfying $\pi\circ\psi={\rm id},$ determines a unique morphism
$\psi^{i}:{\mathscr{H}}^{i}({\mathscr{F}}^{\bullet})\to{\mathscr{H}}^{i}({\mathscr{J}}^{1}({\mathscr{F}}^{\bullet}))={\mathscr{H}}^{i}({\mathscr{F}}^{\bullet})\,\tilde{\oplus}\,\Omega^{1}({\mathscr{H}}^{i}({\mathscr{F}}^{\bullet})),$
such that the composition with the projection ${\mathscr{H}}^{i}({\mathscr{J}}^{1}({\mathscr{F}}^{\bullet}))\overset{\pi^{i}}{\to}{\mathscr{H}}^{i}({\mathscr{F}}^{\bullet})$ is the identity on ${\mathscr{H}}^{i}({\mathscr{F}}^{\bullet}).$ That is, $\psi^{i}$ is a holomorphic connection on the sheaf ${\mathscr{H}}^{i}({\mathscr{F}}^{\bullet}).$
We set $\eta^{j}$ for the morphism of abelian sheaves defined by the inclusion in the direct sum $\eta^{j}:{\mathscr{H}}^{j}({\mathscr{F}}^{\bullet})\rightarrow{\mathscr{H}}^{j}({\mathscr{F}}^{\bullet})\oplus\Omega^{1}({\mathscr{H}}^{j}({\mathscr{F}}^{\bullet})).$ Hence, $\pi^{j}(\psi^{j}-\eta^{j})=0$, and thus $\psi^{j}-\eta^{j}$ defines a morphism of abelian sheaves
(3.13) $\vartheta^{j}:{\mathscr{H}}^{j}({\mathscr{F}}^{\bullet})\to{\Omega^{1}}({\mathscr{H}}^{j}({\mathscr{F}}^{\bullet})),$
which, by (2.9), satisfies the Leibniz rule. That is,
Proposition 8.
The gauge field $\psi$ on the brane ${\mathscr{F}}^{\bullet}$ determines on each ${\mathscr{O}}_{X}$-module ${\mathscr{H}}^{j}({\mathscr{F}}^{\bullet})$ a holomorphic connection $\vartheta^{j}$.
Remark 6.
Let $\psi,\phi$ be two gauge fields on ${\mathscr{F}}$, we set
$\xi:=\phi-\psi\in{\rm Hom}_{D^{b}(X)}\big{(}{\mathscr{F}}^{\bullet},\,{\mathscr{J}}^{1}({\mathscr{F}}^{\bullet})\big{)}.$
Thus, $\xi$ determines a well defined morphism of ${\mathscr{O}}_{X}$-modules between the cohomologies, $\xi^{j}:{\mathscr{H}}^{j}({\mathscr{F}}^{\bullet})\to{\mathscr{H}}^{j}({\mathscr{J}}^{1}({\mathscr{F}}^{\bullet})).$
We denote by $\vartheta^{j}$ and $\chi^{j}$ the connections on ${\mathscr{H}}^{j}({\mathscr{F}}^{\bullet})$ determined by $\psi$ and $\phi$, respectively. Since $\xi^{j}=(\phi^{j}-\eta^{j})-(\psi^{j}-\eta^{j})$, it follows that $\xi^{j}=\vartheta^{j}-\chi^{j}$. In short, $\xi^{j}$ is the “variation” on the connection $\vartheta^{j}$ induced by the “variation” $\xi$ of the gauge field $\psi$.
The result deduced in the following paragraph gives us a suggestion for the definition of the Yang-Mills functional on the gauge fields on a brane over a Kähler manifold.
3.2.1. An Euler-Poincaré mapping.
Let ${\mathscr{A}}$ be a coherent sheaf on the Kähler manifold $X,$ and $\alpha:{\mathscr{A}}\to\Omega^{1}({\mathscr{A}})$ a holomorphic connection on ${\mathscr{A}}$. Denoting by ${\mathcal{S}}_{\mathscr{A}}$ the singular set of ${\mathscr{A}}$, on $Y\setminus{\mathcal{S}}_{\mathscr{A}}$ we define differential form
$\Phi({\mathscr{A}},\alpha):={\rm tr}\big{(}K_{\alpha}\wedge\star K_{\alpha}\big{)}\in\Gamma(X\setminus{\mathcal{S}}_{\mathscr{A}},\,\Omega^{\rm top}\big{)},$
where $K_{\alpha}$ is the curvature of $\alpha$, considered as an ${\rm End}(A)$-valued $2$-form.
By $\mathfrak{C}$, we denote the category whose objects are pairs $({\mathscr{A}},\,\alpha)$. A morphism $f:({\mathscr{A}},\,\alpha)\to({\mathscr{B}},\,\beta)$ is a morphism of coherent sheaves compatible with the connections; i.e. such that $(1\otimes f)\circ\alpha=\beta\circ f$.
Proposition 9.
[…]
$\Phi({\mathscr{B}},\beta)=\Phi({\mathscr{A}},\alpha)+\Phi({\mathscr{C}},\gamma),$
where ${\mathcal{S}}$ is the union of the singular sets of ${\mathscr{A}},$ ${\mathscr{B}},$ and ${\mathscr{C}}$.
Proof. Let $x_{0}\in X\setminus{\mathcal{S}}$. As the exact sequence splits locally on $X\setminus{\mathcal{S}}$, there exists an open neighborhood $U$ of $x_{0}$ such that $g|_{U}$, in the sequence of locally free modules $0\to{\mathscr{A}}|_{U}\overset{f|_{U}}{\to}{\mathscr{B}}|_{U}\overset{g|_{U}}{\to}{\mathscr{C}}|_{U}\to 0,$ has a right inverse $h$.
Let $a$ be a frame for ${\mathscr{A}}|_{U}$, then $\alpha(a)={\sf A}\cdot a$, where ${\sf A}$ is a matrix of $1$-forms on $U$. Furthermore, $a$ can be chosen so that ${\sf A}(x_{0})=0$. Similarly, let $c$ be a frame for ${\mathscr{C}}|_{U}$, then $\gamma(c)={\sf C}\cdot c$ and we choose $c$ so that ${\sf C}(x_{0})=0$. From the splitting, it follows that $\{f(a),\,h(c)\}$ is a frame for ${\mathscr{B}}|_{U}$. By the compatibility of the connections with $f$ and $g$,
$\beta(f(a))=(1\otimes f)(\alpha(a))=(1\otimes f)({\sf A}\cdot a)={\sf A}\cdot f(a).$
On the other hand, $\beta(h(c))={\sf R}\cdot f(a)+{\sf S}\cdot h(c)$, with ${\sf R}$ and ${\sf S}$ matrices of $1$-forms. But,
${\sf C}\cdot c=\gamma(c)=\gamma(gh(c))=(1\otimes g)(\beta(h(c))=(1\otimes g)\big{(}{\sf R}\cdot f(a)+{\sf S}\cdot h(c)\big{)}.$
As $g\circ f=0$ and $g\circ h=1$, it follows that ${\sf C}={\sf S}.$ That is, the matrix of the connection $\beta$ in the frame $\{f(a),\,h(c)\}$ is
(3.14) ${\sf M}:=\begin{pmatrix}{\sf A}&{\sf R}\\ 0&{\sf C}\end{pmatrix}$
Since ${\sf A}(x_{0})=0$ and ${\sf C}(x_{0})=0$, the matrix of $K_{\alpha}(x_{0})$, of the curvature of $\alpha$ at the point $x_{0}$, is $d{\sf A}$. Analogously, the matrix of $K_{\gamma}(x_{0})$ is $d{\sf C}$. The one of $K_{\beta}(x_{0})$ is the exterior derivative of (3.14), since ${\sf M}\wedge{\sf M}=0$ at $x_{0}$. Then
${\rm tr}\big{(}K_{\beta}(x_{0})\wedge\star K_{\beta}(x_{0})\big{)}$ $={\rm tr}(d{\sf A}\wedge\star d{\sf A})+{\rm tr}(d{\sf C}\wedge\star d{\sf C})$
$={\rm tr}\big{(}K_{\alpha}(x_{0})\wedge\star K_{\alpha}(x_{0})\big{)}+{\rm tr}\big{(}K_{\gamma}(x_{0})\wedge\star K_{\gamma}(x_{0})\big{)}.$
As $x_{0}$ is an arbitrary point of $X\setminus{\mathcal{S}}$, it follows the proposition. ∎
Let $({\mathscr{G}}^{\bullet},\,{\delta}^{\bullet})$ be a bounded complex of coherent sheaves on the manifold $X$. Let $\nabla^{\bullet}$ be a family of holomorphic connections, compatible with the operators $\delta^{\bullet}$. That is, $\nabla^{i}:{\mathscr{G}}^{i}\rightarrow\Omega^{1}({\mathscr{G}}^{i})$ is a holomorphic connection on the coherent sheaf ${\mathscr{G}}^{i}$ such that $(1\otimes\delta^{i})\nabla^{i}=\nabla^{i+1}\delta^{i}.$ Hence, $\nabla^{i}({\mathscr{K}er}({\delta}^{i}))\subset{\mathscr{K}er}(1\otimes{\delta}^{i})$ and a similar relation for the image ${\mathscr{I}m}(\delta^{i-1})$. It follows that $\nabla^{i}$ induces a connection $\theta^{i}$ on the cohomology $\theta^{i}:{\mathscr{H}}^{i}({\mathscr{G}}^{\bullet})\to\Omega^{1}({\mathscr{H}}^{i}({\mathscr{G}}^{\bullet})).$ Obviously, the restrictions of $\nabla^{i}$ determine connections on ${\mathscr{K}er}(\delta^{i})$ and ${\mathscr{I}m}(\delta^{i+1})$, respectively. One has the exact sequence
(3.15) $0\to\big{(}{\mathscr{K}er}(\delta^{i}),\,\nabla^{i}\big{)}\to\big{(}{\mathscr{G}}^{i},\,\nabla^{i}\big{)}\to\big{(}{\mathscr{I}m}(\delta^{i}),\,\nabla^{i+1}\big{)}\to 0$
in the category $\mathfrak{C}$. Similarly, we have the exact sequence
(3.16) $0\to\big{(}{\mathscr{I}m}(\delta^{i-1}),\,\nabla^{i}\big{)}\to\big{(}{\mathscr{K}er}(\delta^{i}),\,\nabla^{i}\big{)}\to\big{(}{\mathscr{H}}^{i},\,\theta^{i}\big{)}\to 0$
Corollary 3.
Denoting with ${\mathcal{S}}$ the union of the singular sets of the sheaves ${\mathscr{G}}^{i}$, then $\,\sum_{i}(-1)^{i}{\rm tr}\big{(}K_{\nabla^{i}}\wedge\star K_{\nabla^{i}}\big{)}=\sum_{i}(-1)^{i}{\rm tr}\big{(}K_{\theta^{i}}\wedge\star K_{\theta^{i}}\big{)}$ on $X\setminus{\mathcal{S}}.$
Proof. From Proposition 9 together with (3.15), it follows
$\Phi({\mathscr{G}}^{i},\,\nabla^{i})=\Phi({\mathscr{K}er}(\delta^{i}),\,\nabla^{i})+\Phi({\mathscr{I}m}(\delta^{i}),\,\nabla^{i+1}).$
From (3.16), one obtains an analogous relation. Taking the alternate sums
$\sum_{i}(-1)^{i}\Phi({\mathscr{G}}^{i},\,\nabla^{i})=\sum_{i}(-1)^{i}\Phi({\mathscr{H}}^{i},\,\theta^{i}).\qed$
3.2.2. The Yang-Mills functional
We propose a definition for the Yang-Mills functional over gauge fields on a brane. This proposal is based on the following considerations:
  1. (1)
    It is reasonable to require that this definition generalizes the one for coherent sheaves.
  2. (2)
    As a gauge field $\psi$ on ${\mathscr{F}}^{\bullet}$ is a class of “roofs” [9, p. 148] from the complex ${\mathscr{F}}^{\bullet}$ to ${\mathscr{J}}^{1}({\mathscr{F}}^{\bullet}),$ and equivalent roofs determine the same morphisms between the cohomologies, it seems convenient to move on the cohomology of these complexes.
  3. (3)
    Let $E^{\bullet}$ be a bounded complex of Hermitian vector bundles over the Kähler manifold $Y,$ and $\nabla^{\bullet}$ a family of connections compatible with the Hermitian metrics and the coboundary operators. Denoting by ${\mathscr{H}}^{i}(E^{\bullet})$ the cohomology sheaves, there exist connections $\theta^{i}$ on those sheaves, induced by the family $\nabla^{\bullet}$. By Corollary 3 together with (3.5), one has the following equality of Euler-Poincaré type.
    $\sum_{i}(-1)^{i}\lVert K_{\nabla^{i}}\rVert^{2}=\sum_{i}(-1)^{i}\lVert(K_{\theta^{i}}\rVert^{2}.$
On the basis of the above considerations, it seems appropriate to define the value of the Yang-Mills functional on the gauge $\psi$ on the brane ${\mathscr{F}}^{\bullet}$ as $\sum_{i}(-1)^{i}\lVert K_{\vartheta_{i}}\rVert^{2}.$ More precisely, taking into account Proposition 8, we adopt the following definitions.
Definition 4.
The brane $({\mathscr{F}}^{\bullet},\,\delta^{\bullet})$ is called a Hermitian brane, if the cohomology sheaves ${\mathscr{H}}^{j}$ are Hermitian ${\mathscr{O}}_{X}$-modules.
Let $({\mathscr{F}}^{\bullet},\,{\delta}^{\bullet})$ be a Hermitian brane on the Kähler manifold $X$. Given a gauge field $\psi$ on the brane $({\mathscr{F}}^{\bullet},\,\delta^{\bullet})$, by Proposition 8, one has the family of curvatures ${\mathcal{K}}_{\vartheta^{i}}$ of the connections induced on the cohomologies, whose norms $\lVert{\mathcal{K}}_{\vartheta^{i}}\rVert$ are defined in accordance with (3.4).
Definition 5.
Given a gauge field $\psi$ on the Hermitian $B$-brane $({\mathscr{F}}^{\bullet},\,\delta^{\bullet})$, if the sheaves ${\mathscr{H}}^{i}({\mathscr{F}}^{\bullet})$ are reflexive, we define the value of the Yang-Mills functional at $\psi$ by
(3.17) $\mathcal{YM}(\psi)=\sum_{i}(-1)^{i}\lVert{\mathcal{K}}_{\vartheta_{i}}\rVert^{2}.$
Thus, $\mathcal{YM}(\psi)$ is a kind of Euler characteristic of the gauge field. The Yang-Mills fields on the brane $({\mathscr{F}}^{\bullet},\,{\delta}^{\bullet})$ are the critical points of the functional $\psi\mapsto\mathcal{YM}(\psi).$
Note that if $({\mathscr{F}}^{\bullet},\,\delta^{\bullet})$ is an acyclic complex, then the Yang-Mills functional for this complex is identically zero.
Let ${\mathscr{A}}^{\bullet}:=\big{(}{\mathscr{A}}^{\bullet},\,d_{A}^{\bullet},\,\alpha^{\bullet}\big{)}$ be a complex in the category ${\mathfrak{C}}$; i.e, a complex of coherent sheaves with a family of holomorphic connections compatible with the coboundary operator $d_{A}$. Let $f:=(f^{\bullet})$ a morphism $f^{\bullet}:{\mathscr{A}}^{\bullet}\to{\mathscr{B}}^{\bullet}$ in ${\mathfrak{C}}$; that is, $f$ is a morphism of complexes compatible with the connections. Let ${\mathscr{C}}^{\bullet}$ denote the mapping cone of $f$. Thus, ${\mathscr{C}}^{\bullet}=\big{(}{\mathscr{A}}^{\bullet}[1]\oplus{\mathscr{B}}^{\bullet},d_{C}^{\bullet},\nabla^{\bullet},\big{)}$, with $d_{C}(a,\,b)=\big{(}d(a),\,(-1)^{{\rm degree}\,a}f(a)+db\big{)}$ and $\nabla(a,\,b)=(\alpha(a),\,\beta(b)).$ In fact, $(1\otimes d_{C})\circ\nabla=\nabla\circ d_{C}$ and thus ${\mathscr{C}}^{\bullet}$ is a complex of the category ${\mathfrak{C}}$
For each $i$ one has the following exact sequence in the category ${\mathfrak{C}}$
$0\to{\mathscr{B}}^{i}\to{\mathscr{C}}^{i}\to{\mathscr{A}}^{i}[1]\to 0$
From Proposition 9, $\Phi({\mathscr{B}}^{i})+\Phi({\mathscr{A}}^{i+1})=\Phi({\mathscr{C}}^{i})$. Multiplying by $(-1)^{i}$ and summing
(3.18) $\sum_{i}(-1)^{i}{\rm tr}(K_{\beta^{i}}\wedge\star K_{\beta^{i}})+$ $\sum_{i}(-1)^{i}{\rm tr}(K_{\alpha^{i+1}}\wedge\star K_{\alpha^{i+1}})$
$=\sum_{i}(-1)^{i}{\rm tr}(K_{\nabla^{i}}\wedge\star K_{\nabla^{i}}).$
Let us assume that
  • ${\mathscr{A}}^{i}$ and ${\mathscr{B}}^{i}$ Hermitian sheaves for all $i$.
  • $\alpha^{i}$ and $\beta^{i}$ are Hermitian gauge fields (i.e., compatible with the metric) on ${\mathscr{A}}^{i}$ and ${\mathscr{B}}^{i}$, respectively.
Then one defines on ${\mathscr{C}}^{i}$ the metric $\langle(a,b),\,(a^{\prime},b^{\prime})\rangle:=\langle a,\,a^{\prime}\rangle+\langle b,\,b^{\prime}\rangle$. The connection $\nabla^{i}$ is compatible with this metric. From the equality (3.18) together with (3.5), one deduces the following proposition.
Proposition 10.
With the above notations and under the above hypotheses, $\alpha$ and $\beta$ determine in a natural way a gauge field $\nabla$ on the mapping cone of $f^{\bullet}$ satisfying
(3.19) $\mathcal{YM}(\beta)-\mathcal{YM}(\alpha)=\mathcal{YM}(\nabla).$
On the other hand, in the context of the branes theory, the fact that the branes ${\mathscr{A}}^{\bullet},\,{\mathscr{B}}^{\bullet}$ and ${\mathscr{C}}^{\bullet}$ are the members of the distinguished triangle ${\mathscr{A}}^{\bullet}\to{\mathscr{B}}^{\bullet}\to{\mathscr{C}}^{\bullet}\overset{+1}{\to}$ means that ${\mathscr{A}}^{\bullet}$ and ${\mathscr{C}}^{\bullet}$ can potentially bind together to form the membrane ${\mathscr{B}}^{\bullet}$ [2, Section 6.2.1]. Thus, the additive nature of equation (3.19) is consistent with this interpretation.
From now on, we assume that ${\mathscr{F}}^{\bullet}$ is a Hermitian $B$-brane such that the cohomology sheaves ${\mathscr{H}}^{i}({\mathscr{F}}^{\bullet})$ are reflexive.
Let us suppose that the set of gauge fields on the brane ${\mathscr{F}}^{\bullet}$ is nonempty. Let $m$ be the dimension of the vector space ${\rm Ext}^{0}({\mathscr{F}}^{\bullet},\,\Omega^{1}({\mathscr{F}}^{\bullet})).$ We denote by $\xi_{1},\dots,\xi_{m}$ a basis of this vector space. According to Proposition 1, any gauge field $\psi$ on the brane can be expressed
$\psi=\tilde{\psi}+\sum_{a}\lambda_{a}\xi_{a}$
$\tilde{\psi}$ being a fixed gauge field and $\lambda_{a}\in{\mathbb{C}}.$ Hence, the connections on the cohomology sheaves ${\mathscr{H}}^{i}$ can be written in the form (see Remark 6)
$\vartheta^{i}=\tilde{\vartheta}^{i}+\sum_{a}\lambda_{a}\xi^{i}_{a},$
with $\xi^{i}_{a}\in{\rm Hom}({\mathscr{H}}^{i},\,\Omega^{1}({\mathscr{H}}^{i}))$. The corresponding curvatures satisfy
$K_{\vartheta^{i}}=K_{\tilde{\vartheta}^{i}}+\sum_{a}\lambda_{a}\tilde{\vartheta}^{i}(\xi^{i}_{a})+\sum_{a,b}\lambda_{a}\lambda_{b}\xi^{i}_{a}\wedge\xi^{i}_{b}.$
Therefore, $\lVert K_{\vartheta^{i}}\rVert^{2}$ is a polynomial $P^{i}(\lambda_{1},\dots,\lambda_{m})$ of degree $\leq 4.$ Thus, the critical points of the Yang-Mills functional correspond to the points $(\lambda_{1},\dots,\lambda_{m})\in{\mathbb{C}}^{m}$ which satisfy the equations $\frac{\partial P}{\partial\lambda_{a}}=0,$ where $P$ is the polynomial $\sum_{i}(-1)^{i}P^{i}(\lambda_{1},\dots,\lambda_{m})$. We have the following result, which generalizes Theorem 3.
Theorem 4.
Assumed the set of gauge fields on the brane ${\mathscr{F}}^{\bullet}$ is nonempty and $m={\rm dim}\,{\rm Ext}^{0}({\mathscr{F}}^{\bullet},\,\Omega^{1}({\mathscr{F}}^{\bullet}))$. Then the set of Yang-Mills fields on ${\mathscr{F}}^{\bullet}$ are in bijective correspondence with the points of a subvariety of ${\mathbb{C}}^{m}$ defined by $m$ polynomials of degree $\leq 3$.
Let $({\mathscr{F}}^{\bullet},\,\delta^{\bullet})$ be a $B$-brane, such that the ${\mathscr{H}}^{i}({\mathscr{F}}^{\bullet})$ are reflexive sheaves. If $\phi$ and $\psi$ are gauge fields on ${\mathscr{F}}^{\bullet}$ and $\xi=\phi-\psi$, using the notations introduced in Remark 6, the connections on the cohomologies induced by $\phi$ and $\psi$ satisfy $\chi^{j}(\xi)=\vartheta^{j}+\xi^{j}$, with
$\xi^{j}\in\Gamma(Y\setminus{\mathcal{S}},\,\Omega^{1}\otimes_{\mathscr{O}}{\mathscr{E}nd}(H^{j})),$
${\mathcal{S}}$ being the union of the singularity sets of the reflexive sheaves ${\mathscr{H}}^{j}.$
With the mentioned notation, an infinitesimal variation $\psi_{\epsilon}$ of $\psi$ is an element of the form $\epsilon\xi,$ where $\epsilon\in{\mathbb{C}}$ and $\xi\in{\rm Ext}^{0}({\mathscr{F}}^{\bullet},\,\Omega^{1}({\mathscr{F}}^{\bullet}).$ The relation $\psi_{\epsilon}=\psi+\epsilon\xi$ gives rise to the following equality between the connections on the cohomologies
$\vartheta^{j}_{\epsilon}=\vartheta^{j}+\epsilon\xi^{j}.$
Furthermore, on $Y\setminus{\mathcal{S}}$ the curvatures satisfy
$K_{\vartheta^{j}_{\epsilon}}=K_{\vartheta^{j}}+\epsilon\vartheta^{j}({\xi^{j}})+O(\epsilon^{2}),$
$\vartheta^{j}({\xi^{j}})$ being the covariant derivative of ${\xi^{j}}$ considered as a section of $\Omega^{1}\otimes_{\mathscr{O}}{\mathscr{E}nd}(H^{j})$.
From (3.9) together with Definition 5, the functional $\mathcal{YM}$ takes at the gauge field ${\psi}$ a stationary value if
(3.20) $(1/2)\frac{d}{d\epsilon}\mathcal{YM}(\psi_{\epsilon})\Big{|}_{\epsilon=0}=\sum_{j}(-1)^{j}\langle K_{\vartheta^{j}},\,\vartheta^{j}(\xi^{j})\rangle=0,$
for all any variation of $\psi$. In particular, if $\vartheta^{i}$ es a Yang-Mills field for all $i,$ then by (3.10) $\langle K_{\vartheta^{i}},\,\vartheta^{i}(\xi^{i})\rangle=0,$ and we have the following proposition.
Proposition 11.
Let $\psi$ gauge field on the brane $({\mathscr{F}}^{\bullet},\,\delta^{\bullet}).$ If $\vartheta^{i}$ is a Yang-Mills field on ${\mathscr{H}}^{i}$ for all $i$, then $\psi$ is a Yang-Mills field on the brane.
The following theorem is a partial converse to Proposition 11.
Theorem 5.
If each ${\mathscr{F}}^{i}$ is a semisimple object in the category of coherent sheaves on $X$ and $\nabla^{\bullet}$ is a Yang-Mills field on ${\mathscr{F}}^{\bullet},$ then the connection $\vartheta^{j}$ induced on ${\mathscr{H}}^{j}$ is a Yang-Mills field on this sheaf.
Proof. Since ${\mathscr{F}}^{i}$ is semisimple the following short exact sequence of coherent ${\mathscr{O}}_{X}$-modules
$0\to{\mathscr{K}er}(\delta^{i})\to{\mathscr{F}}^{i}\to{\mathscr{C}oim}(\delta^{i})\to 0$
splits. That is, ${\mathscr{F}}^{i}\simeq{\mathscr{K}er}(\delta^{i})\oplus{\mathscr{C}oim}(\delta^{i}).$ Since ${\mathscr{K}er}(\delta^{i})$ is semisimple the exact sequence $0\to{\mathscr{I}m}({\delta}^{i-1})\to{\mathscr{K}er}(\delta^{i})\to{\mathscr{H}}^{i}\to 0$ also splits Thus,
(3.21) ${\mathscr{F}}^{i}\simeq{\mathscr{H}}^{i}\oplus{\mathscr{G}}^{i},$
where ${\mathscr{G}}^{i}$ is isomorphic to the direct sum of ${\mathscr{C}oim}(\delta^{i})$ and ${\mathscr{I}m}(\delta^{i-1})$.
On the other hand, the coboundary operator $\delta^{i}:{\mathscr{F}}^{i}\to{\mathscr{F}}^{i+1}$ induces via the isomorphisms (3.21) to the morphism
(3.22) $\delta^{i}:{\mathscr{H}}^{i}\oplus{\mathscr{G}}^{i}\to{\mathscr{H}}^{i+1}\oplus{\mathscr{G}}^{i+1},\;\;\;(a,\,b)\mapsto(0,\,\delta^{i}b).$
Given $\xi\in{\rm Hom}_{D^{b}(X)}\big{(}{\mathscr{F}}^{\bullet},\,\Omega^{1}\otimes_{\mathscr{O}}{\mathscr{F}}^{\bullet}\big{)}$, according to Remark 6, it determines $\xi^{i}\in{\rm Hom}\big{(}{\mathscr{H}}^{i},\,\Omega^{1}\otimes_{\mathscr{O}}{\mathscr{H}}^{i}\big{)}.$ As $\nabla^{\bullet}$ is, by hypothesis, a Yang-Mills field then (3.20) is satisfied.
Given $j$, a general “variation” of $\vartheta^{j}$ is defined by an element $\tau\in{\rm Hom}\big{(}{\mathscr{H}}^{j},\,\Omega^{1}\otimes_{\mathscr{O}}{\mathscr{H}}^{j}\big{)}.$ Under the hypotheses of the proposition, we need to prove that
$\langle K_{\vartheta^{j}},\,\vartheta^{j}(\tau)\rangle=0,$
for any variation $\tau$. The morphism $\tau$ can be extended to a morphism
$C^{i}:{\mathscr{H}}^{i}\oplus{\mathscr{G}}^{i}\to\Omega^{1}\otimes_{\mathscr{O}}\big{(}{\mathscr{H}}^{i}\oplus{\mathscr{G}}^{i}\big{)},$
where
$C^{i}(a,\,b)=\begin{cases}(\tau(a),\,0),\;\;\text{if}\;\,i=j\\ (0,\,0),\;\;\text{if}\;\,i\neq j\end{cases}$
[…]
$0=\sum_{i}(-1)^{i}\langle K_{\vartheta^{i}},\,\vartheta^{i}(\xi^{i})\rangle=(-1)^{j}\langle K_{{\vartheta}^{j}},\,\vartheta^{j}(\xi^{j})\rangle=(-1)^{j}\langle K_{{\vartheta}^{j}},\,\vartheta^{j}(\tau)\rangle.$
This holds for any “variation” $\tau$ of $\vartheta^{j}$. That is, by (3.10), $\vartheta^{j}$ is a Yang-Mills field on ${\mathscr{H}}^{j}$. ∎
4. Final Remarks
The following are some aspects and assumptions made in the article that could be considered to generalize or extend the derived results.
Remark 7.
In the definition of the Yang-Mills functional on the gauge fields on brane ${\mathscr{F}}^{\bullet}$ we assumed that the cohomology sheaves ${\mathscr{H}}^{j}({\mathscr{F}}^{\bullet})$ are reflexive. It would be interesting to extend the definition to holomorphic gauge fields on an arbitrary $B$-brane.
Remark 8.
The semisimplicity assumption in Theorem 5 is a strong hypothesis that would be desirable to weaken.
Remark 9.
In Section 2.3.2, we considered the variety of complete flags in ${\mathbb{C}}^{3}$ and proved Theorem 2, which provides an upper bound for the number of gauge fields on this variety. The restriction to this flag variety arises from the fact that, in this case, any Schubert subvariety $Z$ is smooth. Thus, from the vanishing of the first Betti number, we deduce that $H^{1,0}(Z)=0,$ which is a necessary condition for Proposition 6 to hold. Following an alternative approach, could it be possible to prove a version of Theorem 2 for any complete flag variety?
Remark 10.
The definitions introduced in the article and the basic results pertain to $B$-branes over a complex manifold. One may wonder which parts of the article’s content are generalizable to the context of branes over varieties or schemes. It is expected that anything admitting a formulation in categorical language, such as the definition of jet sheaf or Proposition 1, can be translated into the framework of scheme theory. However, the translation of other concepts originating in differential geometry, such as the Yang-Mills functional, is not straightforward.